### v1

# Energy Conditions and Quantum Information

Abstract
The concept of energy lies at the foundation of physical science. In general relativity and quantum field theory, the positivity and conservation of energy are encapsulated by the so-called energy-momentum tensor and the energy conditions. In recent efforts to unify fundamental physics with quantum information, the energy conditions have come to play a crucial role in establishing numerous theorems. In this article, we review the basics of energy conditions in general relativity and their applications in gravitational physics, quantum field theory, and the holographic principle. Through these applications, we explore the profound connection between the energy conditions and quantum information.
preprint: NU-QG-10, OU-HET-1283
1 Introduction
The notion of energy is fundamental to every scientific discipline and all technological advancements, ranging from the development of smart devices to our understanding of the universe. It is not just actual material systems that require energy; even a single piece of information needs energy to be stored or to be transferred. Quantum information has recently emerged as another pivotal concept for the next generation of science and technology. As exemplified by progress in quantum computing and quantum communication, recent efforts to integrate physical science and technology with this field are rapidly advancing. Even at the most fundamental level, such an attempt to unify physics and information has begun to open up a new research area where the very fabric of spacetime itself is envisioned to be described in terms of quantum information. The basic underlying idea is that energy must be assigned to a qubit—the unit of quantum information—and as a source of gravitation, energy produces spacetime curvature, according to the general theory of relativity. The involvement of spacetime curvature, or gravitation, as an impetus for this attempt appears to be rather inevitable, as the discovery of the Hawking radiation Hawking:1974rv ; Hawking:1975vcx and the associated black hole information paradox Hawking:1976ra can be viewed as having initiated the crucial link between gravity, quantum, and information. According to Hawking, a black hole emits thermal blackbody radiation due to quantum field effects, loses its mass, and eventually evaporates away. However, the entire process of black hole formation and complete evaporation appears to contradict the unitarity of quantum mechanics, posing the paradox: where does the quantum information stored inside a black hole go?
Any physically sensible system is expected to have positive definite energy, or at least a lower bound, to ensure a stable ground state. Also important in any physical theory is the law of the conservation of energy. In general relativity, the effects of matter fields on spacetime and conservation laws [^1] The energy-momentum tensor expresses a local conservation of energy and momentum but it does not immediately lead to global properties of energy and momentum of a matter field. In order to discuss the positivity and conservation laws of “total energy” of a system, one needs to introduce the notion of asymptotic flatness in order to define an isolated system (see Theorem 2.2.4 and Sec. 2.6 below). are captured by the “stress-energy-momentum tensor.” The positivity of energy density is then prescribed by the “energy conditions,” which are inequalities to be imposed on the energy-momentum tensor. As will be seen, depending on the choice of observers who measure energy density and also on the specific circumstances of their imposition, one can consider various different types of energy conditions, among which the most widely used are the weak (WEC), the dominant (DEC), the strong (SEC), and the null energy conditions (NEC) HE73 ; Wald84 . In other words, the choice of energy conditions is the selection of what one thinks physically reasonable matter fields are. In any case, roughly speaking, positive energy density generates the attractive nature of gravitation via the Einstein equations, leading to the focusing of families of world-lines of test particles or photons due to gravitational pull. This focusing effect—i.e., the convergence of causal geodesic curves—is key to understanding the physical consequences of gravitation and to establishing a number of important theorems in general relativity. For example, the singularity theorems of Hawking and Penrose Penrose:1964wq ; Hawking:1970zqf , which predict the existence of spacetime singularities inside a black hole or at the beginning of our universe, exploit certain energy conditions and geodesic focusing in an essential way in their proof. The laws of black hole thermodynamics are shown by using a combination of various energy conditions Hawking:1976de ; Wald:1999vt . Furthermore, by using the energy conditions, one can prohibit some exotic phenomena and structures from occurring, such as time-machines, wormholes, and faster-than-light travel Morris:1988tu ; in other words, such an exotic or pathological spacetime structure necessitates some sort of negative energy density.
The positivity of energy of a physical material system is, however, not a trivial condition, especially when quantum effects are involved; there is no shortage of examples of negative energy in quantum mechanical systems. One of the most well-known examples is the Casimir effect. The Hawking radiation mentioned above also violates the positivity of energy. In fact, any locally (pointwisely) defined energy conditions can, in principle, be violated by quantum field effects [^2] It should also be noticed that even though local energy conditions can fail to hold under some quantum effects, they often admit lower bounds, hence satisfying what is called “quantum energy inequality.” See, e.g., Yurtsever:1994wc ; Fewster:2012yh ; Kontou:2020bta for quantum energy inequalities and their relations to the (averaged) energy conditions. . For this reason, various types of non-local modifications of energy conditions have been proposed, such as those by taking the average of a local energy condition along an observer’s world-line (called the averaged energy conditions), or by including non-local quantum field theoretic quantities, e.g., entanglement entropy of quantum fields (called the quantum null energy conditions). Such non-locally modified energy conditions turn out to play a key role in connecting quantum information with spacetime geometry.
Our view of the connection between spacetime geometry and quantum matter has recently undergone revolutionary changes due to the rise of the holographic principle tHooft:1993dmi ; Susskind:1994vu , which also emerged from the study of thermodynamic aspects of black holes. The basic assertion of the holographic principle is that the physics in a spacetime region can be completely described in terms of the degrees of freedom on the boundary of that region. It is fair to say that the holographic idea is still a conjecture, but a number of models exemplifying this idea have been proposed. So far the most successful and concrete model is the AdS/CFT correspondence Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj (or gauge/gravity duality), which asserts that gravitational theory in a spacetime with negative curvature (called AdS) is equivalent to a certain type of quantum field theory (called CFT) residing on the boundary of the AdS spacetime. In particular, this correspondence makes it possible to analyze strongly coupled quantum fields in terms of weakly coupled classical gravity. Therefore the AdS/CFT correspondence can be a powerful tool to prove, for example, some of the energy conditions, as performed in Kelly:2014mra .
The purpose of this article is to review the basics of energy conditions in general relativity, their uses in gravitational physics, their recent applications in quantum field theory and holography, and their relations to quantum information. This article is not intended to cover the entire range of relevant subjects, to be comprehensive, or to provide a complete list of existing literature. Instead, we will introduce key ideas and methods of the uses of energy conditions and quantum information, by focusing on some specific topics. In the first part, we mainly focus on the four locally defined energy conditions (WEC, DEC, SEC, and NEC), with a particular emphasis on their role in several theorems in general relativity. In the second part, we first discuss how quantum effects lead to the violation of these locally defined energy conditions. Then we review some non-local modifications of the above mentioned energy conditions (ANEC and QNEC). We also discuss the role of (A)NEC in quantum field theory and in the context of the AdS/CFT correspondence.
This article is organized as follows. In the next section, we discuss various types of the pointwise energy conditions and their uses in the context of classical theory of general relativity. In particular, we describe how the energy conditions are used in the singularity theorems, black hole mechanics and topology theorems. Most of the discussion in Sec. 2 is based on Hawking-Ellis HE73 and Wald Wald84 . We also discuss the violation of the pointwise energy conditions and introduce the averaged energy conditions, such as ANEC. A number of examples of exotic spacetime structures, such as wormholes, superluminal travel, and their connection to the violation of the energy conditions can be found in literature, e.g., Lobo:2017cay . More complete list of existing literature on the violation of locally defined, as well as averaged energy conditions can be found in, e.g., Fewster:2012yh ; Kontou:2020bta ; Martin-Moruno:2017exc . In Sec. 3, we review the energy conditions and their derivations in quantum field theory in Minkowski spacetime. We begin with the basic notions of quantum information such as the reduced density matrix and modular Hamiltonian for a subsystem, and introduce entanglement entropy and relative entropy as measures of quantum information. We derive the Bekenstein bound and ANEC from the positivity and monotonicity of relative entropy by following Casini:2008cr and Faulkner:2016mzt , respectively. We also outline the derivation of quantum null energy conditions (QNEC) by using relative entropy in Ceyhan:2018zfg . In Sec. 4, we review the connection between the energy conditions and the holographic principle. In particular, we explain the black hole information paradox in more detail in the context of the AdS/CFT correspondence, introducing the recent advancements such as the holographic entanglement entropy or the Ryu-Takayanagi formula Ryu:2006bv ; Ryu:2006ef , quantum extremal surface (QES) Faulkner:2013ana ; Engelhardt:2014gca , and Island formula Almheiri:2019hni ; Penington:2019npb ; Almheiri:2019psf . Then, we discuss the holographic proof of ANEC and its generalization by focusing on a role of causality in the AdS/CFT correspondence. Section 5 is devoted to summary and discussions.
2 Energy conditions in general relativity
In this section, we will review basic properties of the energy conditions and their applications in general relativity and gravitational physics, including the singularity theorems, the laws of black hole mechanics, and the topological censorship. Our main focus is on the uses of four main, locally (pointwisely) defined energy conditions, but we also discuss their violations and non-local generalizations at the end of this section.
2.1 The Einstein equations
Spacetime geometry can be described in terms of the pair of the metric tensor $g_{\mu\nu}$ with Lorentzian signature and a manifold $M$. The Einstein equations are a set of second-order differential equations for the metric tensor, expressed as
$R_{\mu\nu}-\dfrac{1}{2}Rg_{\mu\nu}=8\pi GT_{\mu\nu}\,$, (1)
where $R_{\mu\nu}$ and $R$ are the Ricci tensor and scalar curvature, respectively, constructed from the Riemann tensor333 The Riemann tensor is defined as follows: $R^{\lambda}{}_{\sigma\mu\nu}=\partial_{\mu}\Gamma^{\lambda}_{\sigma\nu}-\partial_{\nu}\Gamma^{\lambda}_{\sigma\mu}+\Gamma^{\lambda}_{\alpha\mu}\Gamma^{\alpha}_{\sigma\nu}-\Gamma^{\lambda}_{\alpha\nu}\Gamma^{\alpha}_{\sigma\mu}\,$, (2) where $\Gamma^{\lambda}_{\mu\nu}=\dfrac{1}{2}g^{\lambda\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ is the Christoffel symbol and $R_{\mu\nu}:=R^{\alpha}{}_{\mu\alpha\nu}$ and $R:=g^{\mu\nu}R_{\mu\nu}$. For geometric quantities, we follow the convention and notations of Wald’s book Wald84 . . On the right-hand side are the energy-momentum (or stress-energy) tensor $T_{\mu\nu}$ for cosmic matter fields and $G$ denotes Newton’s gravitational constant. Due to the Bianchi-identity the divergence of the left-hand side is automatically vanishing, and accordingly, the divergence of the right-hand side is also required to satisfy,
$\nabla_{\nu}T^{\mu\nu}=0\,$, (3)
where $\nabla_{\mu}$ denotes the covariant derivative operator [^4] The covariant derivative acts on a scalar field $\phi$ as $\nabla_{\mu}\phi=\partial\phi/\partial x^{\mu}=:\partial_{\mu}\phi$ and on a vector field $A^{\mu}$ as $\nabla_{\nu}A^{\mu}=\partial_{\nu}A^{\mu}+\Gamma^{\mu}_{\nu\lambda}A^{\lambda}$. From these formulas, one can deduce how $\nabla_{\mu}$ acts on more general tensor fields. . This is the local law of conservation of material energy.
For example, for a perfect fluid (no heat conduction and viscosity), the energy-momentum tensor is given by
$T_{\mu\nu}=(\rho+P)u_{\mu}u_{\nu}+Pg_{\mu\nu}\,$, (4)
where $u^{\mu}$ is a unit timelike vector field, $g_{\mu\nu}u^{\mu}u^{\nu}=-1$, of the $4$-velocity of the fluid, and where $\rho$ and $P$, respectively, are the energy density and the pressure of the fluid as measured in its rest-frame. Then, from the local law of conservation above, one obtains general relativistic versions of the conservation of mass and the Euler equation.
More generally, when an action $S_{M}$ for a matter field is given, the energy-momentum tensor $T_{\mu\nu}$ is obtained by taking a variation of $S_{M}$ with respect to the metric tensor:
$T_{\mu\nu}:=\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{M}}{\delta g^{\mu\nu}}\,$, (5)
where $g:=\det(g_{\mu\nu})$. For example, the action for a single classical scalar field $\phi$ is given by
$S_{\phi}=\int{\rm d}^{4}x\sqrt{-g}\left[-\dfrac{1}{2}g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi-V(\phi)\right]\,$, (6)
where $V(\phi)$ denotes the potential. Then, the energy-momentum tensor is
$T^{(\phi)}_{\mu\nu}=\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{\phi}}{\delta g^{\mu\nu}}=\nabla_{\mu}\phi\nabla_{\nu}\phi-\dfrac{1}{2}g_{\mu\nu}\left[g^{\alpha\beta}\nabla_{\alpha}\phi\nabla_{\beta}\phi+2V(\phi)\right]\,.$ (7)
2.2 Pointwise energy conditions
Given the energy-momentum tensor for some ordinary matter field, one may expect that the energy density measured by an observer with $4$-velocity $v^{\mu}$ should be positive, namely,
$T_{\mu\nu}v^{\mu}v^{\nu}\geqslant 0\,.$ (8)
For example, this implies $\rho\geqslant 0$ for the perfect fluid in the rest-frame $v^{\mu}=u^{\mu}$. This may be viewed as the simplest form of energy conditions.
The positivity of energy can, in principle, depend on an observer who measures energy and momentum. Besides the positivity, also important is the causality; in relativistic theory, energy and momentum associated with a particle form a single four-vector (four-momentum $p^{\mu}$), which is required to be timelike or null. Depending upon the choice of an observer, one may consider different types of energy conditions on the given energy-momentum tensor, as will be seen below. However, before going into more details of possible energy conditions, we should discuss why we need to impose them.
There are two main reasons for considering energy conditions, both of which are motivated to choose physically reasonable energy-momentum tensors:
  1. 1
    . Defining physically meaningful spacetime metrics.
    The first reason is that if no restrictions are imposed on the energy-momentum tensor (other than it satisfying conservation laws), virtually any metric could be considered a solution to the Einstein equations. This is because for any given metric, one can simply view the Einstein equations as the defining equations for an energy-momentum tensor. However, such an arbitrary definition would not represent a physically meaningful spacetime. Instead, we must first impose realistic conditions on the matter fields. Only then should the Einstein equations be used to determine a physically valid metric.
    In this context, we are particularly concerned with the Einstein tensor, $G_{\mu\nu}$, as it (divided by $8\pi G$) is viewed as the energy-momentum tensor itself. The energy conditions most relevant to this discussion are the Weak Energy Condition (WEC), which requires positivity of energy, and the Dominant Energy Condition (DEC), which requires both positivity and causality.
  2. 2
    . Probing spacetime by geodesics.
    The second reason relates to how we probe spacetime. The behavior of causal geodesic congruences—which represent a collection of observers or probe particles—is often used to understand the structure of spacetime. Energy conditions directly influence how these geodesic congruences behave through the Einstein equations. In particular, we are interested in causal geodesic congruences. As we will see below, energy conditions determine the focusing of nearby geodesics due to gravitational attractive force. Combined with energy conditions, we can formulate focusing theorems for causal geodesics, which can be then used to prove, in an essential way, a number of important theorems in general relativity, such as singularity theorems, black hole mechanics, and topological censorship.
    In this context, we are primarily concerned with the Ricci tensor, $R_{\mu\nu}$, as it directly appears in the Raychaudhuri equation as we will explain below. The energy conditions most relevant here are the Strong Energy Condition (SEC), which relates to timelike convergence, and the Null Energy Condition (NEC), which relates to null convergence.
Now we state the four standard energy conditions, WEC, DEC, SEC, and NEC, locally defined within the classical general relativity. The energy-momentum tensor $T_{\mu\nu}$ itself can be classified essentially into four types based on the eigenvalue problem for the components $T^{(a)(b)}$ with respect to a local Lorentz/orthonormal frame $\{e^{(a)}_{\mu}\}$, $\eta_{ab}e^{(a)}_{\mu}e^{(b)}_{\nu}=g_{\mu\nu}$:
$\det(T^{(a)(b)}-\lambda\eta^{ab})=0\,.$ (9)
The Hawking-Ellis classification HE73 consists of Type I – Type IV. We can characterize each condition, WEC, DEC, SEC, or NEC, in terms of the eigenvalues $\lambda$ for (9). The most generic case is Type I, for which
$T^{(a)(b)}=\rho e^{(a)}_{0}e^{(b)}_{0}+\sum_{i=1}^{3}p_{i}e^{(a)}_{i}e^{(b)}_{i}\,$, (10)
so that the Lorentz invariant eigenvalues of $T^{(a)}{}_{(b)}$ are given by $\{-\rho,p_{1},p_{2},p_{3}\}$ HE73 ; Martin-Moruno:2017exc . Many of classically and semiclassically interesting matter fields are of Type I, encompassing, e.g., a perfect fluid (4) and a massive scalar field (7) with the potential $2V(\phi)=m^{2}\phi^{2}$. Type II – IV involves more complications (see HE73 ; Martin-Moruno:2017exc , and see also Maeda:2018hqu ; Maeda:2022vld for recent elaborations). In the following, for simplicity we characterize each of WEC, DEC, SEC, NEC, only in terms of Type I energy-momentum tensor.
{itembox}
[l]Definition 2.2.1 (The Weak Energy Condition (WEC))
$T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
  • For Type I case $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$.
  • This condition simply implies that the energy density measured by any timelike observer should be non-negative and is generally believed to hold for all physically reasonable classical matter fields.
{itembox}
[l]Definition 2.2.2 (The Dominant Energy Condition (DEC)) For arbitrary future-directed timelike vector $\xi^{\mu}$, $T^{\mu}{}_{\nu}\xi^{\nu}$ becomes a future-directed timelike or null vector.
  • For Type I case $\rho\geqslant|p_{i}|\geqslant 0$.
  • Equivalently, for any future- (or past-) directed timelike vectors $\xi_{1}^{\mu}$ and $\xi_{2}^{\mu}$, $T_{\mu\nu}\xi_{1}^{\mu}\xi_{2}^{\nu}\geqslant 0$.
  • DEC implies WEC.
  • DEC implies that the local energy density is non-negative, and the local energy flow vector is non-spaclike. This ensures the following conservation law:
    {itembox}
    [l]Proposition 2.2.3 (Conservation law) Suppose the energy-momentum tensor for matter fields $T_{\mu\nu}$ satisfies the local conservation $\nabla_{\nu}T^{\mu\nu}=0$ and the DEC. If $T_{\mu\nu}=0$ on a closed, achronal hypersurface $\Sigma$, then $T_{\mu\nu}=0$ everywhere in the domain of dependence $D(\Sigma)$.
    • Lemma. 4.3.1 HE73 .
    • For achronal surface $\Sigma$, and the domain of dependence $D(\Sigma)$, see Sec. 2.5 below.
  • Another important consequence of the DEC is the positivity of the total energy of an isolated system. In general relativity, an isolated system is defined in terms of asymptotic flatness, and the total energy defined at spatial infinity is called the “Arnowitt-Deser-Misner (ADM)” energy Arnowitt:1962hi , whose positivity was shown by Schoen-Yau Schon:1981vd , Witten Witten:1981mf : {itembox}[l]Theorem 2.2.4 (Positive energy theorem) The total energy is positive for initial data on an asymptotically flat Cauchy surface, provided the DEC holds.
    • As the total energy of a given isolated system, the ADM energy is a constant. One can also consider the so-called “Bondi” energy Bondi:1962px , which is defined at null infinity (see Sec. 2.6) and describes the remaining energy at an instant of retarded time after the emission of gravitational radiation. For precise definitions of asymptotic flatness, as well as the ADM and the Bondi energy, see Sec. 11.2 of Wald’s book Wald84 .
    • The positivity of the Bondi energy was shown by Horowitz-Perry Horowitz:1981uw , Schoen-Yau Schon:1982re , Ludvigsen-Vickers Ludvigsen:1982 , again under the DEC. The positivity of these two, in particular, the Bondi energy, ensures the stability of an isolated system in general relativity.
    • This theorem was generalized to include black holes and also for the asymptotically AdS case by Gibbons et al Gibbons:1982jg .
{itembox}
[l]Definition 2.2.5 (The Strong Energy Condition (SEC))
$\left(T_{\mu\nu}-\frac{1}{2}T^{\lambda}{}_{\lambda}g_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\quad\mbox{for arbitrary timelike vector $\xi^{\mu}$.}$
  • For Type I case $\rho+p_{i}\geqslant 0$, $\rho+\sum_{i=1}^{3}p_{i}\geqslant 0$.
  • Similar to WEC, SEC is introduced to study the focusing behavior of causal geodesics. Through the Einstein equations, SEC is equivalent to the condition $R_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$, hence leading to the focusing of timelike geodesic congruences (Theorem 2.4.4 below).
  • For a cosmological constant, the energy-momentum tensor can be written as $8\pi GT^{(\Lambda)}_{\mu\nu}=-\Lambda g_{\mu\nu}$ and hence SEC implies
    ${\Lambda}g_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0\,.$ (11)
    Therefore, if a positive cosmological constant is dominant as a cosmic ingredient, the SEC is not satisfied. Instead of focusing, the timelike geodesic congruence has a positive expansion, expressing that a repulsive force is acting on world-lines of timelike test particles. This repulsive nature, if applying to cosmology, leads to an accelerating expansion of the universe, as necessary to describe an inflationary universe or dark energy dominated universe. For this reason, the interest in SEC lies more in its violation rather than its satisfaction. To be more concrete, let us consider the energy-momentum tensor for a scalar field (7). In the vacuum state (i.e., no excitation of $\phi$, and the derivative terms can be ignored), we virtually obtain $T^{(\phi)}_{\mu\nu}\approx-V(\phi)g_{\mu\nu}$. This is equivalent to the perfect fluid with $\rho=-P=V(\phi)$. Since $T_{\mu\nu}-(1/2)T^{\lambda}{}_{\lambda}g_{\mu\nu}\approx V(\phi)g_{\mu\nu}$, as far as $V(\phi)$ is positive, the SEC is violated.
  • There is no containment relationship between SEC and WEC. For example, for a negative cosmological constant, $\Lambda<0$, SEC is satisfied, but WEC fails to hold.
{itembox}
[l]Definition 2.2.6 (The Null Energy Condition (NEC))
$T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ for arbitrary null vector $k^{\mu}$.
  • For Type I case $\rho+p_{i}\geqslant 0$.
  • Since NEC is equivalent to the condition $R_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ via the Einstein equations, by applying it to the Raychaudhuri equation (see (23) below), one can derive the focusing theorem for a null geodesic congruence (Theorem 2.4.2 below). In this sense, this is also called the Null Convergence Condition (NCC).
  • NEC is equivalent to WEC with a timelike vector $\xi^{\mu}$ replaced by a null vector $k^{\mu}$. If WEC holds, then by continuity, NEC also holds. Similarly, if SEC conditions holds, then for any $\xi^{\mu}$, $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant(1/2)T^{\lambda}{}_{\lambda}\xi_{\mu}\xi^{\nu}$, and hence, by continuity, NEC also holds.
  • The violation of NEC (NCC) implies the violation of WEC. Suppose NEC is violated. Then, for any future-directed null vector $k^{\mu}$, there exists a constant $\delta>0$ such that
    $T_{\mu\nu}k^{\mu}k^{\nu}=-\delta<0\,.$
    Let $u^{\mu}$ be an appropriate future-directed timelike vector. Then, $\xi^{\mu}:=k^{\mu}+\epsilon u^{\mu},\,(\epsilon>0)$ can also be a timelike vector. For this timelike vector $\xi^{\mu}$,
    $T_{\mu\nu}\xi^{\mu}\xi^{\nu}=-\delta+2\epsilon T_{\mu\nu}k^{\mu}u^{\nu}+O(\epsilon^{2})\,.$
    Therefore for sufficiently small $\epsilon$, $T_{\mu\nu}\xi^{\mu}\xi^{\nu}<0$.
We summarize the above four locally defined energy conditions in Table 1.
Table 1: Four locally defined energy conditions. Here $\xi^{\mu}$ and $k^{\mu}$ denote any timelike and null vector, respectively, and “causality restriction” means that superluminal flow of energy is not allowed.
Definition Purpose and effects Type I case
WEC    $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ positivity restriction    $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$
DEC $T_{\mu\nu}\xi_{1}^{\mu}\xi_{2}^{\nu}\geqslant 0$    positivity and causality restriction    $\rho\geqslant|p_{i}|\geqslant 0$
SEC $R_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ timelike focusing   $\rho+\sum_{i}p_{i}\geqslant 0$, $\rho+p_{i}\geqslant 0$
NEC $R_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ null focusing   $\rho+p_{i}\geqslant 0$
2.3 Geodesic congruences on curved spacetime
Given the metric which solves the Einstein equations, we can learn a lot about the properties of spacetime not only by directly looking into the metric tensor itself, but also—sometimes more effectively—by inspecting the behavior of causal geodesic curves, which are world-lines of probe particles such as massive test particles or photons. In this section, we first review timelike and null geodesics, and derive the formulas governing the rate of change of expansion for a family (congruence) of causal geodesics. Then, we review the (classical) focusing theorem.
Let us consider a timelike geodesic curve $\gamma$, whose unit tangent vector is expressed by $u^{\mu}={\rm d}x^{\mu}/{\rm d}\tau$ with $\tau$ being the proper time. The tangent vector satisfies $g_{\mu\nu}u^{\mu}u^{\nu}=-1$ and the geodesic equation,
$u^{\nu}\nabla_{\nu}u^{\mu}=0\,.\quad$ (12)
We consider a geodesic congruence of $\gamma$, which is a bundle of mutually non-intersecting timelike geodesics nearby $\gamma$. Let us define the tensor field,
$h_{\mu\nu}:=g_{\mu\nu}+u_{\mu}u_{\nu}\,.$ (13)
It immediately follows that $h^{\mu}{}_{\sigma}h^{\sigma}{}_{\nu}=h^{\mu}{}_{\nu}$ and $h^{\mu}{}_{\nu}u^{\nu}=0$, and thus the tensor field $h^{\mu}{}_{\nu}$ projects any vectors to the $3$-dimensional subspace orthogonal to $u^{\mu}$. It also satisfies $u^{\sigma}\nabla_{\sigma}h^{\mu}{}_{\nu}=0$. Then, we can decompose the tensor field, $\nabla_{\nu}u_{\mu}$, into its trace-part $\theta$, trace-free symmetric part $\sigma_{\mu\nu}$, and anti-symmetric part $\omega_{\mu\nu}$, as follows:
$\nabla_{\nu}u_{\mu}=\dfrac{1}{3}\theta h_{\mu\nu}+\sigma_{\mu\nu}+\omega_{\mu\nu}\,$,
$\theta:=h^{\mu\nu}\nabla_{\nu}u_{\mu}=\nabla_{\mu}u^{\mu}\,,\quad\sigma_{\mu\nu}:=\nabla_{(\nu}u_{\mu)}-\dfrac{1}{3}\theta h_{\mu\nu}\,,\quad\omega_{\mu\nu}:=\nabla_{[\nu}u_{\mu]}\,$, (14)
where $\theta$ is called the expansion of the congruence of $\gamma$, $\sigma_{\mu\nu}$ the shear, and $\omega_{\mu\nu}$ the rotation. Note that by definition $u^{\mu}\sigma_{\mu\nu}=0=u^{\mu}\omega_{\mu\nu}$. The geometric meaning of these quantities can be read off from Figure 1.
Figure 1: The geometric interpretation of the expansion $\theta$, the shear-tensor $\sigma_{\mu\nu}$, and the rotation $\omega_{\mu\nu}$.
Most important in the context below is the behavior of the expansion $\theta$, since it expresses the focusing of geodesics due to gravitational attraction. We are often concerned with hypersurface orthogonal geodesic congruences, for which $\omega_{\mu\nu}$ must vanish according to the Frobenius theorem [^5] If a vector field $n^{\mu}(\neq 0)$ satisfies $n_{[\alpha}\nabla_{\beta}n_{\gamma]}=0\,$, then the orbits of $n^{\mu}$ are orthogonal to some hypersurface. Conversely, if $n^{\mu}$ is hypersurface orthogonal, then it satisfies the above equation. .
Let us consider a congruence of timelike geodesic $\gamma$ orthogonal to $3$-dimensional hypersurface $\Sigma$ with vanishing rotation $\omega_{\mu\nu}=0$. If we denote by $A$ an infinitesimal cross-sectional ($3$-dimensional) area element of the congruence, then
$\theta=\dfrac{1}{A}\dfrac{{\rm d}A}{{\rm d}\tau}\,$, (15)
and thus describes the local rate of change of cross-sectional area along the geodesic $\gamma$. Now, by taking the derivative of $\nabla_{\nu}u_{\mu}$ along $\gamma$, i.e., $u^{\alpha}\nabla_{\alpha}(\nabla_{\nu}u_{\mu})$, and taking its trace, we obtain the Raychaudhuri equation
$\dfrac{{\rm d}\theta}{{\rm d}\tau}=-\dfrac{1}{3}\theta^{2}-\sigma_{\mu\nu}\sigma^{\mu\nu}-R_{\mu\nu}u^{\mu}u^{\nu}\,$, (16)
where $R_{\mu\nu}$ is the Ricci tensor. We can also derive the equations for $\sigma_{\mu\nu}$ by taking the symmetric trace-free part of $u^{\alpha}\nabla_{\alpha}(\nabla_{\nu}u_{\mu})$,
$u^{\alpha}\nabla_{\alpha}\sigma_{\mu\nu}$ $=$ $-\frac{2}{3}\theta\sigma_{\mu\nu}-\left(\sigma_{\mu\sigma}\sigma_{\nu}{}^{\sigma}-\frac{1}{3}\sigma_{\alpha\beta}\sigma^{\alpha\beta}h_{\mu\nu}\right)$ (17)
$-C_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}+\frac{1}{2}\left(h_{\mu}{}^{\alpha}h_{\nu}{}^{\beta}-\frac{1}{3}h^{\alpha\beta}h_{\mu\nu}\right)R_{\alpha\beta}\,$,
where
$C_{\mu\alpha\nu\beta}:=R_{\mu\alpha\nu\beta}+g_{\mu[\beta}R_{\nu]\alpha}+g_{\alpha[\nu}R_{\beta]\mu}+(1/3)Rg_{\mu[\nu}g_{\beta]\alpha}$ (18)
is the Weyl tensor. Note that we omit terms related to $\omega_{\mu\nu}$ since $u^{\mu}$ is assumed to be hypersurface orthogonal.
We can derive similar formulas for null geodesic congruences. Again, we focus on hypersurface orthogonal null geodesics, since a number of interesting spacetime structures can be prescribed by a null hypersurface, generated by null geodesics. A subtle point is that null geodesic generators of a null hypersurface $\cal N$} are tangent to, and simultaneously, normal to $\cal N$}. For this reason, even though $\cal N$} itself is a $3$-dimensional hypersurface, a spacelike cross-section $S$ of $\cal N$}, to which the null congruence is orthogonal, is $2$-dimensional. Let $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ be an affinely parametrized null geodesic vector field, so that
$k^{\nu}\nabla_{\nu}k^{\mu}=0\,,\quad k^{\mu}k_{\mu}=0\,.$ (19)
We also assume that $k^{\mu}$ is null-hypersurface orthogonal. Let $A$ be an infinitesimal cross-sectional ($2$-dimensional) area element of the congruence. Then, the expansion of the congruence is defined by
$\theta=\nabla_{\mu}k^{\mu}=\dfrac{1}{A}\dfrac{{\rm d}A}{{\rm d}\lambda}\,.$ (20)
We can also define the shear $\sigma_{\mu\nu}$ in a manner similar to that or the timelike geodesic case. For this purpose, we introduce an auxiliary null vector field $l^{\mu}$, parallelly transported by $k^{\mu}$, so that it satisfies, $l^{\mu}l_{\mu}=0$, $l^{\mu}k_{\mu}=-1$, and $k^{\nu}\nabla_{\nu}l^{\mu}=0$. Then, we define the projection tensor $p^{\mu}{}_{\nu}$ by
$p_{\mu\nu}:=g_{\mu\nu}+k_{\mu}l_{\nu}+l_{\mu}k_{\nu}\,.$ (21)
It satisfies $p^{\mu}{}_{\sigma}p^{\sigma}{}_{\nu}=p^{\mu}{}_{\nu},\,p^{\mu}{}_{\nu}k^{\nu}=0=p^{\mu}{}_{\nu}l^{\nu},\,k^{\sigma}\nabla_{\sigma}p^{\mu}{}_{\nu}=0$. Note that the auxiliary null vector field $l^{\mu}$ cannot be uniquely chosen, and accordingly $p^{\mu}{}_{\nu}$ is not uniquely defined. Then, the shear and the rotation, are defined by
$\sigma_{\mu\nu}=\nabla_{(\mu}k_{\nu)}-\dfrac{1}{2}\theta p_{\mu\nu}\,,\quad\omega_{\mu\nu}=\nabla_{[\mu}k_{\nu]}\,.$ (22)
Again, since we are concerned with surface orthogonal null geodesics, in what follows we set $\omega_{\mu\nu}=0$. Then, we can derive the corresponding Raychaudhuri equation
$\dfrac{{\rm d}\theta}{{\rm d}\lambda}=-\dfrac{1}{2}\theta^{2}-\sigma_{\mu\nu}\sigma^{\mu\nu}-R_{\mu\nu}k^{\mu}k^{\nu}\,.$ (23)
We can also derive the equation for the shear tensor as
$k^{\alpha}\nabla_{\alpha}\sigma_{\mu\nu}=-\theta\sigma_{\mu\nu}-\left(\sigma_{\mu\sigma}\sigma_{\nu}{}^{\sigma}-\frac{1}{2}\sigma_{\alpha\beta}\sigma^{\alpha\beta}p_{\mu\nu}\right)-p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\,$, (24)
where $C^{\mu}{}_{\alpha\nu\beta}$ is the Weyl tensor.
2.4 Focusing theorems and energy conditions
Now that we have formulas of the expansion and shear for both timelike and null geodesic congruence, we discuss the focusing effects. Below we consider first the null geodesic case in detail, and then derive essentially the same conclusion for the timelike case. We immediately see that the first and second terms in the right-hand side of (23) are non-positive definite, hence always tend to decrease the expansion $\theta$ and cause the convergence of the null geodesic congruence. The third term, $R_{\mu\nu}k^{\mu}k^{\nu}$, is non-positive if the NEC holds, thus also tends to decrease $\theta$ under the NEC. Consider the case that both the expansion and shear tensors vanish initially. If $R_{\mu\nu}k^{\mu}k^{\nu}\neq 0$ at some point of $\gamma$, then via the Raychaudhuri equation, $\theta$ becomes non-vanishing, which, again via the Raychaudhuri equation, makes $\theta$ start to decrease. Suppose now $R_{\mu\nu}k^{\mu}k^{\nu}=0$ (vacuum spacetime), and initially $\theta=0=\sigma_{\mu\nu}$. Even in this case, if the third term of equation (24) is non-vanishing, $p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0$, it makes the shear tensor non-vanishing by (24) and then produces, via (23), non-vanishing expansion. Then, again from (23), $\theta$ starts to decrease. Now we state the following condition. {itembox}[l]Definition 2.4.1 (Null generic condition) A spacetime is said to satisfy the null generic condition if every null geodesic (whose tangent vector field is denoted by $k^{\mu}$) possesses at least one point where
$k_{[\alpha}R_{\mu]\beta\nu[\gamma}k_{\sigma]}k^{\beta}k^{\nu}\neq 0$.
This condition is equivalent to
$\mbox{either (i) }\,R_{\mu\nu}k^{\mu}k^{\nu}\neq 0\,,\,\,\,\mbox{or (ii) }\,p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0\,$, (25)
at at least a single point on any null geodesic $\gamma$ and is interpreted as a non-vanishing tidal force acting on the photon described by $\gamma$.
Suppose that our null geodesic curve $\gamma$ with the tangent vector $k^{\mu}$ is complete with respect to the affain parameter $\lambda$; namely $\gamma$ can be extended to take arbitrary value of $\lambda$ on the spacetime considered. Then, from the above observation, it is inevitable that under the generic conditions and the NEC, the expansion eventually negatively diverges $\theta\rightarrow$-$\infty$. In fact, under the NEC, we have from the Raychaudhuri equation (23),
$\frac{{\rm d}\theta}{{\rm d}\lambda}\leqslant-\frac{1}{2}\theta^{2}<0\,.$ (26)
Assuming the initial value $\theta_{0}$ at $\lambda=0$, integrating the above inequality, we obtain
$\frac{1}{\theta}\geqslant\frac{1}{\theta_{0}}+\frac{\lambda}{2}\,.$ (27)
In particular, if the initial value of the expansion takes some negative value $\theta_{0}<0$, then
$\theta\leqslant-\frac{2}{2/|\theta_{0}|-\lambda}\,.$ (28)
Thus, we arrive at the following result (see also Figure 2):
{itembox}
[l]Theorem 2.4.2 (Null focusing) Under the NEC, any hypersurface-orthogonal null geodesic congruence is focusing:
$\dfrac{{\rm d}\theta}{{\rm d}\lambda}\leqslant 0$.
Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\lambda=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the affine length
$\lambda\leqslant\dfrac{2}{|\theta_{0}|}$.
Figure 2: A null geodesic congruence whose expansion takes initially negative value $\theta_{0}<0$ converges, and within the affine length $\lambda\leqslant 2/|\theta_{0}|$, the cross-section area $A$ of the congruence vanishes, where by definition $\theta\rightarrow$-$\infty$.
Suppose the above null geodesic congruence containing $\gamma$ is emanating from a point $p\in\gamma$ at some $\lambda<0$, attains the negative value $\theta_{0}<0$ at $\lambda=0$, and as claimed above, reaches a point $q\in\gamma$ where $\theta\rightarrow$-$\infty$. Then, in general, $q$ is said to be conjugate to $p$. Then, one can claim that under the null generic condition and NEC, $\gamma$ must possess a pair of conjugate points (see, proposition 9.3.7 Wald). We can also consider a congruence of null geodesics with tangent vector field $k^{\mu}$ which are orthogonal to a $2$-dimensional spacelike surface $S$. There exist, in general, two families of such surface orthogonal null geodesics (often referred to as “ingoing” and “outgoing” null geodesics). In this case, a point $q$ away from $S$ along $\gamma$ where $\theta\rightarrow$-$\infty$ is said to be conjugate to $S$. The occurrence of a conjugate point $q\in\gamma$ to $p\in\gamma$ (or to $S$) is a natural consequence under the attractive nature of gravitation. An important and useful aspect of the occurrence of a conjugate point $q$ is that a null geodesic $\gamma$ fails to remain on the boundary of causal future (or past) of $p$ (or $S$). We shall discuss this aspect in more detail in the next subsection.
For a complete timelike geodesic congruence, whose tangent vector field is $u^{\mu}={\rm d}x^{\mu}/{\rm d}\tau$ with $\tau$ being the proper time, we can obtain (in more straightforward manner) results similar to the null geodesic focusing as summarized below. Since for a timelike geodesics, the Raychaudhuri equation contains the term $R_{\mu\nu}u^{\mu}u^{\nu}$, we use, instead of NEC, the SEC as it implies, $R_{\mu\nu}u^{\mu}u^{\nu}=8\pi G(T_{\mu\nu}-(1/2)Tg_{\mu\nu})u^{\mu}u^{\nu}\geqslant 0$ via the Einstein equations.
{itembox}
[l]Definition 2.4.3 (Timelike generic condition) A spacetime is said to satisfy the timeilke generic condition if every timelike geodesic (whose tangent vector field is denoted by $u^{\mu}$) possesses at least one point where
$R_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}\neq 0$.
Consider two nearby freely falling particles with the unit $4$-velocity vector fields $u^{\mu}=(\partial/\partial\tau)^{\mu}$, and let $\eta^{\mu}$ be the deviation vector (Jacobi field) between the two. Then the geodesic deviation equations are
$u^{\alpha}\nabla_{\alpha}(u^{\beta}\nabla_{\beta}\eta^{\mu})=-(R^{\mu}{}_{\alpha\nu\beta}u^{\alpha}u^{\beta})\eta^{\nu}\,.$ (29)
Thus, the above generic condition implies that a non-vanishing tidal force acts on every freely falling test particle.
{itembox}
[l]Theorem 2.4.4 (Timelike focusing) Under the SEC, any hypersurface-orthogonal timelike geodesic congruence is focusing:
$\dfrac{{\rm d}\theta}{{\rm d}\tau}\leqslant 0$.
Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\tau=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the proper time
$\tau\leqslant\dfrac{3}{|\theta_{0}|}$.
Similar to the null geodesic case, we can say that a point $q\in\gamma$ is conjugate to a point $p$ (or to $3$-dimensional hypersurface $\Sigma$) when the congruence is emanating from the point $p\in\gamma$ (or orthogonal to $\Sigma$). We can also claim that under the timeilke generic condition and the SEC, $\gamma$ must possess a pair of conjugate points. (See, proposition 9.3.1 Wald). An important aspect of the occurrence of a conjugate point $q$ for timelike geodesics is that a timelike geodesic $\gamma$ fails to be a local maximum of proper time between $q$ and $p$ (or $\Sigma$).
2.5 Singularity theorems and energy conditions
One of the most important applications of the energy conditions in general relativity is the singularity theorems. In this section, we will discuss how the energy conditions are used in the proof of the theorems.
A spacetime singularity represents some extreme situation where general relativistic description of spacetime as a manifold breaks down due to infinite energy density or infinite spacetime curvature. This, in turn, implies that spacetime singularity should not be considered as part of the spacetime manifold. This leads to the idea that if a freely falling observer reaches a singularity, her/his world-line suddenly ends at a finite proper time and cannot be extended any further. Such a world-line (i.e., endless but never attains arbitrary large values in its proper time) can be described as an inextendible, incomplete timelike geodesic curve. The singularity theorems employ, as the criterion for the presence of a singularity, the existence of an inextendible incomplete causal (timelike or null) geodesic curve. For further motivation, satisfactory and unsatisfactory aspects of this definition of spacetime singularities, see, e.g., HE73 ; Wald84 .
[…]
{itembox}
[l]Proposition 2.5.1 Any achronal null geodesic (hence, any achronal null hypersurface) does not admit a pair of conjugate points on it.
Figure 3 illustrates what happens when a conjugate point appears.
Figure 3: Null geodesics orthogonally emanating from $S$. One such null geodesic $\gamma$ is passing through $p$, where its expansion changes sign from positive to negative, and then has a conjugate point $r$ in $J^{+}(p)$. Beyond this conjugate point, $\gamma$ fails to remain on the boundary $\partial J^{+}(S)$ and enters the interior $I^{+}(S)$. Since any point $q$ on $\gamma$ within $I^{+}(S)$ can be connected to $S$ by a timelike curve, $\gamma$ no longer is achronal beyond the conjugate point $r$. See also Figure 4 below.
An important example of achronal set is given by, {itembox}[l]Proposition 2.5.2 The boundary $\partial J^{+}(S)$ of the causal future of $S$ is a closed, $3$-dimensional, achronal submanifold.
Let $\Sigma$ be any closed achronal set. The future domain of dependence of $\Sigma$, denoted $D^{+}(\Sigma)$, is defined by the set of all points $p\in M$ such that every past-directed inextendible causal curve from $p$ intersects $\Sigma$. The past domain of dependence of $\Sigma$, denoted $D^{-}(\Sigma)$ is defined analogously. One often considers the union $D(\Sigma):=D^{+}(\Sigma)\cup D^{-}(\Sigma)$. In general, when $D(\Sigma)$ is a proper subset of $M$, i.e., $D(\Sigma)\varsubsetneqq M$, the boundary $\partial D^{\pm}(\Sigma)$ is called the future (past) Cauchy horizon of $\Sigma$, denoted by $H^{\pm}(\Sigma)$, respectively. A physical importance of the notion of the domain of dependence is that the physics inside $D(\Sigma)$ can be determined by the initial data on $\Sigma$, without any other information, e.g., boundary conditions outside $D(\Sigma)$.
If $D(\Sigma)=M$ (i.e., $H^{\pm}(\Sigma)=\emptyset$), then $\Sigma$ is called a Cauchy surface of $M$. If $M$ admits a Cauchy surface, then $M$ is said to be globally hyperbolic. An important role of the global hyperbolicity in the context of the singularity theorems is concerning the length of causal curves. For a smooth causal curve $\lambda$ between two points $p,q\in M$ with tangent $t^{\mu}=(\partial/\partial t)^{\mu}$, the length of $\lambda$ is defined as HE73 ; Wald84 ,
$L[\lambda]=\int_{p}^{q}{\rm d}t\,(-g_{\mu\nu}t^{\mu}t^{\nu})^{1/2}\,.$ (30)
Then, one has {itembox}[l]Proposition 2.5.3 (Existence of maximum length curve) Let $N$ be globally hyperbolic (possibly as a subset of $M$). Then, for any two points $p,q\in N$ with $q\in J^{+}(p)$, there exists a causal geodesic curve $\mu$ from $p$ to $q$ which attains the maximum length among all causal curves from $p$ to $q$.
  • Proposition 6.7.1 HE73 .
This is a consequence of the following three facts: firstly, the set $C(p,q)$ of all causal curves which connect two points $p,q$ in globally hyperbolic region $N$ is compact, secondly, the length $L[\lambda]$ is an upper semicontinuous function on $C(p,q)$, and thirdly, any upper semicontinuous function on a compact space has a maximum value. $\Box$
It is rather easy to see what happens if the assumption of global hyperbolicity is eliminated. For example, consider timelikely separated two points $p,q$ in Minkowski spacetime with a single point $r$ just on a straight line segment connecting the two points removed. In this case, the maximum value (in the sense of $d(p,q)$ below) is the length of the segment but there does not exist a causal curve attaining the maximum value. If we consider timelikely separated two points in anti-de Sitter (AdS) spacetime, and in particular, if these two are not lying in a globally hyperbolic subregion, then the maximum value itself does not exist Ishibashi:2012xk .
As mentioned briefly just below the focusing theorems in the previous subsection, if a timelike geodesic connecting two points $p,q\in M$ contains a pair of conjugate points between the two, then there exists another causal curve whose length (proper time) is larger than the original timelike geodesic. The Lorentzian distance function $d:M\times M\rightarrow{\mathbb{R}}\cup\infty$ between two points $p,q\in M$ is defined by $d(p,q):=\sup\{L(\lambda):\lambda\in C(p,q)\}$. If $p,q\in I^{+}(p),\>p\neq q$, then $d(p,q)>0$, and if $q\in E^{+}(p)$ or not causally related, then $d(p,q)=0$. For any $q,r\in I^{+}(p),\>q\in I^{+}(r)$,
$d(p,q)\geqslant d(p,r)+d(r,q)\,.$ (31)
This is the reverse triangle inequality in Lorentzian distance, which is nothing but the inequality one may see in the so-called “twin paradox.” As can be directly read off from Figure 4, if a timelike geodesic $\gamma$ connecting points $p$ and $q$ contains a point $r$ conjugate to $p$, then there is another timelike geodesic $\gamma^{\prime}$ that also connects $p$ and $q$ with a longer length (proper time), according to (31).
Figure 4: (a) Reverse triangle inequality. (b) A timelike geodesic $\gamma$ from $p$ to $q$ admits a point $r$ conjugate to $p$ between $p$ and $q$. (c) A timelike geodesic $\gamma^{\prime}$ from $p$ to $q$, constructed by deforming $\gamma$ so that $\gamma^{\prime}$ has a larger value of the proper time than $\gamma$. Compare Figure 3.
Similarly, if a null geodesic connecting two points $p,q\in M$ contains a pair of conjugate points on it, then it cannot be achronal since $p,q$ can be then connected by some timelike curve. In other words, it fails to be the fastest causal curve. Combining this observation with Prop. 2.5.3, we have the following:
{itembox}
[l]Proposition 2.5.4 Any causal geodesic curve which connects two points $p$ and $q$, and which attains the maximum length between them cannot have a pair of conjugate points between $p$ and $q$.
The singularity theorems predict the existence of an inextendible, incomplete causal (either timelike or null) geodesic curve, under the following three conditions concerning:
  • (i)
    causal structure,
  • (ii)
    strong gravity,
  • (iii)
    geodesic focusing.
By strengthening or weakening these three conditions, one can formulate different versions of the singularity theorems, but the proof in all cases essentially proceeds to derive a contradiction, under the assumption that every causal geodesic be complete.
Let us discuss the implications of the above three conditions (i)–(iii) in more detail regarding singularity formation.
  • (i)
    Causal structure.
    The first singularity theorem by Penrose Penrose:1964wq assumes the global hyperbolicity, or equivalently the existence of a Cauchy surface, as a condition for global causal structure (Penrose’s theorem actually requires a non-compact Cauchy surface). However, the global hyperbolicity may be considered to be too strong a causality condition. After demonstrating contradictory consequences, instead of claiming that the assumption of the geodesic completeness fails, one may rather be able to conclude that the assumption of the global hyperbolicity must fail. In this respect, it is worth noting that under the same conditions of Penrose’s theorem, except the assumption of the global hyperbolicity, one can construct geodesically complete spacetimes, among which most well-known example is the Bardeen’s regular black hole.
    For this reason, after Penrose’s theorem Hawking:1970zqf , Hawking and Penrose proposed the improved version of the singularity theorem, in which the unwanted assumption of the global hyperbolicity is dropped, and instead, the chronology condition—which demands that no closed timelike curves exist—is imposed. This is a reasonable condition as to consider a physically realistic universe.
  • (ii)
    Strong gravity.
    This property is described by the requirement of the trapped set defined as follows: {itembox}[l]Definition 2.5.5 (Future trapped set) A closed achronal subset $S$ is said to be a future trapped set if $E^{+}(S)$ is compact. The past trapped set is defined analogously. A simple example is a $2$-dimensional closed surface $S$ in the static Einstein universe $M={\mathbb{R}}\times S^{3}$ with the metric,
    ${\rm d}s^{2}=-{\rm d}T^{2}+{\rm d}R^{2}+\sin^{2}R\,({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (32)
    as depicted in Figure 5. In the cosmological context, a point $p$ whose past light cone starts converging again toward the early universe is a past trapped set (see Figure 50 in HE73 ).
    Figure 5: An achronal subset $S$ in spatially compact universe with $\Sigma$ being a compact Cauchy surface. The null geodesics (thick curves) generating the future horismos $E^{+}(S)$ have the two future endpoints due to the compactness of the universe, and accordingly $E^{+}(S)$ also is compact, thus $S$ is a trapped set. One may think of $S$ as any $2$-sphere of constant $T$ and $R$ in the static Einstein universe. (The two points labeled as $S$ in the figure represent the single $2$-sphere $S$.) $E^{+}(S)$ has the two future endpoints; one for the “ingoing” null geodesics, and the other for the “outgoing” ones, which are orthogonal to $S$. It is important to note, however, that in a closed universe, one cannot generally distinguish between ingoing and outgoing.
    In the context of gravitational collapse and black hole physics, the most important example of a trapped set is the closed trapped surface, which is a compact $2$-dimensional surface $\cal T$} such that the expansions $\theta_{\pm}$ of the two families of future (or past) null geodesic congruence orthogonal to $\cal T$} both are negative. An example is any $2$-sphere of constant area radius inside the Schwarzschild (horizon) radius $r_{h}:=2G{\cal M}$. To be concrete, let us express the Schwarzschild metric666 In the standard coordinates $x^{\mu}=(t,r,\vartheta,\varphi)$, the Schwarzschild metric is given as ${\rm d}s^{2}=-f(r)\,{\rm d}t^{2}+f^{-1}(r)\,{\rm d}r^{2}+r^{2}({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (33) where $f(r):=1-2G{\cal M}/r$, ${\cal M}=const.$. In one of the exterior region, U<0,V>0U<0,\>V>0, the two coordinates are related as $U=-e^{-(t-r_{*})/2r_{h}},\>V=e^{(t+r_{*})/2r_{h}}$, where $r_{*}=r+r_{h}\log(r-r_{h})$. in the Kruskal coordinates $\bar{x}^{\mu}=(U,V,\vartheta,\varphi)$,
    ${\rm d}s^{2}=-\dfrac{4r_{h}^{3}e^{-r/r_{h}}}{r}{\rm d}U{\rm d}V+r^{2}({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (34)
    where the area radius $r$ is now regarded as the function of $U$ and $V$. This coordinate system covers the entire (maximally extended) Schwarzschild spacetime (see, e.g., Sec. 6.4 of Wald’s book Wald84 ). In this Kruskal coordinate system, the center $r=0$ of the spherical symmetry is at $UV=1$, the horizon (i.e., $r=r_{h}$) is at $UV=0$, the interior region ($0<r<r_{h}$) consists of the black hole region U>0,V>0U>0,\;V>0 and the white hole region U<0,V<0U<0,\;V<0, while the exterior region ($r>r_{h}$) consists of two causally disjoint regions: U<0,V>0U<0,\>V>0 and U>0,V<0U>0,\>V<0 (see Figure 6). The expansions $\theta_{\pm}$ of the two families (out-going and in-going) of future directed null geodesic congruences orthogonal to any $2$-sphere of constant $U$ and $V$ are calculated as
    $\theta_{+}=-\dfrac{U}{r_{h}r}\,,\quad\theta_{-}=-\dfrac{V}{r_{h}r}\,.$ (35)
    Figure 6: The maximally extended Schwarzschild spacetime in Kruskal coordinates $(U,V)$. The future event horizon is at $U=0,V>0$ (and $V=0,U>0$), the $2$-sphere $\sigma_{B}$ at $U=0=V$ is the bifurcate surface. Any $2$-sphere $\cal T$} inside the black hole region, $U>0,V>0$, is a future trapped surface, and any $2$-sphere ${\cal T}_{0}$ on the future event horizon $U=0,V>0$ is a MOTS.
    Inside the black hole region $U>0,V>0$, both $\theta_{\pm}<0$ on $\cal T$}, and therefore $\cal T$} is indeed a closed trapped surface. If every null geodesic were to be complete, then it would follow from the null focusing theorem 2.4.2 (with the NEC obviously holding in vacuum spacetime) that every future-directed null geodesic orthogonally emanating from $\cal T$} must have an endpoint (conjugate point to $\cal T$}), making $E^{+}({\cal T})$ compact, hence $\cal T$} is a trapped set. In the present case, the actual “endpoints” of $E^{+}({\cal T})$, where $\theta_{\pm}\rightarrow$-$\infty$, are located at “r=0r=0” as explicit in (35), and thus not part of the spacetime manifold but correspond to the central singularity. Therefore null geodesics along $E^{+}({\cal T})$ are incomplete. Note also that for any $2$-sphere ${\cal T}_{0}$ on the future event horizon, say $U=0,V>0$, for the in-going null $\theta_{-}<0$, while for the out-going null $\theta_{+}=0$ and its outward derivative $\partial_{r}\theta_{+}\geqslant 0$. Such a closed surface ${\cal T}_{0}$ characterizing locally the outer boundary of trapped region is called the stably marginally outer trapped surface (MOTS) Andersson:2005gq (see, Sec. 2.6.2 below) or more loosely the apparent horizon.
  • (iii)
    Geodesic focusing.
    The null and timelike focusing theorems (Theorems 2.4.2 and 2.4.4) are a direct result of the NEC and SEC, and the Raychaudhuri equations as we have seen in the previous section. As discussed before, the null and timelike focusing theorems ensures that any complete causal geodesics inevitably contain a pair of conjugate points.
Let us state Penrose’s version of the singularity theorem Penrose:1964wq . {itembox}[l]Theorem 2.5.6 (Penrose 1965) Spacetime $(M,g_{\mu\nu})$ cannot be null geodesically complete if:
  • (i)
    there is non-compact Cauchy surface $\Sigma$ in $M$,
  • (ii)
    there is a closed trapped surface $\cal T$} in $M$,
  • (iii)
    the NEC holds.
The proof goes roughly as follows. Suppose $M$ is null geodesically complete. From (ii) and (iii) (thus null focusing theorem), it follows that $E^{+}({\cal T})$ is compact. Furthermore, since $M$ is globally hyperbolic by assumption (i), $E^{+}({\cal T})=\partial J^{+}({\cal T})$. It immediately follows from Prop. 2.5.2 that $E^{+}({\cal T})$ is a compact $3$-dimensional achronal, topological manifold. Again, from the global hyperbolicity of $M$ and the achronality of $E^{+}({\cal T})$, one can find a continuous one-to-one map $\psi$ from $E^{+}({\cal T})$ to $\Sigma$. Then, its image $\psi(E^{+}({\cal T}))$ must have a boundary in $\Sigma$ since $E^{+}({\cal T})$ is compact, whereas $\Sigma$ is non-compact. However, this contradicts the fact that $E^{+}({\cal T})$ is a compact $3$-dimensional achronal topological manifold without boundary. $\Box$
Next, consider Hawking and Penrose’s improved version of the theorem Hawking:1970zqf , in which the assumption of the global hyperbolicity (i.e., the existence of a Cauchy surface) is eliminated: {itembox}[l]Theorem 2.5.7 (Hawking-Penrose 1970) The following three conditions cannot all hold:
  • (i)
    the chronology condition holds,
  • (ii)
    there is a trapped set $\cal S$},
  • (iii)
    every inextendible causal geodesic contains a pair of conjugate points.
  • This is an alternative expression HE73 of the Hawking-Penrose theorem Hawking:1970zqf .
  • Condition (iii) results from the NEC and the SEC, the generic conditions 2.4.1 and 2.4.3, the focusing theorems 2.4.2 and 2.4.4, and the hypothetical assumption of causal geodesic completeness.
The proof goes roughly as follows. If the above three conditions all hold, one can construct a globally hyperbolic subregion $N$ which contains inextendible timelike curves [^7] One can always find a future inextendible timelike curve $\gamma$ in $D^{+}(E^{+}({\cal S}))$, since $E^{+}({\cal S})$ is compact as $\cal S$} is a trapped set, whereas $H^{+}:=\partial D^{+}(E^{+}({\cal S}))$ is non-compact under (i) and (iii) (Lemma 8.2.1 HE73 ). This is essentially because if $H^{+}$ were to be compact, it would have to contain inextendible null geodesic generators, which however must contain a pair of conjugate points due to condition (iii), thus contradicting the achronality of $H^{+}$ (Prop. 2.5.1). Consider now the compact set $F:=E^{+}({\cal S})\cap\overline{J^{-}(\gamma)}$. Since $F$ turns out to be past trapped, $E^{-}(F)$ is compact. Hence, in a similar manner to $\gamma$, one can show that $D^{-}(E^{-}(F))$ contains a past inextendible timelike curve $\lambda$. Then, one can construct the desired globally hyperbolic subregion as $N:=D(E^{-}(F))$, which, by construction, contains both future and past inextendible timelike curves (constructed by suitably deforming and connecting $\gamma$ and $\lambda$ if necessary). . Among them, there must exist an inextendible timelike geodesic curve $\mu$ which attains the maximum length between any pair of two points on it, according to Prop. 2.5.3. Then, as such, $\mu$ cannot admit a pair of conjugate points on it according to Prop. 2.5.4. However, this contradicts (iii). $\Box$
2.6 Black holes and energy conditions
A black hole obeys the laws of mechanics, which correspond precisely to the ordinary laws of thermodynamics. This correspondence was first established within classical general relativity as a mathematical resemblance Bardeen:1973gs , and was soon promoted to a perfect correspondence by the discovery of Hawking radiation Hawking:1975vcx . This implies that a stationary black hole can be completely characterized by only a few parameters, just like a thermodynamic equilibrium system can be described by a few macroscopic variables. This property is established as the black hole uniqueness theorem, which states that any asymptotically flat, stationary rotating black hole that is regular everywhere on and outside the event horizon, and that solves the electrovacuum Einstein equations, must be described by the Kerr-Newmann metric. In this section, we see the key roles the energy conditions (WEC, DEC, SEC, and NEC) play in the proof of the classical black hole mechanics.
[…]
$B=M\setminus J^{-}(\mathscr{I}^{+})\,.$ (36)
[…]
2.6.1 Black hole mechanics and energy conditions
[…]
{itembox}
[l]Proposition 2.6.1 If the NEC holds, then a closed trapped surface $\cal T$}, if it exists, cannot intersect $J^{-}(\mathscr{I}^{+})$.
  • Propositoin 9.2.1 HE73 .
  • This implies that a closed trapped surface $\cal T$} cannot be seen from an ideal distant observer $\mathscr{I}^{+}$, and therefore indicates that $\cal T$} should be hidden inside the black hole region $B$.
  • Under the DEC, Schoen-Yau (1983) Schoen:1983tiu showed that if a sufficiently large amount of matter is condensed in a small enough region, a closed trapped surface $\cal T$} must occur. This, combined together with the singularity theorems and the above proposition, indicates that complete gravitational collapse results in the formation of a black hole and a singularity inside.
The proof goes roughly as follows. To derive a contradiction, suppose ${\cal T}\in J^{-}(\mathscr{I}^{+})$. Then, a null geodesic generator $\mu$ of $\partial J^{+}({\cal T})$ has an intersection with $\mathscr{I}^{+}$ in $\tilde{M}$. This implies on one hand that $\mu$ is future complete, and on the other hand that $\mu$ has the past endpoint at $\cal T$}. Then, a null geodesic congruence of $\mu$ orthogonal to $\cal T$} must have a negative expansion $\theta_{+}<0$ at $\cal T$}, since $\cal T$} is a closed trapped surface. Then, by Theorem 2.4.2, $\mu$ must have a conjugate point to $\cal T$} in the future of $\cal T$}. This, however, contradicts the fact that $\mu$ is the generator of the achronal hypersurface $\partial J^{+}({\cal T})$ (Proposition 2.5.1). $\Box$
Applying a similar type of argument to the event horizon itself, one obtains the following result:
{itembox}
[l]Theorem 2.6.2 (Area theorem: The second law) Under the NEC, the cross-sectional area of the event horizon never decreases toward future.
  • Propositoin 9.2.1 HE73 , Theorem 12.2.6 Wald84 .
  • One-quarter of the horizon cross-sectional area can be interpreted as the entropy of the black hole Bekenstein:1973ur (see Sec 4 and Formula 4.3.1), and accordingly the above property of the horizon area increase is phsyically interpreted as the second law of the black hole thermodynamics.
[…]
Next we consider stationary black holes, since at a sufficiently late time, any black hole formed by gravitational collapse is expected to settle down to a stationary black hole spacetime, which corresponds to an equilibrium thermodynamic system.
An asymptotically flat spacetime is said to be stationary, if there is a Killing vector field, $t^{\mu}$ whose orbits are complete (i.e., $t^{\mu}$ generates an isometry group) and timelike, at least near infinity $\mathscr{I}$. Similarly, a spacetime is said to be axisymmetric, if there exists a Killing vector field $\varphi^{\mu}$ whose orbits are closed, spacelike curves. We also assume that the two Killing vector fields $t^{\mu}$ and $\varphi^{\mu}$ commute. An important consequence is that: {itembox}[l]Theorem 2.6.3 (Rigidity theorem) If an asymptotically flat, analytic, stationary spacetime solving the electrovacuum Einstein equations contains a black hole, then the event horizon of the black hole must be a Killing horizon. Furthermore, if a stationary black hole is rotating, it must also be axisymmetric.
  • Proposition 9.3.6 HE73 . See also Hollands:2006rj ; Hollands:2008wn for more general analyses.
  • A null hypersurface $\cal N$} is said to be a Killing horizon, if there exists a Killing vector field $\chi^{\mu}$ which is normal to $\cal N$} on $\cal N$}.
[…]
  • For an asymptotically flat, stationary axisymmetric black hole, we normalize $t^{\mu}$ such that $t^{\mu}t_{\mu}\rightarrow-1$ at infinity and $\varphi^{\mu}$ so that its closed orbits are $2\pi$-periodic.
[…]
From the rigidity theorem 2.6.3, the horizon normal Killing vector field $\chi^{\mu}$ is related to the stationarity $t^{\mu}$ and axisymmetry $\varphi^{\mu}$ as
$\chi^{\mu}=t^{\mu}+\Omega_{H}\varphi^{\mu}\,$, (37)
where the constant $\Omega_{H}$ is the angular velocity of the horizon with respect to the asymptotic stationary observers along the orbits of $t^{\mu}$. The parameter $v$, satisfying $\chi^{\mu}\nabla_{\mu}v=1$, is called the Killing parameter of $\chi^{\mu}$, which is in general different from the affine parameter $\lambda$ along the null geodesic generators of the Killing horizon.
For any Killing horizon with the associated Killing vector field $\chi^{\mu}$, we can define a surface gravity $\kappa$ as
$\kappa^{2}=-\dfrac{1}{2}(\nabla^{\mu}\chi^{\nu})\nabla_{\mu}\chi_{\nu}\,.$ (38)
Although the surface gravity $\kappa$, defined as above, is a local notion, the following result holds: {itembox}[l]Theorem 2.6.4 (The zero-th law) The surface gravity $\kappa$ of a stationary black hole must be a constant over its event horizon.
  • Bardeen-Carter-Hawking (1973) Bardeen:1973gs with DEC.
  • As will be discussed in Sec. 4, the surface gravity $\kappa$ of a stationary black hole can be interpreted as the temperature of the black hole.
[…]
$\delta M=\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}t^{\nu}\,,\quad\delta J=-\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}\varphi^{\nu}\,$, (39)
where $\sigma$ denotes a horizon cross-section with the area element ${\rm d}S$, and $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ the tangent to the horizon null geodesic generator with $\lambda$ being an affine parameter so that $k^{\mu}=\chi^{\mu}{\rm d}v/{\rm d}\lambda=\chi^{\mu}/(\kappa\lambda)=(t^{\mu}+\Omega_{H}\varphi^{\mu})/(\kappa\lambda)$. At this stage, we regard $T_{\mu\nu}$ as a small perturbation, which generates non-vanishing expansion and shear via the Raychaudhur equation. In the linear order of perturbation, neglecting the quadratic terms of the expansion and shear, we can write the Raychaudhuri equation for the horizon null geodesic generators as
$\dfrac{{\rm d}\theta}{{\rm d}\lambda}\simeq-8\pi GT_{\mu\nu}k^{\mu}k^{\nu}=-\dfrac{8\pi G}{\kappa\lambda}T_{\mu\nu}(t^{\mu}+\Omega_{H}\varphi^{\mu})k^{\nu}\,.$ (40)
Then, by integrating the Raychaudhuri equation along the horizon generators [^9] By multiplying $\kappa\lambda$ on both side of (40), and taking integration over $\sigma$ and along the null generator, the right-hand side immediately yields $8\pi G(\delta M-\Omega_{H}\delta J)$, while the left-hand side gives $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,(\kappa\lambda)\,\dfrac{{\rm d}\theta}{{\rm d}\lambda}=\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\left[\dfrac{{\rm d}(\lambda\theta)}{{\rm d}\lambda}-\theta\right]=-\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,\theta=-\kappa\delta A\,$, (41) where we have used the fact that before and sufficiently after one has dropped matter fields into the horizon, $\theta=0$ as a stationary state. , as well as horizon cross-section $\sigma$, we have {itembox}[l]Theorem 2.6.5 (The first law)
$\dfrac{\kappa}{8\pi G}\delta A=\delta M-\Omega_{H}\delta J$ (42)
  • Bardeen-Carter-Hawking (1973) Bardeen:1973gs .
  • In general relativity $M$ is equivalent to energy. As mentioned before, $\kappa$ and $A$ are proportional to the black hole’s temperature and entropy, respectively. The formula can then be interpreted as expressing the first law of black hole thermodynamics, with $\Omega_{H}\delta J$ viewed as the work term.
If the DEC holds, the integrand of (39) satisfies $T_{\mu\nu}k^{\mu}t^{\nu}\geqslant 0$, and the change of the black hole mass $\delta M$ is non-negative ($\delta M\geqslant 0$). This makes sense since the DEC implies the positivity of energy density and at the same time, it prohibits superluminal energy flow. This implies that matter with positive energy density can be swallowed by the black hole but never be able to escape from the event horizon. On the other hand, we have seen in the area theorem 2.6.2 that under the NEC, if we throw a positive energy density into the horizon, $\delta A>0$. Then, from the first law (42), we have,
$\delta M>\Omega_{H}\delta J\,.$ (43)
As a concrete example, let us consider the Kerr metric, which describes a stationary, rotating vacuum black hole with the two parameters $(M,J)$, whose surface gravity is given by,
$\kappa=\dfrac{1}{2GM}\dfrac{\sqrt{1-(J/GM^{2})^{2}}}{\left(1+\sqrt{1-(J/GM^{2})^{2}}\right)}\,.$ (44)
According to Theorem 2.6.4, $\kappa$ corresponds to the temperature, and the absolute zero-temerature (extreme black hole) state could be achieved in the limit
$J\rightarrow GM^{2}\,.$ (45)
Now let us attempt to attain this zero-temperature state by throwing a particle into a Kerr black hole of nearly zero temperature. Since near $\kappa=0$, the horizon angular velocity is $\Omega_{H}\simeq 1/2GM$, we have from (43),
$\delta(GM^{2})>\delta J\,.$ (46)
Thus, the change of the mass is always larger than that of the angular momentum, indicating that the extremal limit (45) can never be achieved by this process. This implies the third law of black hole thermodyamics to hold under the NEC and DEC. For more elaborate analysis and discussion, see, e.g., Israel:1986gqz ; Wald:1974hkz .
2.6.2 Topology and energy conditions
Let us discuss the constraints that the energy conditions can place on the possible topology of spacetime. We first discuss the topology of black hole horizon and then the topology of the exterior region of an asymptotically flat black hole spacetime.
To discuss horizon topology, we first recall the notion of stably marginally outer trapped surface (MOTS), which we have introduced in the previous section, as any $2$-sphere on the future event horizon of the Schwarzschild black hole. To define the notion of MOTS in more general setting, consider a closed, connected $2$-surface ${\cal T}_{0}$. Let $k^{\mu}$ be the tangent vector field for an out-going101010 In general, the notion of “out-going” or “outward” direction is not uniquely given. For an asymptotically flat spacetime with a single asymptotic region, if ${\cal T}_{0}$ separates a Cauchy surface $\Sigma$ into two disconnected parts $\Sigma_{\rm in}$ and $\Sigma_{\rm out}$, then “out-going” direction is defined as the one, say $\Sigma_{\rm out}$, that contains the asymptotic region. null geodesic congruence orthogonally emanating from ${\cal T}_{0}$, and $\theta_{+}$ be the associated expansion. Let $\ell^{\mu}$ be a past-directed, outward null vector field on ${\cal T}_{0}$. If $\theta_{+}=0$ and $\ell^{\nu}\nabla_{\nu}\theta_{+}\geqslant 0$ on ${\cal T}_{0}$, then ${\cal T}_{0}$ is said to be a MOTS. The following theorem determines the topology of MOTS.
{itembox}
[l]Theorem 2.6.6 (Topology of horizon cross-section) Under the DEC, the horizon cross-section ${\cal T}_{0}$ of a stationary black hole must be topologically $2$-sphere.
  • Propositoin 9.3.2 HE73 .
The proof consists of two parts. The first part is to show that any horizon cross-section of a stationary black hole is a MOTS, which is essentially the same as the first step of the proof of the rigidity theorem 2.6.3 (Proposition 9.3.1 of HE73 ). The second part is to show that MOTS must be a topologically $2$-sphere under DEC. Let us consider a horizon cross-section ${\cal T}_{0}$, which is assumed to be compact $2$-dimensional surface. By using the Einstein equations and the properties of the MOTS, $\theta_{+}=0$ and $\ell^{\nu}\nabla_{\nu}\theta_{+}\geqslant 0$, one can find the following inequality:
$\oint_{{\cal T}_{0}}\!\!{\rm d}S\,{\cal R}\geqslant 8\pi T_{\mu\nu}k^{\mu}(-\ell^{\nu})\geqslant 0\,$, (47)
[…]
Next, let us consider the topology of the black hole exterior region. We first describe two theorems concerning topological constraints on globally hyperbolic, asymptotically flat spacetimes.
{itembox}
[l]Theorem 2.6.7 (Topology and singularity) Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime, on which the NEC holds. Suppose that $M$ contains a non-simply connected Cauchy surface $\Sigma$. Then, $M$ must be geodesically incomplete.
  • Gannon (1975) Gannon:1975 .
  • The asymptotic flatness implies that the topologically non-trivial structure is isolated.
If a non-trivial topological structure is isolated, a natural question is whether it can be detected. To make this more concrete, let us imagine a situation where a bounded region enclosed by a sphere $\sigma$ in an asymptotically flat $3$-dimensional Cauchy surface $\Sigma$ contains a “handle” or a “wormhole” structure (see Figure 7).
Figure 7: (a) A “wormhole (handle-structure)” exists in the interior of a sphere $\sigma$ within a $3$-dimensional space $\Sigma$. (b) This situation can be constructed by excising two 33-dimensional balls from the interior of $\sigma$ and identifying their 22-sphere boundaries.
Can a distant observer actually detect this wormhole (handle) structure within that bounded region inside $\sigma$? For example, we would attempt to send a probe rocket into this bounded region so that the rocket would orbit the handle, become snagged by its non-trivial topological structure, and then return to a distant location to report its findings. Again under the NEC, we have the following theorem. {itembox}[l]Theorem 2.6.8 (Topological censorship) Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime. If the NEC holds in $M$, then any causal curve $\gamma$ from a point on $\mathscr{I}^{-}$ to a point on $\mathscr{I}^{+}$ can be continuously deformed to a timelike curve $\gamma_{0}$ which is contained in a simply connected neighborhood of $\mathscr{I}^{+}\cup\mathscr{I}^{-}$.
  • Friedman-Schleich-Witt (1993) Friedman:1993ty .
  • It is equivalent to say that $M$ must have as simple a topological structure as the neighborhood of $\mathscr{I}$. Since $M$ is globally hyperbolic, there exists a Cauchy surface $\Sigma$ so that $M\simeq{\mathbb{R}}\times\Sigma$. Then, it implies $\Sigma$ must be simply connected.
  • In fact, this theorem was shown in Friedman:1993ty under a weaker condition, the averaged null energy condition (ANEC), which we will discuss in the next subsection.
  • Here, we can consider $\gamma_{0}$ as the world-line of an observer who stays in the asymptotic region near infinity, and $\gamma$ as the world-line of a probe rocket that departs from the asymptotic region, explores the topology of spacetime $M$, and returns to the asymptotic region again. Two curves, $\gamma_{0}$ and $\gamma$, are said to be homotopic if one can be continuously deformed into the other. This means that the probe $\gamma$ does not get caught on a non-trivial topological structure like a “handle.”
  • “Censorship” is used metaphorically here to describe how physical laws prohibit the observation of topology.
Figure 8: (a) Consider a bundle of light rays (i.e., null geodesic congruence) traveling in a flat spacetime, which serves as a topologically trivial space. An ingoing light ray bundle perpendicular to 22-sphere $\sigma$ has a negative expansion $\theta<0$, and converges at the origin (which corresponds to the conjugate point to $\sigma$ and the expansion $\theta$ of the ingoing light rays changes from -$\infty$ to $\infty$ there) and then emerges from $\sigma$ as an outgoing bundle with a positive expansion $\theta>0$. (b) Now, consider a bundle of light rays entering the interior of $\sigma$ containing a wormhole (handle-structure), depicted in Figure 7, and subsequently exiting from $\sigma$ as an outgoing light ray. The cross-sectional area of the light rays entering one mouth of the wormhole decreases, having a negative expansion. However, after passing through the wormhole and exiting from the other mouth, the light ray bundle travels outward from $\sigma$ with its cross-sectional area increasing as it propagates towards infinity. This behavior of the light ray bundle leads to a contradiction.
The idea of the proof of Theorem 2.6.8 is as follows. Suppose an ingoing null geodesic congruence is emitted orthogonally from a $2$-sphere in the vicinity of past null infinity $\mathscr{I}^{-}$ toward inner region enclosed by $\sigma$ in which wormhole (handle) structure is confined. As depicted in Figure 8 (b), an ingoing null geodesic congruence perpendicular to $\sigma$ enters one mouth of the handle structure, and then exits from the other mouth as an outgoing bundle, propagating outward from $\sigma$, toward future null infinity $\mathscr{I}^{+}$. As the null congruence travels toward the entrance of the handle structure, the cross-section area $A$ of the congruence decreases (i.e., its expansion is negative $\theta<0$), but as it exits from the other mouth, $A$ increases (i.e., $\theta>0$). However, the congruence does not pass through a zero point, $A=0$ (i.e., the conjugate point where $\theta=\pm\infty$), while traversing the handle structure. Such a behavior of the expansion $\theta$ (i.e, changing its sign from negative to positive, without passing a conjugate point) contradicts the null focusing theorem 2.4.2, under the NEC. $\Box$
In conclusion, under the NEC, a distant observer cannot see an isolated non-trivial topology. This does not mean that asymptotically flat spacetime as a whole can only have a simple topology. Instead, it implies that if a non-trivial topological structure like a handle exists, it must be hidden within a region that is not visible from null infinity—namely, inside a black hole, just like the case of a closed trapped surface $\cal T$}, which also cannot intersect $J^{-}(\mathscr{I}^{-})$ (Prop. 2.6.1). In this respect, it is intriguing to note that Theorem 2.6.7 is analogous to Penrose’s singularity theorem (Theorem 2.5.6) with condition (ii) replaced with the existence of an isolated non-trivial topology in an asymptotically flat Cauchy surface.
As the exterior region of an asymptotically flat black hole, we define the domain of outer communications by
${\rm DOC}:=I^{+}(\mathscr{I}^{-})\cap I^{-}(\mathscr{I}^{+})\,$, (48)
as illusrated in Figure 9.
Figure 9: The domain of outer communications of an asymptotically flat black hole spacetime.
Applying Theorem 2.6.8 to $\rm DOC$}, one can find the following constraints on the horizon topology:
{itembox}
[l]Theorem 2.6.9 Suppose $(M,g_{\mu\nu})$ is a stationary, asymptotically flat spacetime containing a single asymptotically flat region, whose $\rm DOC$} is globally hyperbolic. Suppose the NEC holds. Then, (i) $\rm DOC$} is simply connected, and (ii) each connected component of the horizon cross-section is homeomorphic to a $2$-sphere.
  • Chrusciel-Wald (1994) Chrusciel:1994tr .
  • The first statement (i) results from the topological censorship (Theorem 2.6.8) and the second statement (ii) is proven by cobordism.
  • Compare with Theorem 2.6.6, in which the DEC is used, while here NEC.
[…]
Note that our discussion so far focuses on $4$-dimensional spacetimes. Similar techniques using NEC and cobordism in Theorem 2.6.9 can be applied to higher dimensional black holes, but they yield only weaker constraints Helfgott:2005jn . The generalization of the argument in Theorem 2.6.6 by using the DEC to higher dimensions yields that MOTS must admit a Riemannian metric of positive scalar curvature, accordingly the topology of MOTS must be of positive Yamabe type Galloway:2005mf ; Galloway:2006ws , which allows for various types of topologies other than the spherical one (see Emparan:2001wn for a seminal example of the “black ring” and Hollands:2010qy for thorough analysis, and see also, e.g., Hollands:2012xy ; Ida:2011jm for more general discussions of black holes and their topology in $4$ and higher dimensions).
[…]
Figure 10: Penrose diagram of the maximally extended Schwarzschild spacetime. Each point describes a $2$-sphere. There are two asymptotic regions, L and R, which are causally disconnected divided by the event horizon, and $\sigma_{\rm B}$ is the bifurcate surface or Einstein-Rosen bridge (ER). $S$ is a $2$-sphere in ${\rm DOC}_{\rm L}$, and $\theta_{\rm L}$ and $\theta_{\rm R}$, are the expansions of, respectively, the out-going (toward $\mathscr{I}^{+}_{\rm L}$) and in-going null geodesic congruences perpendicular to $S$. The in-going null geodesics with negative expansion $\theta_{\rm R}<0$ goes toward the central singularity but never enters ${\rm DOC}_{\rm R}$.
However, since the two DOCs are spacelike separated with ER located at $\partial{\rm DOC}_{\rm R}\cap\partial{\rm DOC}_{\rm L}$, any causal curve cannot travel from inside one DOC to the other, passing through ER. From this observation, to make our discussion more precise we define a traversable wormhole structure (TW) as follows. Let $(M,g_{\mu\nu})$ be a connected, globally hyperbolic asymptotically flat spacetime with two (mutually disconnected) null infinities, $\mathscr{I}_{\rm R}$ and $\mathscr{I}_{\rm L}$. (The global hyperbolicity is not strictly necessary, but is assumed to simplify the argument.) Consider ${\rm DOC}_{\rm R}$ and ${\rm DOC}_{\rm L}$, associated with $\mathscr{I}_{\rm R}$ and $\mathscr{I}_{\rm L}$, respectively, as defined above. If there exists a non-empty set
${\rm TW}:={\rm int}{\rm DOC}_{\rm R}\cap{\rm int}{\rm DOC}_{\rm L}\neq\emptyset\,$, (49)
then, $M$ is said to contain a traversable wormhole.
Then, applying a similar argument of the proof of Theorem 2.6.8, we immediately obtain the following:
{itembox}
[l]Proposition 2.6.10 (No traversable wormhole) The traversable wormhole structure (TW) must violate the NEC.
The idea of the proof is as follows. The essence of the proof of Theorem 2.6.8 is that any structure which allows the expansion $\theta$ of a null geodesic congruence to change its sign from negative to positive without passing a conjugate point contradicts the null focusing theorem 2.4.2, under the NEC. As illustrated in Figure 11. consider a $2$-sphere $S$ in ${\rm DOC}_{\rm L}\cap I^{-}({\rm TW})$ in a neighborhood sufficiently close to $\mathscr{I}^{-}_{\rm L}$ so that a future-directed ingoing (toward $\rm TW$}) null geodesic $\gamma$ perpendicularly emanating from $S$ has a negative expansion $\theta_{\rm R}<0$ on $S$. Passing through $\rm TW$}, the null geodesic $\gamma$ enters ${\rm DOC}_{\rm R}$ and eventually approaches $\mathscr{I}^{+}_{\rm R}$. In the neighborhood of $\mathscr{I}^{+}_{\rm R}$, the expansion of $\gamma$ must turn to be positive $\theta_{\rm R}>0$. This leads to the contradiction mentioned above. $\Box$
Figure 11: Penrose diagram of an asymptotically flat, traversable wormhole spacetime. There are two asymptotic regions, L and R. $S$ is a $2$-sphere in ${\rm DOC}_{\rm L}\cap I^{-}({\rm TW})$, but not in ${\rm DOC}_{\rm R}$. If $S$ is sufficiently close to $\mathscr{I}^{-}_{\rm L}$, the “in-going” null geodesics orthogonally emanating from $S$ have a negative expansion $\theta_{\rm R}<0$ in the region ${\rm DOC}_{\rm L}\setminus{\rm TW}$, then passing through the $\rm TW$}, eventually approach $\mathscr{I}^{+}_{\rm R}$ as “out-going” null geodesics with a positive expansion $\theta_{\rm R}>0$. From $\mathscr{I}^{+}_{\rm R}$, $S$ may look like a “closed trapped surface” which, however, should not be visible from $\mathscr{I}^{+}_{\rm R}$ under the NEC, according to Prop. 2.6.1.
2.7 Averaged energy conditions
So far we have reviewed the pointwise (locally defined) energy conditions (WEC, DEC, SEC, NEC) and their roles in classical general relativity. However, these conditions can be violated even within classical frameworks. For example, consider a scalar field $\phi$ with potential $V(\phi)$ with the energy-momentum tensor (7) in a spatially homogeneous and isotropic (FLRW) universe. Let $t^{\mu}=(\partial/\partial t)^{\mu}$ be a unit timelike vector orthogonal to any homogeneous and isotropic hypersurface. Then, due to the symmetry, $\phi$ can only be dependent on $t$, and we have $\left(T^{(\phi)}_{\mu\nu}-({1}/{2})T^{(\phi)}{}^{\lambda}{}_{\lambda}g_{\mu\nu}\right)t^{\mu}t^{\nu}=\dot{\phi}^{2}-V(\phi)$. Thus, if $\dot{\phi}^{2}<V$, which can occur for e.g., a massive free scalar field, or inflation/dark energy models, the SEC fails to hold.
In the semiclassical approach, quantum field effects can be incorporated in the expectation value of the renormalized energy-momentum tensor $\langle T_{\mu\nu}\rangle$, which replaces the classical counterpart in the right-hand side of the Einstein equations (1). There are a number of known examples in which the expectation value $\langle T_{\mu\nu}\rangle$ violates the locally defined energy conditions, even when their classical counterparts do not violate the locally defined energy conditions. In fact, Casimir effects violate all local energy conditions Roman:1986tp . For example, consider two parallel reflecting planes orthogonal to $z$-axis in Minkowski spacetime. Then the vacuum expectation value of the energy-momentum tensor for a scalar field between the two planes takes the form of $\langle 0|T_{\mu\nu}|0\rangle=-(C^{2}/L^{4})\mathrm{diag}(1,-1,-1,3)$ with $C^{2}$ being a positive constant and $L$ the separation distance between the two planes Birrell:1982ix , for which all the four locally defined energy conditions are violated.
While the pointwise energy conditions may fail in some restricted local regions, they may still hold in the rest of the universe. This observation suggests that non-local forms of energy conditions, which are defined by taking a suitable average over spacetime, may be more robust under many circumstances of interest. For example, in the case of Casimir effect above, although between the two planes, the NEC is violated along $z$-axis, since the two planes supply positive energy, if we integrate $\langle 0|T_{\mu\nu}k^{\mu}k^{\nu}|0\rangle$ along $z$-axis long enough passing through the two planes, we can have a positive energy value. In what follows, we consider generalizations of the energy conditions to the averaged forms. As we have seen in Sec. 2.4, 2.5, and 2.6, the energy conditions play a key role mainly through the focusing theorems, we will focus on the averaged SEC and NEC. As a weaker condition than SEC, the averaged strong-energy condition (ASEC) is simply formulated as follows.
{itembox}
[l]Definition 2.7.1 (ASEC) For every inextendible timelike geodesic curve with proper time $\tau$ and the tangent $\xi^{\mu}$, the integral along the timelike geodesic,
$\int{\rm d}\tau\left(T_{\mu\nu}-\dfrac{1}{2}Tg_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\,.$ (50)
  • Integrating the local timelike focusing ${\rm d}\theta/{\rm d}\tau\leqslant-R_{\mu\nu}\xi^{\mu}\xi^{\nu}$ in a finite interval $0,\tau]$, and using the Einstein equations, we have
    $\theta(\tau)\leqslant\theta(0)-8\pi G\int_{0}^{\tau}\!\!{\rm d}\tau\left(T_{\mu\nu}-\dfrac{1}{2}Tg_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\,.$ (51)
    Thus, the ASEC expresses the timelike geodesic focusing $\theta(\tau)\leqslant\theta(0)$ in averaged sense.
As for a weaker condition than NEC, we have the averaged null energy condition (ANEC)
{itembox}
[l]Definition 2.7.2 (ANEC) For every inextendible null geodesic curve with affine parameter $\lambda$ and the tangent $k^{\mu}$, the integral along the null geodesic,
$\int{\rm d}\lambda\,T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0\,.$ (52)
  • This is the condition used in the original statement of the topological censorship Friedman:1993ty (Theorem 2.6.8).
  • Hawking-Penrose singularity theorem (c.f., Theorem 2.5.7) is generalized by using ASEC and ANEC Tipler:1977zza ; Tipler:1978zz ; Fewster:2010gm . See also Borde:1987qr , for the focusing theorems under ASEC and ANEC, and further generalizations of the singularity theorems with averaged energy conditions.
ANEC was shown to hold in various situations (see, e.g., Yurtsever:1994wc ; Klinkhammer:1991ki ; Yurtsever:1990gx ; Wald:1991xn ; Ford:1995gb ; Kontou:2012ve ). In Minkowski background, ANEC was proven by a completely field theoretic argument Faulkner:2016mzt ; Hartman:2016lgu . The ANEC on curved spacetime which solves the semiclassical Einstein equations is called self-consistent ANEC, and has been shown to hold for pure and mixed states under certain conditions on the curvature scale Flanagan:1996gw .
There have also been discovered a number of cases in which the ANEC can be violated. For example, the ANEC for free scalar field can be violated if the background is locally flat but topologically non-trivial (e.g., torus compactification) Klinkhammer:1991ki . In curved spacetime, there are many known cases of the ANEC violation. For example, ANEC can be violated for a conformally coupled free scalar field in conformally flat spacetime Urban:2009yt (see also Visser:1994jb ; Ishibashi:2019nby ). If one considers a null geodesic of circular orbit in a Schwarzschild spacetime, (e.g. a null geodesic orbiting at the constant area radius $r=3G{\cal M}$), ANEC is not satisfied for the Hartle-Hawking state Visser:1996iw . Such a null geodesic of circular orbit is not achronal, while in many cases of interest in applications, e.g., the singularity theorems, the achronality of null geodesics plays a role. For this reason, there has been proposed the notion of achronal ANEC (aANEC), which asserts that the ANEC should hold for every complete achronal null geodesic but not necessarily on chronal null geodesics Wald:1991xn ; Graham:2007va (see also for the case of the aANEC violation Visser:1994jb ; Urban:2010vr ; Kontou:2015yha ).
The ANEC has recently attracted increasing attention, not only in general relativity, but also in quantum field theory and in the context of the holographic principle. We will discuss ANEC and its further generalizations in more detail in the subsequent sections 3 and 4.
3 Energy conditions and quantum field theory
In this section, we turn to energy conditions in quantum field theory in Minkowski spacetime. After giving a brief review of quantum information in a bipartite system, we consider the modular Hamiltonian $K_{A}$ associated with the subsystem $A$ in the vacuum. When the boundary of $A$ is planar, the Bisognano-Wichmann theorem Bisognano:1976za states that $K_{A}$ generates the Rindler boost, hence be written by smearing the stress-energy tensor on $A$. This theorem generalizes to the case when $A$ has the boundary on a null hypersurface Casini:2017roe , which allows us to derive the ANEC from the monotonicity of the vacuum modular Hamiltonian Faulkner:2016mzt . We also consider the quantum null energy condition (QNEC), which bounds the null-null component of the stress tensor by the second null derivative of entanglement entropy Bousso:2015mna ; Wall:2011kb . We sketch the derivation of QNEC following Kudler-Flam:2023hkl and leave the complete proof to Ceyhan:2018zfg ; Hollands:2025glm .
3.1 Quantum information measures
We will consider a bipartite system with the Hilbert space $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$. For a density matrix $\rho$ on $\mathcal{H}$, the reduced density matrix $\rho_{A}$ on the Hilbert space $\mathcal{H}_{A}$ of the subsystem $A$ is defined by
$\rho_{A}\equiv{\rm tr}_{\bar{A}}\left[\,\rho\,\right]\ ,$ (53)
[…]
$|].\displaystyle\rho_{A}={\rm tr}_{R}\left[\,\ket{\rho}\bra{\rho}\,\right$]\ . (54)
[…]
where we denote the union of the two subsystems $A$ and $R$ as $AR$.
The entanglement entropy $S(\rho_{A})$ of the subsystem $A$ is defined as the von Neumann entropy of the reduced density matrix $\rho_{A}$:
$S(\rho_{A})\equiv-{\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]\ .$ (56)
Entanglement entropy is known to satisfy the following inequality known as the strong subadditivity (SSA) Lieb1973cp :
$S(\rho_{AB})+S(\rho_{BC})\geq S(\rho_{A})+S(\rho_{C})\ .$ (57)
For a pair of two density matrices $\rho_{A}$ and $\sigma_{A}$ on $\mathcal{H}_{A}$, the (quantum) relative entropy is defined by
$S(\rho_{A}||\sigma_{A})\equiv{\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]-{\rm tr}_{A}\left[\rho_{A}\,\log\sigma_{A}\right]\ .$ (58)
Roughly speaking, the relative entropy measures the “distance” between two quantum states $\rho_{A}$ and $\sigma_{A}$. Indeed, it is non-negative
$S(\rho_{A}||\sigma_{A})\geq 0\ ,$ (59)
and vanishes only when the two states are the same:
$S(\rho_{A}||\sigma_{A})=0\qquad\Longleftrightarrow\qquad\rho_{A}=\sigma_{A}\ .$ (60)
Moreover, the relative entropy decreases monotonically as one shrinks the size of the subsystem: lindblad1975completely ; Araki1976zv ; uhlmann1977relative ; witten2018notes-8d6 : [^11] The monotonicity of the relative entropy is equivalent to the strong subadditivity of entanglement entropy (57).
$B\subset A\quad\rightarrow\quad S(\rho_{B}||\sigma_{B})\leq S(\rho_{A}||\sigma_{A})\ .$ (61)
For a density matrix $\sigma_{A}$, we can define the modular Hamiltonian as an analogue of energy:
$K^{(\sigma)}_{A}=-\log\sigma_{A}\ .$ (62)
Then, the relative entropy can be rewritten as
$S(\rho_{A}||\sigma_{A})=\Delta K^{(\sigma)}_{A}-\Delta S_{A}\ ,$ (63)
where
$\begin{aligned} \Delta K^{(\sigma)}_{A}&={\rm tr}_{A}\left[\rho_{A}\,K^{(\sigma)}_{A}\right]-{\rm tr}_{A}\left[\sigma_{A}\,K^{(\sigma)}_{A}\right]\ \ ,\\ \Delta S_{A}&=S(\rho_{A})-S(\sigma_{A})\ .\end{aligned}$ (64)
In fact, this expression coincides with the free energy $F=\Delta E-T\Delta S$, if $\sigma_{A}$ is the thermal density matrix.
When the system $\bar{A}$ is a purifying system of both $\rho_{A}$ and $\sigma_{A}$, we can simplify $\Delta K^{(\sigma)}_{A}$ to the following form:
[…]
3.2 Modular Hamiltonian in quantum field theory
The modular Hamiltonian defined by (62) does not have a local expression written by an integral of a local operator in general, unlike a Hamiltonian. However, when $\sigma$ is the vacuum density matrix $\Omega$, i.e.  $|\sigma=\Omega=\ket{\Omega}\bra{\Omega}$, the modular Hamiltonian (62) can be “geometric”, which is a local integral of the stress tensor, generating the time evolution. A well-known example is the Bisognano-Wichmann theorem Bisognano:1976za . Let region $A$ be the right half-space $A=\{t=0,x_{1}\geq 0\}$. We take the coordinate in Minkowski spacetime as
$t=r\,\sinh\tau\ ,\qquad x_{1}=r\,\cosh\tau\ .$ (66)
That is, the coordinate $(r,\tau)$ implies the Minkowski coordinate of the observer who is uniformly accelerating in the Minkowski spacetime, and covers only $|t|\leq x_{1}$, known as the (right) Rindler wedge (see the left panel of figure 12). The Rindler wedge is the causal domain (diamond) $D[A]$ of the region $A$. The causal domain $D[A]$ is the set of points $p$ such that all causal curves through $p$ intersect $A$, and the information at the point $p$ is completely determined by that of $A$. The Hamiltonian $K$ for the observer with respect to the Rindler time $\tau$ is given by
$K=\int_{A=\{t=0,\,x_{1}\geq 0\}}{\rm d}^{d-1}x\,x_{1}\,T_{tt}\ .$ (67)
The Bisognano-Wichmann theorem states that the vacuum modular Hamiltonian is the Hamiltonian for the uniformly accelerating observer, i.e. $K_{A}^{(\Omega)}=2\pi\,K$. The physical interpretation of this theorem is as follows. Starting from $x_{1}\geq 0$ at the initial time, the uniformly accelerating observer cannot obtain the information about $x_{1}\leq 0$ at the same time, since even the signal with the speed of light never reaches the observer (for this observer, $x_{1}<0$ is regarded as the inside of the black hole). Thus, the system with this observer is described by the vacuum reduced density matrix $\sigma_{A}$, even if the total system is in a pure state (vacuum). Furthermore, $\Omega_{A}=e^{-2\pi K}$ implies that the Rindler observer perceives a thermal bath at inverse temperature $\beta=2\pi$.
Figure 12: [Left] The right Rindler wedge is the causal domain $D[A]$ of the region $A$ as shown in gray color. The vacuum modular Hamiltonian coincides with the boost operator $K_{A}^{(\Omega)}=2\pi\,K$ due to the Bisognano-Wichmann theorem. [Right] The right Rindler wedge is mapped to the causal domain of the spherical region by the conformal transformation (68).
While the Bisognano-Wichmann theorem holds in any relativistic quantum field theory, there is another expression for the modular Hamiltonian in conformal field theory (CFT). Let region $A$ be the solid ball of radius $R$ centered at the origin: $A=\{X^{0}=0,\,0\leq r\leq R\}$. Here, $r\equiv\sqrt{\sum_{i=1}^{d-1}(X^{i})^{2}}$ is the radius of the spherical coordinates in the constant-time slice. The Rindler wedge is mapped to the causal domain $D[A]$ of the solid ball $A$, by conformal transformation, called the CHM map Casini:2011kv (figure 12):
$X^{\mu}=\frac{x^{\mu}+(x\cdot x)\,c^{\mu}}{1+2\,(x\cdot c)+(x\cdot x)\,(c\cdot c)}-2\,R^{2}\,c^{\mu}\ ,$ (68)
where $c^{\mu}=(0,1/(2R),\vec{0}\,)$. This means that, in CFT, the vacuum reduced density matrix of the right half-space is related to $\Omega_{A}$ by a unitary transformation. Then, the modular Hamiltonian of $\Omega_{A}$ is given by Casini:2011kv ; Hislop:1981uh
$K^{(\Omega)}_{A}=2\pi\int_{X^{0}=0,\,0\leq r\leq R}{\rm d}^{d-1}X\,\frac{R^{2}-r^{2}}{2R}\,T_{00}\ .$ (69)
We can consider the null deformed Rindler wedge as the causal domain $D[A]$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ as shown in figure 13. The modular Hamiltonian associated with such a region $A$ is also shown in Casini:2017roe to have a local expression similar to (67)
$K^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{\gamma_{A}(\bm{x}_{\perp})}^{\infty}{\rm d}x^{+}(x^{+}-\gamma_{A}(\bm{x}_{\perp}))\,T_{++}\left(x^{+},x^{-}=0,\bm{x}_{\perp}\right)\ .$ (70)
Figure 13: The null deformed Rindler wedge as the causal domain $D[A]$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ .
Having the applications to QFTs in mind, it is useful to extend the monotonicity of relative entropy (61) to the operator inequality Borchers:1995zg ; Blanco:2013lea :
$D[B]\subset D[A]\quad\rightarrow\quad\hat{K}^{(\rho)}_{B}\leq\hat{K}^{(\rho)}_{A}\$, (71)
where $D[A]$ is the causal domain of a spacelike hypersurface $A$, and $\hat{K}^{(\rho)}_{A}$ is the full modular Hamiltonian
$\hat{K}^{(\rho)}_{A}$ $\equiv K^{(\rho)}_{A}\otimes{\bf 1}_{\bar{A}}-{\bf 1}_{A}\otimes K^{(\rho)}_{\bar{A}}\$. (72)
3.3 Bekenstein bound
Bekenstein proposed the following conjecture (Bekenstein bound) Bekenstein:1980jp :
$S\lesssim R\,E\ ,$ (73)
where $S$ and $E$ are the entropy of the matter and energy in the region $A$ with typical size $R$. This bound arose through the following thought experiment. Consider an object with energy $E$ and entropy $S$, and drop into the black hole of radius $R$ in the three-dimensional space for simplicity (a similar argument holds in an arbitrary dimension). It is known that the black hole entropy is proportional to its area. As the radius of the black hole increases by $\Delta R\propto E$, the entropy also increases by $\Delta S_{\text{BH}}\propto R\Delta R$. The second law of thermodynamics implies that the entropy of the total system including an object and the black hole is bounded: $S\leq\Delta S_{\text{BH}}\propto R\,E$.
However, the Bekenstein bound becomes ambiguous in QFT, because the energy $E$ and the entropy $S$ have ultraviolet divergence. That is, the bound (73) is ill-defined.
To give a well-defined formulation of the Bekenstein bound, we apply the Bisognano-Wichmann theorem Casini:2008cr . Suppose that the region $A$ is the solid ball of the radius $R$ in CFT. Consider the reduced density matrices of an excited state and the vacuum state: $\rho_{A},\Omega_{A}$. The positivity of the relative entropy (59) and the relation to modular Hamiltonian (63) show
$\Delta S_{A}\leq\Delta K^{(\Omega)}_{A}\$, (74)
where the $K^{(\Omega)}_{A}$ in the right hand side is the vacuum modular Hamiltonian in CFT (69). Let $E$ be the energy of the excited state in the region $A$, then we can estimate $T_{tt}\sim E/R^{d-1}$, since $T_{tt}$ is the energy density. Integrating over the region $A$, we obtain $\Delta K_{A}^{(\Omega)}\sim R\,E$. Therefore, we can think of (74) as the quantum version of the Bekenstein bound (73). In addition, a different form of the Bekenstein bound was discovered using the monotonicity of the relative entropy in Blanco:2013lea .
In the next two sections, we show the applications of the monotonicity of the relative entropy (61) (or the modular Hamiltonian (71)) to deriving energy conditions in quantum field theory.
3.4 Averaged null energy condition
[…]
where we use the light-cone coordinates $x^{\mu}=(x^{+},x^{-},\bm{x}_{\perp})\ ,~x^{\pm}\equiv t\pm x_{1}$ in Minkowski spacetime (see also Definition 2.7.2). The ANEC holds for any QFT in Minkowski spacetime, which was proved in special cases (free theory, two dimensions, holography) Klinkhammer:1991ki ; Wald:1991xn ; Folacci:1992xg ; Verch:1999nt ; Kelly:2014mra , and in interacting theories Faulkner:2016mzt ; Hartman:2016lgu . Here, we give a proof using the monotonicity of the relative entropy along the line of Faulkner:2016mzt .
The modular Hamiltonian (70) for the null deformed Rindler wedge can also be written as
$K^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{0}^{\infty}{\rm d}x^{+}x^{+}\,T_{++}\left(x^{+}-\gamma_{A}(\bm{x}_{\perp}),\,x^{-}=0,\,\bm{x}_{\perp}\right)\ .$ (76)
We also consider another region $B$, whose boundary satisfies $\partial B=\{x^{-}=0,x^{+}=\gamma_{B}(\bm{x}_{\perp})\}$ and $\Delta\gamma(\bm{x}_{\perp})\equiv\gamma_{B}(\bm{x}_{\perp})-\gamma_{A}(\bm{x}_{\perp})>0$ (figure 14). Then, since the causal diamonds have a relation $D[B]\subset D[A]$, the monotonicity of the full modular Hamiltonian (71) yields the inequality
$\begin{aligned} 0&\leq\hat{K}^{(\Omega)}_{A}-\hat{K}^{(\Omega)}_{B}\\ &=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{-\infty}^{\infty}{\rm d}x^{+}\,\Delta\gamma(\bm{x}_{\perp})\,T_{++}(x^{+},x^{-}=0,\bm{x}_{\perp})\ .\end{aligned}$ (77)
[…]
Figure 14: [Left] The averaged null energy condition (ANEC). [Right] The regions $A$ and $B$ used to prove the ANEC.
ANEC has some applications. For example, $(3+1)$-dimensional CFTs have two central charges characterizing conformal anomaly (which is the violation of the conformal symmetry in a curved background), and the ratio of those is bounded from ANEC Hofman:2008ar ; Hofman:2016awc . In addition, the irreversibility of renormalization groups, called the $\mathcal{C}$-theorem, has been rederived in two- and four-dimensional QFTs by means of ANEC Hartman:2023qdn ; Hartman:2023ccw . In CFTs, there are some generalizations of ANEC, such as higher-spin ANEC Hartman:2016lgu ; Kravchuk:2018htv ; Meltzer:2018tnm and ANEC in de Sitter and anti-de Sitter backgrounds Rosso:2019txh .
3.5 Quantum null energy condition
While ANEC is a kind of global energy conditions, there is a local bound on the null energy density. This is called the quantum null energy condition (QNEC) Bousso:2015mna , whose statement is as follows: {itembox}[l]Definition 3.5.1 (QNEC)
[…]
Here $|]\rho_{A}={\rm tr}_{\bar{A}}\left[\,\ket{\rho}\bra{\rho}\,\right$] and $S(\rho_{A})$ is the entanglement entropy of a spacelike hypersurface $A$ in $x^{-}\leq 0$ whose boundary is anchored on $x^{-}=0,x^{+}=\gamma_{A}(\bm{x}_{\perp})$. Note that the second-order null shape derivative of the entanglement entropy in the right-hand side is not necessarily positive. ANEC follows from QNEC by integrating both sides of (78) along, assuming that the surface term $\delta S/\delta\gamma$ vanishes. QNEC arose in the context of quantum focusing conjecture (QFC) Bousso:2015mna , and was proven in free theory Bousso:2015wca ; Malik:2019dpg , in holographic theories Koeller:2015qmn , and in interacting theories Balakrishnan:2017bjg ; Ceyhan:2018zfg ; Hollands:2025glm . [^12] For interacting CFT, QNEC is conjectured to be saturated Leichenauer:2018obf ; Balakrishnan:2019gxl . QNEC also has applications to quenches Mezei:2019sla ; Kibe:2024icu ; Kibe:2025cqc , and renormalization group flow Casini:2023kyj .
Let us give a sketch of a proof of QNEC (our proof mainly follows that of Kudler-Flam:2023hkl ). We will use the expression of the relative entropy (63) and the vacuum modular Hamiltonian (70). By taking $\sigma$ in (63) to be the vacuum reduced density matrix, the second-order null shape derivative of the relative entropy becomes
$\frac{\delta^{2}}{\delta\gamma_{A}(\bm{x}_{\perp})^{2}}S(\rho_{A}||\Omega_{A})$ $=\frac{\delta^{2}}{\delta\gamma_{A}(\bm{x}_{\perp})^{2}}(\Delta K^{(\Omega)}_{A}-\Delta S_{A})$
[…]
$\frac{\delta^{2}S(\rho_{A}||\Omega_{A})}{\delta\gamma_{A}(\bm{x}_{\perp})^{2}}\geq 0\ .$ (80)
This positivity is guaranteed by the “ant-formula”:
[…]
By differentiating (72) with respect to $\gamma_{A}$ and taking the expectation value, the expectation value of the ANEC operator can be written as
[…]
On the other hand, the monotonicity of the relative entropy (61) implies that the relative entropy of the complementary region $\bar{A}$ becomes larger as we increase $\gamma_{A}$:
$\frac{\delta S(\tilde{\rho}_{\bar{A}}||\Omega_{\bar{A}})}{\delta\gamma_{A}(\bm{x}_{\perp})}\geq 0\ .$ (83)
Using (63) and repeating a similar argument as in (3.5), this inequality can be written as
$\frac{\delta S(\tilde{\rho}_{\bar{A}}||\Omega_{\bar{A}})}{\delta\gamma_{A}(\bm{x}_{\perp})}=\frac{\delta\langle K^{(\Omega)}_{\bar{A}}\rangle_{\tilde{\rho}_{\bar{A}}}}{\delta\gamma_{A}(\bm{x}_{\perp})}-\frac{\delta S(\tilde{\rho}_{\bar{A}})}{\delta\gamma_{A}(\bm{x}_{\perp})}\geq 0\ .$ (84)
Moreover, the strong subadditivity $S(\tilde{\rho}_{\bar{A}B})+S(\tilde{\rho}_{BC})\geq S(\tilde{\rho}_{\bar{A}})+S(\tilde{\rho}_{C})$ for the three regions
$\bar{A}$ $=\{x^{-}=0,-\infty\leq x^{+}\leq\gamma_{A}(\bm{x}_{\perp})\}\ ,$
$B$ $=\{x^{-}=0,\gamma_{A}(\bm{x}_{\perp})\leq x^{+}\leq\gamma_{A}(\bm{x}_{\perp})+\delta\}\ ,$ (85)
$C$ $=\{x^{-}=0,\gamma_{A}(\bm{x}_{\perp})+\delta\leq x^{+}\leq\infty\}\ ,$
yields the following inequality in the limit $\delta\rightarrow 0$:
$\frac{\delta S(\tilde{\rho}_{\bar{A}})}{\delta\gamma_{A}(\bm{x}_{\perp})}\geq\frac{\delta S(\rho_{A})}{\delta\gamma_{A}(\bm{x}_{\perp})}\ ,$ (86)
where we used the relation $\tilde{\rho}_{A}=\rho_{A}$ which follows from the definition of the purified density matrix $\tilde{\rho}$ for $\rho_{A}$. Combining (84) and (86), we obtain
$\frac{\delta\langle K^{(\Omega)}_{\bar{A}}\rangle_{\tilde{\rho}_{\bar{A}}}}{\delta\gamma_{A}(\bm{x}_{\perp})}\geq\frac{\delta S(\rho_{A})}{\delta\gamma_{A}(\bm{x}_{\perp})}\ .$ (87)
By adding ${-\delta\langle K^{(\Omega)}_{A}\rangle_{\rho_{A}}}/{\delta\gamma_{A}(\bm{x}_{\perp})}$ to both sides of (87), and using (63), we find
[…]
By combining this inequality with (82) and taking the infimum with respect to the purified state $\tilde{\rho}$, we finally obtain
[…]
The ant-formula follows if one can prove the inequality in the opposite direction. An interested reader may refer to the papers Ceyhan:2018zfg ; Hollands:2025glm for the detail of the proof.
4 Energy conditions and holographic principle
A recent development of the holographic principle, more specifically the AdS/CFT correspondence, serves as another powerful catalyst for integrating fundamental physics with quantum information theory. In this section, we will provide a brief overview of the basic aspects of the holographic principle, its relation to the black hole information paradox, and its applications to the ANEC.
4.1 Black hole entropy and holography
Although our main focus is on the energy conditions, a deeper understanding of their quantum extensions is incomplete without addressing holography.This subsection introduces the holographic principle, which—though seemingly peripheral—sheds crucial light on the emergence of energy, entropy, and spacetime itself in quantum gravity. In particular, black holes provide a striking arena where these ideas converge and force us to reconsider basic physical notions.
The path to holography begins with the thermal nature of black holes. Hawking’s discovery Hawking:1974rv ; Hawking:1975vcx that black holes emit blackbody radiation implies that they possess a well-defined temperature. From the laws of black hole mechanics, first formulated by Bekenstein Bekenstein:1973ur and Hawking, a consistent thermodynamic description emerged:
$T_{H}$ $=\dfrac{\hbar\kappa}{2\pi k_{\rm B}}\quad\mbox{(Hawking temperature)}$ (90)
$S_{\rm BH}$ $=\frac{A}{4G\hbar}\quad\mbox{(Bekenstein-Hawking entropy)}$ (91)
where $T_{\rm H}$ is called the Hawking temperature given by the surface gravity $\kappa$, and $S_{\rm BH}$ called the Bekenstein-Hawking entropy formula, proportional to the area $A$ of a cross-section of the event horizon as seen in Sec. 2.6.
This thermodynamic interpretation of black holes, though initially counterintuitive, finds natural footing in quantum field theory in curved spacetime. As clarified by the work of Bogoliubov and Unruh Unruh:1976db , particle content depends on the observer and the causal structure of spacetime. When a causal horizon is present, such as a black hole event horizon, it limits access to information and thus redefines the vacuum state and the notion of particles.
By dimensional analysis, since Schwarzschild black hole horizon radius scale as $r_{h}\sim GM$, the corresponding thermal energy satisfies
$k_{\rm B}T_{\rm H}\sim\frac{\hbar}{GM},$ (92)
and applying the first law of thermodynamics yields
$S\sim\frac{k_{\rm B}G}{\hbar}\int M{\rm d}M\sim\frac{k_{B}r_{h}^{2}}{\hbar G},$ (93)
thus reproducing the area law.
What is remarkable about this result is that the entropy scales with area, not volume, suggesting that the fundamental degrees of freedom responsible for this entropy reside not in the bulk, but on the boundary. This observation gave rise to the holographic principle, originally proposed by ’t Hooft tHooft:1993dmi and Susskind Susskind:1994vu , which posits that all the information contained in a volume of space can be represented by degrees of freedom on its boundary.
Since black holes are believed to be the densest and most entropic objects in nature, they offer a natural upper bound on the information content in a given region. This inspired the conjecture that the fundamental theory underlying quantum gravity is holographic: {itembox}[l]Conjecture 4.1.1 (Holographic principle) The complete description of a gravitational system in a bulk spacetime is equivalently formulated as a lower-dimensional, non-gravitational quantum field theory living on its boundary. A concrete and precise realization of this principle was proposed by Maldacena Maldacena:1997re in 1997, now famously known as the AdS/CFT correspondence.
Anti-de Sitter (AdS) spacetime $(M,g_{MN})$ is a maximally symmetric solution to the vacuum Einstein equations with a negative cosmological constant $\Lambda<0$, and its (conformal) boundary $\partial M$ is a spacetime of one dimension lower [^13] As a maximally symmetric spacetime, a $(d+1)$-dimensional AdS spacetime is represented as the hyperboloid $M\simeq S^{1}\times{\mathbb{R}}^{d}$ embedded in $(d+2)$-dimensional flat spacetime with the metric of signature (,,+,+,,+)(-,-$\cdots,+$). In the context of the AdS/CFT correspondence, by “AdS bulk” $(M,g_{MN})$ we mean its universal covering space, i.e., $M\simeq{\mathbb{R}}^{d+1}$, and therefore globally $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. . As discussed in Sec. 2.2 (see below Definition 2.2.5), since a cosmological constant $\Lambda$ can be viewed as a type of matter with the energy-momentum tensor $T^{(\Lambda)}_{MN}=-(\Lambda/8\pi G_{d+1})g_{MN}$, where $G_{d+1}$ denotes the bulk gravitational constant, AdS spacetime itself violates the WEC and the DEC, but satisfies the SEC and the NEC. Hence the timelike and null focusing theorems both hold in AdS spacetime.
Figure 15: [Left] The cylinder depicts a conformally compactified AdS spacetime. The cylinder’s interior corresponds to the AdS bulk $M$ and its surface represents the conformal boundary $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. The diamond-shaped subregion on $\partial M$ corresponds to $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{4},\eta_{\mu\nu})$. This subregion is enclosed by the red curves, which represent the future and past null infinity ($\mathscr{I}^{\pm}\subset\partial M$), and a single point $i^{0}\in\partial M$, which represents spatial infinity. [Right] 22-dimensional section of the conformal diagram of AdS spacetime. Two vertical double lines represent conformal infinity (i.e., the surface of the cylinder). The triangular subregion inside $M$ is covered by the horospherical chart (also known as the Poincare chart) given by (94), where $\infty>z>0$ inside $M$. Each vertical hyperboloid represents a constant-$z$ hypersurface, which is a $d$-dimensional Minkowski spacetime with the metric $(\ell/z)^{2}\eta_{\mu\nu}$. The boundary is located at $z=0$.
The AdS metric $g_{MN}$ in $(d+1)$-dimension can be expressed in the horospherical (Poincare) chart as
${\rm d}s^{2}_{(d+1)}=g_{MN}\,{\rm d}x^{M}{\rm d}x^{N}=\dfrac{\ell^{2}}{z^{2}}\left({\rm d}z^{2}+\eta_{\mu\nu}{\rm d}x^{\mu}{\rm d}x^{\nu}\right)\,$, (94)
where $\ell$ denotes the AdS curvature radius defined by $\ell^{2}=-d(d-1)/2\Lambda$ and where $z\rightarrow 0$ corresponds to infinity. Multiplying $g_{MN}$ by ${z^{2}}/\ell^{2}$ and taking the limit $z\rightarrow 0$, we obtain the $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{d},\eta_{\mu\nu})$ as part of the conformal boundary $\partial M$, which is considered as the arena for a quantum field theory (see Figure 15). As (part of) the $d$-dimensional boundary spacetime $(\partial M,g_{\mu\nu})$, one can also consider different geometries, such as $d$-dimensional de Sitter, AdS, or the static Einstein universe ${\mathbb{R}}\times S^{d-1}$ (see the metric (32) for $d=4$ case), other than the Minkowski spacetime. As for the $(d+1)$-dimensional bulk geometry $(M,g_{MN})$, one can, in principle, consider any asymptotically AdS spacetime, such as those including black holes inside but at large distances approaching the above AdS metric. In general, the AdS/CFT correspondence asserts the following: {itembox}[l]Conjecture 4.1.2 (AdS/CFT correspondence) A duality between a $(d+1)$-dimensional string theory defined in asymptotically Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) without gravity residing on its conformal boundary.
In this framework, string theory or quantum gravity phenomena in the bulk AdS spacetime can be fully encoded in the quantum dynamics of a conventional QFT on the boundary. In fact, the AdS/CFT correspondence is widely interpreted as providing a precise non-perturbative definition of quantum gravity in AdS via a lower-dimensional, non-gravitational field theory. Strictly speaking, the boundary theory must be a matrix-valued quantum field theory, typically a theory of $N\times N$ matrices such as $\mathcal{N}=4$ supersymmetric Yang–Mills theory, and the correspondence becomes exact in the large-$N$ limit. In this limit, the bulk Newton constant scales as $G\propto 1/N^{2}$, meaning that classical gravity emerges as the leading order approximation in a $1/N$ expansion. While these details are crucial for precision, we will not delve further into them here.
As an application to the energy conditions, the AdS/CFT correspondence is practically a powerful tool in computing the renormalized energy-momentum tensor for quantum fields. In general, any asymptotically AdS metric can be expressed near the conformal boundary—similar to (94)—in the following form,
$g_{MN}\,{\rm d}x^{M}{\rm d}x^{N}=\dfrac{\ell^{2}}{z^{2}}\left({\rm d}z^{2}+g_{\mu\nu}(z,x)\,{\rm d}x^{\mu}{\rm d}x^{\nu}\right)\,$, (95)
from whichi the $d$-dimensional metric on the conformal boundary at $z\rightarrow 0$ is read off as $g^{(0)}_{\mu\nu}=\lim_{z\rightarrow 0}g_{\mu\nu}(z,x)$. Then, near the conformal boundary, one can expand the $d$-dimensional metric $g_{\mu\nu}(z,x)$ in $z$-coordinate as
$g_{\mu\nu}(z,x)=\sum_{n=0}^{d}g^{(n)}_{\mu\nu}(x)z^{n}+h^{(d)}_{\mu\nu}(x)\log z^{2}+\cdots\,$, (96)
where the logarithmic term appears only when $d$ is even, and $g^{(2k+1)}_{\mu\nu}(x)=0$ for any integer $k$ satisfying $0\leqslant 2k+1<d$. This is called the Fefferman-Graham (FG) expansion. For given $g^{(0)}_{\mu\nu}(x)$, all the subleading coefficients $g^{(n)}_{\mu\nu}(x)$ with $n<d$ are determined solely by the boundary metric $g^{(0)}_{\mu\nu}(x)$ and the curvature tensors with respect to $g^{(0)}_{\mu\nu}(x)$, while the coefficient $g^{(d)}_{\mu\nu}(x)$ is related to the expectation value of the boundary field stress-energy tensor $\langle T_{\mu\nu}\rangledeHaro$:2000vlm : {itembox}[l]Formula 4.1.3 (Holographic stress-energy tensor) The renormalized stress-energy tensor on the boundary quantum field is given in terms of the FG expansion coefficient as,
$\langle T_{\mu\nu}\rangle=\dfrac{d\ell^{d-1}}{16\pi G_{d+1}}g^{(d)}_{\mu\nu}(x)+{X}_{\mu\nu}\,.$ (97)
Here $X_{\mu\nu}$ represents gravitational anomaly, which vanishes when the boundary dimension $d$ is odd. This formula stems from the fact that the AdS/CFT correspondence essentially equates the partition functions—and thus their effective actions—of the bulk and boundary theories. It is generally a formidable task to derive the left-hand side by using conventional quantum field theoretic methods. In the above formula, this task is considerably simplified by the use of the AdS/CFT correspondence, which replaces it with the FG expansion of the classical metric in the AdS bulk spacetime. In Sec. 4.4 and 4.5, we will discuss how the holographic stress-tensor derived from this formula can be used to examine the ANEC for strongly coupled quantum fields. However, before that, we will discuss more on recent progress in the holographic principle and understanding the black hole information paradox in the next two subsections.
4.2 AdS/CFT correspondence and the black hole information paradox
From the viewpoint of energy, the AdS/CFT correspondence implies that energetic processes in the bulk, such as black hole formation and evaporation, have a dual description in terms of unitary evolution within the boundary CFT. In this sense, the unitarity of the boundary theory guarantees the unitarity of the bulk gravitational dynamics. However, the situation is more subtle. If one takes the AdS/CFT correspondence at face value, then black hole evaporation must indeed proceed as a unitary process, in full accordance with quantum mechanical principles. Yet, it is crucial to emphasize that the AdS/CFT correspondence, despite the overwhelming body of evidence and many successful checks, remains a conjecture rather than a rigorously proven theorem. Moreover, powerful and long-standing arguments challenge the idea of unitary black hole evaporation. These stem from Hawking’s original semiclassical analysis, which predicts that Hawking radiation is purely thermal and uncorrelated with the matter that collapsed to form the black hole. This suggests an irreversible loss of information, in apparent violation of unitarity. This conflict constitutes the core of the black hole information paradox. If unitarity and semiclassical gravity cannot be simultaneously maintained, then one of our foundational assumptions about spacetime or quantum mechanics must give way. Far from a technical inconsistency, the paradox offers a crucial window into the fundamental nature of quantum gravity.
In the remainder of this subsection, we introduce a series of recent developments, particularly the concepts of quantum extremal surfaces and the island formula, that offer a striking resolution to the paradox within the holographic paradigm.
To appreciate the tension between semiclassical gravity and unitarity, let us consider Hawking’s original argument in a concrete setting. Suppose we have a large black hole, such as a supermassive one at the center of a galaxy, with a mass on the order of $10^{5}$ solar masses. The horizon radius in such a case is comparable to the size of the Earth. Consequently, the curvature near the event horizon is extremely small, and the semiclassical approximation, that is, treating spacetime classically while quantizing matter fields, should be valid to high precision.
In this approximation, the region near the horizon effectively resembles an expanding spacetime, in the sense that the proper distance between points just inside and just outside the horizon increases with time. This effective “expansion” is not due to a global cosmological expansion, but rather arises from the diverging world-lines of near-horizon observers: the local geometry induces a relative stretching across the horizon, which facilitates the production of entangled pairs (see Figure 16). Quantum fluctuations in this region give rise to entangled particle pairs, one falling into the black hole and the other escaping to null infinity $\mathscr{I}^{+}$ as Hawking radiation. As discussed in earlier sections, this process generates entanglement between the black hole interior and the outgoing radiation, causing the entanglement entropy of the radiation to increase steadily over time. If the process continues in this fashion throughout the evaporation, the entropy of the Hawking radiation would grow monotonically. Since the geometry near the horizon remains nearly classical during most of the black hole’s lifetime, especially for such large black holes, there appears to be no mechanism within the semiclassical framework that could halt or reverse this entropy growth.
[…]
However, as Page pointed out in Page:1993df ; Page:1993wv , if the evaporation process is ultimately unitary, the entanglement entropy should not increase indefinitely. Rather, it should follow a specific curve, now known as the Page curve: it grows initially but must eventually reach a maximum and then decrease to zero as the black hole disappears, restoring a pure final state. It is worth emphasizing here that this turning behavior—growth followed by decrease—is a natural consequence when one considers the mutual nature of entanglement entropy. The turning point is called the Page time, which roughly corresponds to the moment when the number of degrees of freedom in the emitted radiation equals that of the remaining black hole. This typically occurs when the black hole has lost about half of its initial mass.
This leads to a striking tension: at the Page time, the black hole is still macroscopically large, and the semiclassical approximation near the horizon should remain valid. Yet, the Page curve requires the entanglement entropy to start decreasing at this stage. This apparent contradiction—between the need for unitary evolution implied by the Page curve and the expected validity of semiclassical gravity—is at the core of the black hole information paradox.
A breakthrough in resolving this paradox came with the application of the Ryu-Takayanagi (RT) formula and its quantum-corrected version, which incorporates so called islands in the gravitational path integral. According to the island formula, the entanglement entropy of the Hawking radiation is given not just by the von Neumann entropy of the radiation in quantum field theory, but also includes contributions from a region inside the black hole, the island, bounded by an extremal surface. This surface plays the role of a generalized RT surface, and its inclusion leads to the Page curve behavior in entropy, thus restoring consistency with unitarity.
4.3 Entanglement entropy, the Ryu-Takayanagi formula, and the Island paradigm
A pivotal insight in the effort to resolve the black hole information paradox is that entropy, particularly entanglement entropy, plays a fundamental role in the structure of quantum gravity. In semiclassical gravity, the entropy of a black hole is famously captured by the Bekenstein-Hawking formula, {itembox}[l]Formula 4.3.1 (Bekenstein-Hawking entroy)
$S_{\text{BH}}=\frac{A}{4G\hbar}\,$ (98)
where $A$ denotes the area of the event horizon as mentioned at the beginning of this section. This area-law scaling stands in sharp contrast to the volume-law behavior characteristic of entropy in local quantum field theory, hinting at a radical departure from conventional locality. It suggests that the underlying degrees of freedom in quantum gravity must be organized in a nonlocal fashion.
This conceptual leap gained concrete form in the context of the AdS/CFT correspondence, especially with the introduction of the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef . For a spatial subregion $A$ of the boundary CFT, the RT formula computes its entanglement entropy as {itembox}[l]Formula 4.3.2 (Ryu-Takayanagi formula)
$S_{A}=\min_{\gamma}\left[\frac{\mathrm{Area}(\gamma_{A})}{4G}\right]\,$, (99)
where $\gamma_{A}$ is the codimension-2 minimal surface in the bulk AdS geometry, anchored to $\partial A$ and homologous to $A$. This prescription geometrizes entanglement: it relates a quantum information-theoretic quantity to a geometric property of spacetime, suggesting that entanglement is not merely a diagnostic tool but a foundational principle underlying gravitational physics.
To further appreciate the physical meaning of the RT formula, one can ask: given a bipartition of the boundary system into regions $A$ and $\bar{A}$, what is the corresponding object in the bulk that computes the entanglement entropy $S_{A}$ in the boundary theory? Since entanglement entropy is a quantity defined entirely in terms of the boundary degrees of freedom, its bulk dual must also be determined solely from boundary data. The RT formula proposes that this dual is the area of a minimal surface $\gamma_{A}$ in the bulk, anchored to $\partial A$ and homologous to $A$.
To gain intuition about how this minimal surface behaves, consider a thought experiment in which AdS spacetime contains a black hole in the bulk and on the boundary the size of region $A$ is gradually increased, starting from an infinitesimal region. Initially, $A$ is very small and $\bar{A}$ occupies nearly the entire boundary. In this regime, the corresponding minimal surface $\gamma_{A}$ lies close to the boundary and does not probe deep into the bulk. As the size of $A$ grows, and $\bar{A}$ correspondingly shrinks, the minimal surface $\gamma_{A}$ begins to dip further into the bulk and eventually wraps around the black hole horizon (see Figure 17).
Figure 17: A conformal diagram of a spatial section of AdS spacetime containing a black hole. The entire circle depicts the spatial section of the conformal boundary $\partial M$. (a) When AA (red curve) is small, the minimal surface $\gamma_{A}$ (orange dashed curve) in $M$ anchored to $\partial A$ lies close to the boundary $\partial M$. (b) When AA is large, occupying nearly the entire boundary $\partial M$, $\gamma_{A}$ nearly wraps around the black hole horizon.
In the limiting case where region $A$ covers the entire boundary (i.e., $\bar{A}$ is empty), the entanglement entropy $S_{A}$ corresponds to the von Neumann entropy of the global mixed state seen by the external observer. In a holographic theory with a black hole in the bulk, this entropy must match the Bekenstein-Hawking entropy of the black hole:
$S_{A}=S_{\text{BH}}=\frac{A_{\text{horizon}}}{4G}.$ (100)
This matching of entropies in the fully mixed limit strongly supports the identification of the black hole horizon itself as the minimal surface $\gamma_{A}$ in the RT formula (99).
Thus, the behavior of $\gamma_{A}$ across varying bipartitions encodes a consistent geometric realization of quantum entanglement in holographic theories. The RT prescription not only satisfies the area law for black hole entropy in the appropriate limit, but also ensures a smooth interpolation across partial regions, reflecting the structure of entanglement entropy in quantum field theory. This consistency under varying partitions reinforces the idea that $\gamma_{A}$ serves as a geometric surrogate for boundary entanglement, and is a key element in the broader program of interpreting spacetime geometry as emergent from quantum entanglement.
The RT formula thus supports the now widespread intuition that “entanglement builds geometry.” It suggests that spacetime geometry may emerge from the entanglement pattern in a dual quantum system, embodying the holographic principle. However, the original RT prescription applies only to static configurations and is valid only at leading order in $1/N$, where bulk quantum effects are neglected. To describe dynamical spacetimes, such as those involving black hole formation and evaporation, the RT formula must be generalized to the covariant Hubeny-Rangamani-Takayanagi (HRT) prescription Hubeny:2007xt , which replaces minimal surfaces with extremal ones. Furthermore, to incorporate quantum corrections—crucial in scenarios like black hole evaporation—one must go beyond the classical approximation and include subleading contributions in $1/N$, leading to the quantum extremal surface (QES) prescription.
These bring us to the notion of quantum extremal surfaces (QES), developed by Faulkner, Lewkowycz, and Maldacena Faulkner:2013ana and further formalized by Engelhardt and Wall Engelhardt:2014gca . The QES prescription replaces the classical minimal-area surface with a surface that extremizes the generalized entropy: {itembox}[l]Formula 4.3.3 (Quantum extremal surface)
$S_{\text{gen}}(\gamma)=\mbox{Ext}_{\gamma}\left[\frac{\mathrm{Area}(\gamma)}{4G}+S_{\text{bulk}}(\Sigma_{\gamma})\right]\,$, (101)
where $\Sigma_{\gamma}$ is the bulk region bounded by $\gamma$ and the boundary region $A$, and $S_{\text{bulk}}(\Sigma_{\gamma})$ is the von Neumann entropy of quantum fields in that region. The entanglement entropy of region $A$ is then computed by extremizing this generalized entropy over all candidate surfaces $\gamma$ anchored to $\partial A$. Note that the idea of including quantum corrections from matter fields in this way is conceptually rooted in Bekenstein’s notion of generalized entropy Bekenstein:1974ax . There, Bekenstein proposed that the total entropy should consist of both the area term and the matter entropy outside the horizon, anticipating the structure of $S_{\text{gen}}$ long before the advent of the AdS/CFT correspondence, even before the discovery of Hawking radiation.
This formalism lays the foundation for understanding the entanglement wedge, a concept central to the program of holographic quantum error correction Almheiri:2014lwa ; Dong:2016eik ; Harlow:2016vwg . According to the QES prescription, the region $\Sigma_{\gamma}$ bounded by $A$ and the QES $\gamma$ defines the entanglement wedge of $A$. Bulk operators in $\Sigma_{\gamma}$ can, in principle, be reconstructed from data in $A$.
A major conceptual advance came with the development of the island formula, especially in the context of evaporating black holes coupled to an external non-gravitating bath. This framework was pioneered by Almheiri, Engelhardt, Marolf, and Maxfield Almheiri:2019hni and Penington Penington:2019npb , and further developed by Almheiri:2019psf . In such a setup, the entanglement entropy $S_{R}$ of the Hawking radiation $R$ is given by {itembox}[l]Formula 4.3.4 (Island formula)
$S_{R}=\mbox{Ext}_{I}\left[\frac{\mathrm{Area}(\partial I)}{4G}+S_{\text{bulk}}(R\cup I)\right]\,.$ (102)
Here $I$ denotes a candidate island region inside the black hole, and $\partial I$ is its boundary (see Figure 18). At early times, the minimal configuration has no island, reproducing Hawking’s prediction of monotonically increasing radiation entropy. However, after the Page time, a new configuration appears where a nontrivial island emerges near the black hole horizon that becomes dominant, yielding a smaller generalized entropy. As a result of this island, the entropy of the radiation begins to decrease after the Page time, successfully reproducing the Page curve.
[…]
The reason for this transition from growth to decay in the radiation entropy lies in the emergence of the island near the horizon. Once the island is included in the entanglement wedge of the radiation, the interior degrees of freedom of the black hole—those within the island—are effectively treated as part of the same quantum system as the Hawking radiation itself. Consequently, the entanglement entropy between the radiation and the black hole begins to decrease, as the radiation now purifies the interior. Furthermore, the island formula contains a geometric term—the area of the island’s boundary—as its first contribution. As the black hole evaporates, this area term shrinks accordingly. Since both the area term and the bulk entropy $S_{\text{bulk}}(R\cup I)$ diminish over time, the generalized entropy of the radiation decreases and eventually vanishes. This reflects the restoration of purity in the final state, as required by unitarity.
This transition restores consistency with unitarity and provides a semiclassically controlled resolution to the black hole information paradox. It indicates that, despite the semiclassical appearance of the geometry, the interior of the black hole becomes encoded in the radiation through the quantum entanglement.
The emergence of islands thus implies a striking nonlocality: information that appears to reside deep inside the black hole is accessible from the Hawking radiation outside. In holographic gravity, spacetime is not merely a stage on which entanglement plays out; it is itself a consequence of entanglement. The RT formula, its quantum extension via QES, and the island paradigm together unify geometry, entropy, and quantum information.
Moreover, this flow of insights is not unidirectional. While boundary entanglement patterns determine the bulk geometry, the consistency and causal structure of the bulk spacetime also impose powerful constraints on the boundary theory. A notable example is the derivation of boundary energy conditions, such as the averaged null energy condition (ANEC), from bulk causality and modular flow. In the next section, we explore how this bidirectional logic further deepens our understanding of quantum gravity.
4.4 Causal constraints and the ANEC in holography
A landmark result demonstrating how bulk geometric principles constrain boundary dynamics is the holographic proof of the ANEC on the Minkowski background by Kelly and Wall Kelly:2014mra . Their derivation relies on the requirement of bulk causal consistency, and establishes that the ANEC holds in any boundary CFT admitting a semiclassical gravitational dual.
As we had reviewed before, the ANEC asserts that the null-null component of the stress-energy tensor, integrated along a complete null geodesic, is non-negative:
$\int_{-\infty}^{\infty}{\rm d}\lambda\,\langle T_{\mu\nu}k^{\mu}k^{\nu}(\lambda)\rangle\geqslant 0\,$, (103)
where $k^{\mu}$ is a future-directed null vector and the expectation value is given by the holographic Formula 4.1.1 deHaro:2000vlm . The key insight in the Kelly-Wall argument is that any violation of the ANEC would imply a breakdown of boundary causality via its holographic encoding in bulk geometry. More precisely, if bulk signals could propagate in a manner that violates the causal structure of the boundary theory, then the duality itself would become inconsistent. To formalize this, they consider null deformations of boundary subregions and analyze the corresponding changes in the bulk extremal surfaces computing their entanglement entropy. A violation of the ANEC would induce deformations that allow extremal surfaces to access regions outside the causal domain of the boundary, thereby violating entanglement wedge nesting and the monotonicity of relative entropy.
This perspective is closely related to a foundational result by Gao and Wald Gao:2000ga , who showed that in asymptotically AdS spacetimes satisfying the NEC, bulk causal propagation respects boundary causal structure. More precisely, consider an asymptotically AdS spacetime $(M,g_{MN})$ and its conformal boundary $\partial M$. Suppose a pair of boundary two points $p,q\in\partial M$ are connected by a null geodesic $\gamma$ which is lying entirely in $\partial M$ and is achronal with respect to the boundary metric $g^{(0)}_{\mu\nu}$. If there is a timelike curve through the bulk $M$ which also connects the boundary two points $p$ and $q$, then the entire spacetime $M\cup\partial M$ is said to admit a bulk-shortcut. {itembox}[l]Theorem 4.4.1 (Gao-Wald theorem) Consider an asymptotically AdS spacetime $M$ which has no causal pathology, such as a naked singularity or causality violation, in the bulk. Then, under the (A)NEC and the null generic condition in the bulk, there exists no bulk-shortcut and the fastest possible causal curve connecting two boundary points $p,q$ is an achronal null geodesic $\gamma$ lying entirely on the boundary $\partial M$.
  • If the null generic condition (Definition 2.4.1) is dropped, the exact AdS spacetime is included. For the exact AdS spacetime case, besides the boundary null geodesic $\gamma$, there exists an achronal null geodesic in the bulk which also connects $p,q\in\partial M$.
  • The lack of this no shortcut property results in the violation of the weak cosmic censorship Ishibashi:2019nby .
Gao-Wald (Theorem 2 of Gao:2000ga ) implies that the propagation of light through an asymptotically AdS spacetime will always be delayed relative to propagation through the exact AdS spacetime (see Figure 19). In other words, causal curves in the bulk cannot allow information to travel outside the light-cone defined by the boundary metric. This bulk-to-boundary causality theorem implies that any violation of the ANEC in the boundary theory would necessarily reflect a breakdown of this property in the bulk.
This leads to the remarkable conclusion that in any holographic theory governed by semiclassical Einstein gravity with matter obeying the NEC, boundary causality enforces the ANEC (103). That is, the causal structure of the bulk spacetime, via its encoding in extremal surfaces and entanglement wedges, imposes physical energy constraints on the dual quantum field theory.
Figure 19: A conformal diagram of an asymptotically AdS spacetime $M$ and its conformal boundary $\partial M$. Consider a boundary achronal null geodesic $\gamma$ connecting two points $p,q$ in $\partial M$ entirely lying on $\partial M$, and also a bulk null geodesic $\gamma^{\prime}$ starting from $p$ to a point $r\in J^{+}(q,\partial M)$. Gao-Wald theorem implies that, in general, $r\neq q$ but instead lies in the chronological future of $q$, i.e., $r\in I^{+}(q,\partial M)$. If the null generic condition is dropped, it is possible for $r$ to coincide with $q$ (i.e., $r=q$). This happens only when $\gamma^{\prime}$ is an achronal null geodesic in the bulk, which is the case of the exact AdS bulk. This implies that the propagation of light through an asymptotically AdS spacetime will always be delayed $\varDelta T\geqslant 0$ relative to propagation through the exact AdS spacetime.
4.5 Conformally invariant ANEC
As briefly discussed in Sec. 2.7, it is important to note that the ANEC can be violated in conformally coupled free scalar fields on conformally flat spacetimes Urban:2009yt (see also Visser:1994jb for more general curved backgrounds), as well as in strongly coupled field theories with gravity duals in the holographic setting Ishibashi:2019nby . A key feature of these examples of the ANEC violation is that a local violation of the NEC can be enhanced via conformal transformation, primarily because the ANEC is not conformally invariant. As a result, the ANEC may be violated in the conformally transformed spacetime. In this subsection, as a generalization of the ANEC in the context of the AdS/CFT correspondence, we review a weighted ANEC or conformally invariant ANEC Iizuka:2019ezn ; Iizuka:2020wuj , which applies to spatially compact spacetimes.
A canonical example of spatially compact spacetimes is the static Einstein universe. As explained in Sec. 4.1, the conformal boundary $\partial M$ of $(d+1)$-dimensional AdS spacetime is globally the static Einstein universe in $d$-dimension, $\partial M\simeq{\mathbb{R}}\times S^{d-1}$. Let us consider a pair of two boundary points $p,q\in\partial M$ such that $p$ and $q\in J^{+}(\partial M)$ can be connected by an achronal null geodesic $\gamma$ with the tangent vector $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ lying entirely in $\partial M$, and $q$ corresponds to the antipodal point of $p$ in spatial section $S^{d-1}$, as depicted in Figure 19. According to Gao-Wald theorem, under the bulk NEC, there is no bulk-shortcut connecting $p,q$ and $\gamma$ must be the fastest causal curve connecting the two point. Let $\lambda_{-}$ and $\lambda_{+}$ be the affine parameter values of $\gamma$ at the point $p$ and $q$, respectively, and $\eta$ be the magnitude of the boundary Jacobi field along $\gamma$, which vanishes at $p,q$, i.e., $\eta(\lambda_{\pm})=0$ but is non-vanishing at any other points between $p$ and $q$, due to the achronality of $\gamma$ (Proposition 2.5.1). On this curved boundary, we can evaluate the expectation value of the energy-momentum tensor $\langle T_{\mu\nu}\rangle$ by using the holographic formula (97). Note that when the boundary dimension $d$ is odd, $X_{\mu\nu}$ in (97) identically vanishes, while when $d$ is even, $X_{\mu\nu}$ must be taken into account. Then, applying Gao-Wald theorem, we find the following inequality: {itembox}[l]Definition 4.5.1 (Holographic weighted ANEC)
$\int_{\lambda_{-}}^{\lambda_{+}}\!\!\ {\rm d}\lambda\>\eta^{d}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\geqslant C\,.$ (104)
  • When the boundary dimension $d$ is odd, the lower bound given by constant $C$ in the right-hand side vanishes, and this inequality is manifestly conformally invariantIizuka:2019ezn , and is called the conformally invariant ANEC (CANEC). In the framework of the conformal boundary field theory, all physical quantities should be described in a conformally invariant way. In this sense, the CANEC is expected to be more useful than the ANEC to put restrictions on the boundary stress-energy tensors (see also Rosso:2019txh ; Rosso:2020cub for similar analysis of the weighted ANEC).
  • When the boundary dimension $d$ is even, the lower bound constant $C$ is expressed in terms of the boundary integration along $\gamma$ of a certain combination of boundary geometric quantities such as the boundary curvature tensor and the expansions of the boundary null congruence of $\gamma$ (see Iizuka:2020wuj for the detailed expression).
[…]
${\cal M}\geqslant 0\,.$ (105)
This suggests that the Gao-Wald theorem is related to the positive energy theorem (see Theorem 2.2.4) in asymptotically AdS bulk spacetimes Gibbons:1982jg . This connection arises because the Schwarzschild-AdS spacetime with negative mass ${\cal M}<0$ contains a naked singularity in the bulk, but under the conditions of the Gao-Wald theorem, the occurrence of a bulk naked singularity is forbidden Ishibashi:2019nby . In fact, the non-negative time delay $\varDelta T\geqslant 0$, which is a consequence of Gao-Wald theorem, can be used to derive not only the boundary weighted ANEC (104) but also the positivity of the bulk weighted time-averaged energy Page:2002xn . In this way, the causality of the bulk and boundary spacetimes relates the energy positivity of the bulk and boundary.
5 Summary and discussions
In this article, we reviewed various applications of the energy conditions in general relativity, quantum field theory, and the holographic principle, and discussed their profound connection to quantum information. In Section 2, we first described the locally defined energy conditions: WEC, DEC, SEC, and NEC. The former two primarily concern the positivity aspects of matter field energy-momentum, while the latter two the geometric aspects via the geodesic focusing. We explained how these energy conditions are used—mainly through the focusing theorems—in the proofs of the singularity theorems, the black hole mechanics (or the laws of black hole thermodynamics), and theorems on spacetime topology. We also briefly discussed the violation of these local energy conditions and described their generalizations to the non-local, averaged energy conditions: ASEC and ANEC.
In Section 3, we described the ANEC and QNEC in quantum field theory. The key to their derivations is the positivity and monotonicity of relative entropy. A Rényi generalization of QNEC is proposed and discussed in Lashkari:2018nsl ; Moosa:2020jwt ; Roy:2022yzm , and it would be interesting to see if a similar derivation to QNEC would work for the Rényi generalization. ANEC is also assumed to derive non-trivial constraints in defect CFTs, i.e., CFTs with non-local operators (such as lines, surfaces and boundaries) in e.g., Jensen:2018rxu ; Meineri:2019ycm ; Herzog:2020bqw ; Chalabi:2021jud . It remains open whether ANEC holds in the presence of defect as the proof in subsection 3.4 assumes that the state is in the vacuum and a generalization of ANEC to defect CFTs is left for future investigations.
In Section 4, we first briefly described the holographic principle and the AdS/CFT correspondence. We then reviewed recent progress in understanding the black hole information paradox within the framework of holography, introducing key concepts such as the Ryu-Takayanagi formula, QES, Island formula. We then discussed the relationship between ANEC and holography. The AdS/CFT correspondence equates bulk and boundary field theories, which naturally leads to the question of how their respective energy conditions are related. We showed, via the Gao-Wald theorem Gao:2000ga , that these two energy conditions are related by bulk and boundary causalities. As a key application of this theorem, we described the holographic proof of ANEC in Minkowski background Kelly:2014mra , and also, as its generalization to curved backgrounds, the holographic weighted and conformally invariant ANEC Iizuka:2019ezn ; Iizuka:2020wuj . For the even-boundary dimension case, the holographic weighted ANEC (104) has the non-vanishing lower bound $C$, and thus takes a similar form to the quantum energy inequality (QEI) (see, e.g., Fewster:2012yh ; Kontou:2020bta ). The clarification of the relationship between the holographic weighted ANEC and QEI is still ongoing. In the standard formulation of the AdS/CFT correspondence, the boundary metric is fixed, not considered as a dynamical variable, and so far, all holographic proofs of ANEC are performed on a fixed background. It remains an open question whether the self-consistent ANEC can be shown by using the holographic semiclassical Einstein equations Compere:2008us ; Ishibashi:2023luz .
As we have seen throughout this paper, locally defined energy conditions are often violated by quantum effects, such as vacuum fluctuations or the Casimir effect. It is especially noteworthy that the ANEC is not a universal feature of quantum field theories. Yet, in holographic theories with consistent bulk duals, the ANEC is restored as a consequence of deeper geometric and information-theoretic principles. Perhaps most strikingly, this result reverses the typical logic of holography: rather than using boundary data to constrain the bulk, bulk causality is used to derive nontrivial dynamical constraints on the boundary theory. The argument elegantly ties together extremal surfaces, entanglement entropy, and modular flow to derive a physical energy condition from first principles.
Acknowledgments
We wish to thank Prof. Masahiro Hotta for proposing the idea of this review article. The work of N.I. was supported in part by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. JP21H05182, No. 21H05184. The work of N.I. was also supported in part by NSTC of Taiwan Grant Number 114-2112-M-007-025-MY3. The work of A. I. and K. M. was supported in part by MEXT KAKENHI Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. JP21H05182, No. JP21H05186. The work of K. M. was supported in part by JSPS KAKENHI Grant-in-Aid for Scientific Research (C) No. 25K07306. The work of T. N. was supported in part by the JSPS Grant-in-Aid for Scientific Research (B) No. 24K00629, Grant-in-Aid for Scientific Research (A) No. 21H04469, and Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe” No. 21H05182 and No. 21H05190.

### v2

@@
 Also important in any physical theory is the law of the conservation of energy.
-In general relativity, the effects of matter fields on spacetime and conservation laws [^1] The energy-momentum tensor expresses a local conservation of energy and momentum but it does not immediately lead to global properties of energy and momentum of a matter field.
+In general relativity, the effects of matter fields on spacetime and conservation laws [^1] The energy-momentum tensor allows us to express a local conservation of energy and momentum but it does not immediately lead to global properties of energy and momentum of a matter field.
 In order to discuss the positivity and conservation laws of “total energy” of a system, one needs to introduce the notion of asymptotic flatness in order to define an isolated system (see Theorem 2.2.4 and Sec.
@@
 The positivity of energy of a physical material system is, however, not a trivial condition, especially when quantum effects are involved; there is no shortage of examples of negative energy in quantum mechanical systems.
-One of the most well-known examples is the Casimir effect.
+One of the most well-known examples is the Casimir effect, which is experimentally well-established (see, e.g., Klimchitskaya:2009cw and references therein).
 The Hawking radiation mentioned above also violates the positivity of energy.
-In fact, any locally (pointwisely) defined energy conditions can, in principle, be violated by quantum field effects [^2] It should also be noticed that even though local energy conditions can fail to hold under some quantum effects, they often admit lower bounds, hence satisfying what is called “quantum energy inequality.” See, e.g., Yurtsever:1994wc ; Fewster:2012yh ; Kontou:2020bta for quantum energy inequalities and their relations to the (averaged) energy conditions.
+In fact, any locally (pointwisely) defined energy condition can, in principle, be violated by quantum field effects [^2] It should also be noticed that even though local energy conditions can fail to hold under some quantum effects, they often admit lower bounds, hence satisfying what is called “quantum energy inequality.” See, e.g., Yurtsever:1994wc ; Fewster:2012yh ; Kontou:2020bta for quantum energy inequalities and their relations to the (averaged) energy conditions.
 .
@@
 Therefore the AdS/CFT correspondence can be a powerful tool to prove, for example, some of the energy conditions, as performed in Kelly:2014mra .
-The purpose of this article is to review the basics of energy conditions in general relativity, their uses in gravitational physics, their recent applications in quantum field theory and holography, and their relations to quantum information.
-This article is not intended to cover the entire range of relevant subjects, to be comprehensive, or to provide a complete list of existing literature.
+The purpose of this article is to review the basics of energy conditions in general relativity, their uses in gravitational physics, their recent applications in quantum field theory, and relations to quantum information.
+This article is not intended to cover the entire range of relevant subjects, to be comprehensive, or provide a complete list of existing literature.
 Instead, we will introduce key ideas and methods of the uses of energy conditions and quantum information, by focusing on some specific topics.
-In the first part, we mainly focus on the four locally defined energy conditions (WEC, DEC, SEC, and NEC), with a particular emphasis on their role in several theorems in general relativity.
+In the first part, we mainly focus on the four locally defined energy conditions (WEC, DEC, SEC, NEC), with a particular emphasis on their role in several theorems in general relativity.
 In the second part, we first discuss how quantum effects lead to the violation of these locally defined energy conditions.
@@
 In particular, we explain the black hole information paradox in more detail in the context of the AdS/CFT correspondence, introducing the recent advancements such as the holographic entanglement entropy or the Ryu-Takayanagi formula Ryu:2006bv ; Ryu:2006ef , quantum extremal surface (QES) Faulkner:2013ana ; Engelhardt:2014gca , and Island formula Almheiri:2019hni ; Penington:2019npb ; Almheiri:2019psf .
-Then, we discuss the holographic proof of ANEC and its generalization by focusing on a role of causality in the AdS/CFT correspondence.
+Then, we discuss the holographic proof of ANEC and its generalization by focusing on the role of causality in the AdS/CFT correspondence.
 Section 5 is devoted to summary and discussions.
@@
 Our main focus is on the uses of four main, locally (pointwisely) defined energy conditions, but we also discuss their violations and non-local generalizations at the end of this section.
+Unless specifically stated, the assertions, theorems, and formulas discussed in this section assume a four-dimensional spacetime.
+Most of these formulas and theorems remain valid in general dimensions, either without modification or with obvious dimensional extensions.
 2.1 The Einstein equations
-Spacetime geometry can be described in terms of the pair of the metric tensor $g_{\mu\nu}$ with Lorentzian signature and a manifold $M$.
+Spacetime geometry can be described in terms of the pair of a metric tensor $g_{\mu\nu}$ with Lorentzian signature and a connected manifold $M$.
 The Einstein equations are a set of second-order differential equations for the metric tensor, expressed as
 $R_{\mu\nu}-\dfrac{1}{2}Rg_{\mu\nu}=8\pi GT_{\mu\nu}\,$, (1)
-where $R_{\mu\nu}$ and $R$ are the Ricci tensor and scalar curvature, respectively, constructed from the Riemann tensor333 The Riemann tensor is defined as follows: $R^{\lambda}{}_{\sigma\mu\nu}=\partial_{\mu}\Gamma^{\lambda}_{\sigma\nu}-\partial_{\nu}\Gamma^{\lambda}_{\sigma\mu}+\Gamma^{\lambda}_{\alpha\mu}\Gamma^{\alpha}_{\sigma\nu}-\Gamma^{\lambda}_{\alpha\nu}\Gamma^{\alpha}_{\sigma\mu}\,$, (2) where $\Gamma^{\lambda}_{\mu\nu}=\dfrac{1}{2}g^{\lambda\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ is the Christoffel symbol and $R_{\mu\nu}:=R^{\alpha}{}_{\mu\alpha\nu}$ and $R:=g^{\mu\nu}R_{\mu\nu}$.
+where $R_{\mu\nu}$ and $R$ are the Ricci tensor and scalar curvature, respectively, constructed from the Riemann tensor [^3] The Riemann tensor is defined as follows: $R^{\lambda}{}_{\sigma\mu\nu}=\partial_{\mu}\Gamma^{\lambda}_{\sigma\nu}-\partial_{\nu}\Gamma^{\lambda}_{\sigma\mu}+\Gamma^{\lambda}_{\alpha\mu}\Gamma^{\alpha}_{\sigma\nu}-\Gamma^{\lambda}_{\alpha\nu}\Gamma^{\alpha}_{\sigma\mu}\,$, (2) where $\Gamma^{\lambda}_{\mu\nu}=\dfrac{1}{2}g^{\lambda\sigma}(\partial_{\mu}g_{\sigma\nu}+\partial_{\nu}g_{\sigma\mu}-\partial_{\sigma}g_{\mu\nu})$ is the Christoffel symbol and $R_{\mu\nu}:=R^{\alpha}{}_{\mu\alpha\nu}$ and $R:=g^{\mu\nu}R_{\mu\nu}$.
 For geometric quantities, we follow the convention and notations of Wald’s book Wald84 .
@@
 Then, from the local law of conservation above, one obtains general relativistic versions of the conservation of mass and the Euler equation.
-More generally, when an action $S_{M}$ for a matter field is given, the energy-momentum tensor $T_{\mu\nu}$ is obtained by taking a variation of $S_{M}$ with respect to the metric tensor:
+In general, when an action $S_{M}$ for a matter field is given, the energy-momentum tensor $T_{\mu\nu}$ is obtained by taking a variation of $S_{M}$ with respect to the metric tensor:
 $T_{\mu\nu}:=\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{M}}{\delta g^{\mu\nu}}\,$, (5)
@@
 where $V(\phi)$ denotes the potential. Then, the energy-momentum tensor is
-$T^{(\phi)}_{\mu\nu}=\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{\phi}}{\delta g^{\mu\nu}}=\nabla_{\mu}\phi\nabla_{\nu}\phi-\dfrac{1}{2}g_{\mu\nu}\left[g^{\alpha\beta}\nabla_{\alpha}\phi\nabla_{\beta}\phi+2V(\phi)\right]\,.$ (7)
+$T^{(\phi)}_{\mu\nu}$ $=$ $\dfrac{-2}{\sqrt{-g}}\dfrac{\delta S_{\phi}}{\delta g^{\mu\nu}}$ (7)
+$=$ $\nabla_{\mu}\phi\nabla_{\nu}\phi-\dfrac{1}{2}g_{\mu\nu}\left[g^{\alpha\beta}\nabla_{\alpha}\phi\nabla_{\beta}\phi+2V(\phi)\right]\,.$
 2.2 Pointwise energy conditions
@@
 The positivity of energy can, in principle, depend on an observer who measures energy and momentum.
-Besides the positivity, also important is the causality; in relativistic theory, energy and momentum associated with a particle form a single four-vector (four-momentum $p^{\mu}$), which is required to be timelike or null.
+Besides the positivity, also important is the causality; in relativistic theory, energy and momentum associated with a particle form a single four-vector (four-momentum $p^{\mu}$), which is required to be timelike or null, and also be future pointing.
 Depending upon the choice of an observer, one may consider different types of energy conditions on the given energy-momentum tensor, as will be seen below.
@@
     The first reason is that if no restrictions are imposed on the energy-momentum tensor (other than it satisfying conservation laws), virtually any metric could be considered a solution to the Einstein equations.
-This is because for any given metric, one can simply view the Einstein equations as the defining equations for an energy-momentum tensor.
+This is because for any given metric, one can simply view the Einstein equations as the defining equations for an energy-momentum tensor [^5] This procedure for defining an energy-momentum tensor is sometimes known as the “Synge G-method” or the “Nariai-method” (see, e.g., Ellis:2023css ).
+.
 However, such an arbitrary definition would not represent a physically meaningful spacetime.
@@
 In the following, for simplicity we characterize each of WEC, DEC, SEC, NEC, only in terms of Type I energy-momentum tensor.
-{itembox}
-[l]Definition 2.2.1 (The Weak Energy Condition (WEC))
-$T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
+  Definition 2.2.1 (Weak Energy Condition)   $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
   • For Type I case $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$.
   • This condition simply implies that the energy density measured by any timelike observer should be non-negative and is generally believed to hold for all physically reasonable classical matter fields.
-{itembox}
-[l]Definition 2.2.2 (The Dominant Energy Condition (DEC)) For arbitrary future-directed timelike vector $\xi^{\mu}$, $T^{\mu}{}_{\nu}\xi^{\nu}$ becomes a future-directed timelike or null vector.
+  Definition 2.2.2 (Dominant Energy Condition)   For arbitrary future-directed timelike vector $\xi^{\mu}$, $T^{\mu}{}_{\nu}\xi^{\nu}$ becomes a future-directed timelike or null vector.
   • For Type I case $\rho\geqslant|p_{i}|\geqslant 0$.
@@
   • DEC implies that the local energy density is non-negative, and the local energy flow vector is non-spaclike. This ensures the following conservation law:
-    {itembox}
-    [l]Proposition 2.2.3 (Conservation law) Suppose the energy-momentum tensor for matter fields $T_{\mu\nu}$ satisfies the local conservation $\nabla_{\nu}T^{\mu\nu}=0$ and the DEC.
+      Proposition 2.2.3 (Conservation law)   Suppose the energy-momentum tensor for matter fields $T_{\mu\nu}$ satisfies the local conservation $\nabla_{\nu}T^{\mu\nu}=0$ and the DEC.
 If $T_{\mu\nu}=0$ on a closed, achronal hypersurface $\Sigma$, then $T_{\mu\nu}=0$ everywhere in the domain of dependence $D(\Sigma)$.
@@
   • Another important consequence of the DEC is the positivity of the total energy of an isolated system.
-In general relativity, an isolated system is defined in terms of asymptotic flatness, and the total energy defined at spatial infinity is called the “Arnowitt-Deser-Misner (ADM)” energy Arnowitt:1962hi , whose positivity was shown by Schoen-Yau Schon:1981vd , Witten Witten:1981mf : {itembox}[l]Theorem 2.2.4 (Positive energy theorem) The total energy is positive for initial data on an asymptotically flat Cauchy surface, provided the DEC holds.
+In general relativity, an isolated system is defined in terms of asymptotic flatness, and the total energy defined at spatial infinity is called the “Arnowitt-Deser-Misner (ADM)” energy Arnowitt:1962hi , whose positivity was shown by Schoen-Yau Schon:1981vd , Witten Witten:1981mf :
+      Theorem 2.2.4 (Positive energy theorem)   The total energy is positive for initial data on an asymptotically flat Cauchy surface, provided the DEC holds.
     • As the total energy of a given isolated system, the ADM energy is a constant.
@@
     • This theorem was generalized to include black holes and also for the asymptotically AdS case by Gibbons et al Gibbons:1982jg .
-{itembox}
-[l]Definition 2.2.5 (The Strong Energy Condition (SEC))
-$\left(T_{\mu\nu}-\frac{1}{2}T^{\lambda}{}_{\lambda}g_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\quad\mbox{for arbitrary timelike vector $\xi^{\mu}$.}$
+  Definition 2.2.5 (Strong Energy Condition)   $\left(T_{\mu\nu}-\dfrac{1}{2}T^{\lambda}{}_{\lambda}g_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0$ for arbitrary timelike vector $\xi^{\mu}$.
   • For Type I case $\rho+p_{i}\geqslant 0$, $\rho+\sum_{i=1}^{3}p_{i}\geqslant 0$.
@@
 Since $T_{\mu\nu}-(1/2)T^{\lambda}{}_{\lambda}g_{\mu\nu}\approx V(\phi)g_{\mu\nu}$, as far as $V(\phi)$ is positive, the SEC is violated.
-  • There is no containment relationship between SEC and WEC. For example, for a negative cosmological constant, $\Lambda<0$, SEC is satisfied, but WEC fails to hold.
-{itembox}
-[l]Definition 2.2.6 (The Null Energy Condition (NEC))
-$T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ for arbitrary null vector $k^{\mu}$.
+  • SEC does not imply WEC. For example, for a negative cosmological constant, $\Lambda<0$, SEC is satisfied, but WEC fails to hold.
+  Definition 2.2.6 (Null Energy Condition)   $T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0$ for arbitrary null vector $k^{\mu}$.
   • For Type I case $\rho+p_{i}\geqslant 0$.
@@
 In this sense, this is also called the Null Convergence Condition (NCC).
-  • NEC is equivalent to WEC with a timelike vector $\xi^{\mu}$ replaced by a null vector $k^{\mu}$.
+  • NEC is analogous to WEC, with a timelike vector $\xi^{\mu}$ replaced by a null vector $k^{\mu}$.
 If WEC holds, then by continuity, NEC also holds.
 Similarly, if SEC conditions holds, then for any $\xi^{\mu}$, $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant(1/2)T^{\lambda}{}_{\lambda}\xi_{\mu}\xi^{\nu}$, and hence, by continuity, NEC also holds.
-  • The violation of NEC (NCC) implies the violation of WEC. Suppose NEC is violated. Then, for any future-directed null vector $k^{\mu}$, there exists a constant $\delta>0$ such that
+  • The violation of NEC (NCC) implies the violation of WEC. Suppose NEC is violated. Then, for some future-directed null vector $k^{\mu}$, there exists a constant $\delta>0$ such that
     $T_{\mu\nu}k^{\mu}k^{\nu}=-\delta<0\,.$
-    Let $u^{\mu}$ be an appropriate future-directed timelike vector. Then, $\xi^{\mu}:=k^{\mu}+\epsilon u^{\mu},\,(\epsilon>0)$ can also be a timelike vector. For this timelike vector $\xi^{\mu}$,
+    Let $u^{\mu}$ be an appropriate future-directed timelike vector. Then, $\xi^{\mu}:=k^{\mu}+\epsilon u^{\mu},\,(\epsilon>0)$ is also a timelike vector. For this timelike vector $\xi^{\mu}$,
     $T_{\mu\nu}\xi^{\mu}\xi^{\nu}=-\delta+2\epsilon T_{\mu\nu}k^{\mu}u^{\nu}+O(\epsilon^{2})\,.$
@@
 Table 1: Four locally defined energy conditions.
-Here $\xi^{\mu}$ and $k^{\mu}$ denote any timelike and null vector, respectively, and “causality restriction” means that superluminal flow of energy is not allowed.
-Definition Purpose and effects Type I case
-WEC    $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ positivity restriction    $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$
-DEC $T_{\mu\nu}\xi_{1}^{\mu}\xi_{2}^{\nu}\geqslant 0$    positivity and causality restriction    $\rho\geqslant|p_{i}|\geqslant 0$
+Here $\xi^{\mu}$ and $k^{\mu}$ denote any timelike and null vector, respectively, and “causality” means that superluminal flow of energy is not allowed.
+The timelike vectors $\xi_{1}^{\mu},\>\xi_{2}^{\nu}$ appearing in DEC are assumed to share the same time orientation.
+Definition Effects Type I case
+WEC    $T_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ positivity    $\rho\geqslant 0$, $\rho+p_{i}\geqslant 0$
+DEC $T_{\mu\nu}\xi_{1}^{\mu}\xi_{2}^{\nu}\geqslant 0$    positivity and causality    $\rho\geqslant|p_{i}|\geqslant 0$
 SEC $R_{\mu\nu}\xi^{\mu}\xi^{\nu}\geqslant 0$ timelike focusing   $\rho+\sum_{i}p_{i}\geqslant 0$, $\rho+p_{i}\geqslant 0$
@@
 $\nabla_{\nu}u_{\mu}=\dfrac{1}{3}\theta h_{\mu\nu}+\sigma_{\mu\nu}+\omega_{\mu\nu}\,$,
-$\theta:=h^{\mu\nu}\nabla_{\nu}u_{\mu}=\nabla_{\mu}u^{\mu}\,,\quad\sigma_{\mu\nu}:=\nabla_{(\nu}u_{\mu)}-\dfrac{1}{3}\theta h_{\mu\nu}\,,\quad\omega_{\mu\nu}:=\nabla_{[\nu}u_{\mu]}\,$, (14)
+$\theta:=h^{\mu\nu}\nabla_{\nu}u_{\mu}=\nabla_{\mu}u^{\mu}\,,\quad\sigma_{\mu\nu}:=\nabla_{(\nu}u_{\mu)}-\dfrac{1}{3}\theta h_{\mu\nu}\,$,
+$\omega_{\mu\nu}:=\nabla_{[\nu}u_{\mu]}\,$, (14)
 where $\theta$ is called the expansion of the congruence of $\gamma$, $\sigma_{\mu\nu}$ the shear, and $\omega_{\mu\nu}$ the rotation.
@@
 .
-Let us consider a congruence of timelike geodesic $\gamma$ orthogonal to $3$-dimensional hypersurface $\Sigma$ with vanishing rotation $\omega_{\mu\nu}=0$.
+Let us consider a congruence of timelike geodesics orthogonal to $3$-dimensional hypersurface $\Sigma$ with vanishing rotation $\omega_{\mu\nu}=0$.
 If we denote by $A$ an infinitesimal cross-sectional ($3$-dimensional) area element of the congruence, then
@@
 where $R_{\mu\nu}$ is the Ricci tensor. We can also derive the equations for $\sigma_{\mu\nu}$ by taking the symmetric trace-free part of $u^{\alpha}\nabla_{\alpha}(\nabla_{\nu}u_{\mu})$,
-$u^{\alpha}\nabla_{\alpha}\sigma_{\mu\nu}$ $=$ $-\frac{2}{3}\theta\sigma_{\mu\nu}-\left(\sigma_{\mu\sigma}\sigma_{\nu}{}^{\sigma}-\frac{1}{3}\sigma_{\alpha\beta}\sigma^{\alpha\beta}h_{\mu\nu}\right)$ (17)
-$-C_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}+\frac{1}{2}\left(h_{\mu}{}^{\alpha}h_{\nu}{}^{\beta}-\frac{1}{3}h^{\alpha\beta}h_{\mu\nu}\right)R_{\alpha\beta}\,$,
+$u^{\alpha}\nabla_{\alpha}\sigma_{\mu\nu}$ $=$ $-\frac{2}{3}\theta\sigma_{\mu\nu}-\left(\sigma_{\mu\sigma}\sigma_{\nu}{}^{\sigma}-\frac{1}{3}\sigma_{\alpha\beta}\sigma^{\alpha\beta}h_{\mu\nu}\right)-C_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}+\frac{1}{2}\left(h_{\mu}{}^{\alpha}h_{\nu}{}^{\beta}-\frac{1}{3}h^{\alpha\beta}h_{\mu\nu}\right)R_{\alpha\beta}\,$, (17)
 where
-$C_{\mu\alpha\nu\beta}:=R_{\mu\alpha\nu\beta}+g_{\mu[\beta}R_{\nu]\alpha}+g_{\alpha[\nu}R_{\beta]\mu}+(1/3)Rg_{\mu[\nu}g_{\beta]\alpha}$ (18)
+$C_{\mu\alpha\nu\beta}$ $:=$ $R_{\mu\alpha\nu\beta}+g_{\mu[\beta}R_{\nu]\alpha}+g_{\alpha[\nu}R_{\beta]\mu}+\dfrac{1}{3}Rg_{\mu[\nu}g_{\beta]\alpha}$ (18)
 is the Weyl tensor. Note that we omit terms related to $\omega_{\mu\nu}$ since $u^{\mu}$ is assumed to be hypersurface orthogonal.
@@
 Now we state the following condition.
-{itembox}[l]Definition 2.4.1 (Null generic condition) A spacetime is said to satisfy the null generic condition if every null geodesic (whose tangent vector field is denoted by $k^{\mu}$) possesses at least one point where
-$k_{[\alpha}R_{\mu]\beta\nu[\gamma}k_{\sigma]}k^{\beta}k^{\nu}\neq 0$.
-This condition is equivalent to
-$\mbox{either (i) }\,R_{\mu\nu}k^{\mu}k^{\nu}\neq 0\,,\,\,\,\mbox{or (ii) }\,p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0\,$, (25)
+  Definition 2.4.1 (Null generic condition)   A spacetime is said to satisfy the null generic condition if every null geodesic (whose tangent vector field is denoted by $k^{\mu}$) possesses at least one point where $k_{[\alpha}R_{\mu]\beta\nu[\gamma}k_{\sigma]}k^{\beta}k^{\nu}\neq 0$.
+This condition is equivalent to either (i) or (ii) below holds Wald84 ,
+$\mbox{(i) }\,R_{\mu\nu}k^{\mu}k^{\nu}\neq 0\,,\,\,\mbox{(ii) }\,p_{\mu}{}^{\alpha}p_{\nu}{}^{\beta}C_{\sigma\alpha\rho\beta}k^{\sigma}k^{\rho}\neq 0\,$, (25)
 at at least a single point on any null geodesic $\gamma$ and is interpreted as a non-vanishing tidal force acting on the photon described by $\gamma$.
-Suppose that our null geodesic curve $\gamma$ with the tangent vector $k^{\mu}$ is complete with respect to the affain parameter $\lambda$; namely $\gamma$ can be extended to take arbitrary value of $\lambda$ on the spacetime considered.
+Suppose that our null geodesic curve $\gamma$ with the tangent vector $k^{\mu}$ is complete with respect to the affine parameter $\lambda$; namely $\gamma$ can be extended to take arbitrary value of $\lambda$ on the spacetime considered.
 Then, from the above observation, it is inevitable that under the generic conditions and the NEC, the expansion eventually negatively diverges $\theta\rightarrow$-$\infty$.
@@
 Thus, we arrive at the following result (see also Figure 2):
-{itembox}
-[l]Theorem 2.4.2 (Null focusing) Under the NEC, any hypersurface-orthogonal null geodesic congruence is focusing:
-$\dfrac{{\rm d}\theta}{{\rm d}\lambda}\leqslant 0$.
-Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\lambda=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the affine length
-$\lambda\leqslant\dfrac{2}{|\theta_{0}|}$.
+  Theorem 2.4.2 (Null focusing)   Under the NEC, any hypersurface-orthogonal null geodesic congruence is focusing: $\dfrac{{\rm d}\theta}{{\rm d}\lambda}\leqslant 0$.
+Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\lambda=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the affine length $\lambda\leqslant\dfrac{2}{|\theta_{0}|}$.
 Figure 2: A null geodesic congruence whose expansion takes initially negative value $\theta_{0}<0$ converges, and within the affine length $\lambda\leqslant 2/|\theta_{0}|$, the cross-section area $A$ of the congruence vanishes, where by definition $\theta\rightarrow$-$\infty$.
@@
 Then, in general, $q$ is said to be conjugate to $p$.
-Then, one can claim that under the null generic condition and NEC, $\gamma$ must possess a pair of conjugate points (see, proposition 9.3.7 Wald).
+Then, one can claim that under the null generic condition and NEC, $\gamma$ must possess a pair of conjugate points (see, proposition 9.3.7 Wald Wald84 ).
 We can also consider a congruence of null geodesics with tangent vector field $k^{\mu}$ which are orthogonal to a $2$-dimensional spacelike surface $S$.
@@
 Since for a timelike geodesics, the Raychaudhuri equation contains the term $R_{\mu\nu}u^{\mu}u^{\nu}$, we use, instead of NEC, the SEC as it implies, $R_{\mu\nu}u^{\mu}u^{\nu}=8\pi G(T_{\mu\nu}-(1/2)Tg_{\mu\nu})u^{\mu}u^{\nu}\geqslant 0$ via the Einstein equations.
-{itembox}
-[l]Definition 2.4.3 (Timelike generic condition) A spacetime is said to satisfy the timeilke generic condition if every timelike geodesic (whose tangent vector field is denoted by $u^{\mu}$) possesses at least one point where
-$R_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}\neq 0$.
+  Definition 2.4.3 (Timelike generic condition)   A spacetime is said to satisfy the timelike generic condition if every timelike geodesic (whose tangent vector field is denoted by $u^{\mu}$) possesses at least one point where $R_{\mu\alpha\nu\beta}u^{\alpha}u^{\beta}\neq 0$.
 Consider two nearby freely falling particles with the unit $4$-velocity vector fields $u^{\mu}=(\partial/\partial\tau)^{\mu}$, and let $\eta^{\mu}$ be the deviation vector (Jacobi field) between the two.
@@
 Thus, the above generic condition implies that a non-vanishing tidal force acts on every freely falling test particle.
-{itembox}
-[l]Theorem 2.4.4 (Timelike focusing) Under the SEC, any hypersurface-orthogonal timelike geodesic congruence is focusing:
-$\dfrac{{\rm d}\theta}{{\rm d}\tau}\leqslant 0$.
-Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\tau=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the proper time
-$\tau\leqslant\dfrac{3}{|\theta_{0}|}$.
+  Theorem 2.4.4 (Timelike focusing)   Under the SEC, any hypersurface-orthogonal timelike geodesic congruence is focusing: $\dfrac{{\rm d}\theta}{{\rm d}\tau}\leqslant 0$.
+Furthermore, if the congruence has initially a negative expansion (say, $\theta_{0}<0$ at $\tau=0$), then it continues to converge and reaches a point where $\theta\rightarrow$-$\infty$ within a finite value of the proper time $\tau\leqslant\dfrac{3}{|\theta_{0}|}$.
 Similar to the null geodesic case, we can say that a point $q\in\gamma$ is conjugate to a point $p$ (or to $3$-dimensional hypersurface $\Sigma$) when the congruence is emanating from the point $p\in\gamma$ (or orthogonal to $\Sigma$).
-We can also claim that under the timeilke generic condition and the SEC, $\gamma$ must possess a pair of conjugate points.
-(See, proposition 9.3.1 Wald).
-An important aspect of the occurrence of a conjugate point $q$ for timelike geodesics is that a timelike geodesic $\gamma$ fails to be a local maximum of proper time between $q$ and $p$ (or $\Sigma$).
+We can also claim that under the timelike generic condition and the SEC, $\gamma$ must possess a pair of conjugate points.
+(See, proposition 9.3.1 Wald Wald84 ).
+An important aspect of the occurrence of a conjugate point $q$ for timelike geodesics is that a timelike geodesic $\gamma$ fails to be a local maximum of proper time between the initial point $p$ (or $\Sigma$) and any point beyond $q$ HE73 ; Wald84 .
 2.5 Singularity theorems and energy conditions
@@
 In this section, we will discuss how the energy conditions are used in the proof of the theorems.
-A spacetime singularity represents some extreme situation where general relativistic description of spacetime as a manifold breaks down due to infinite energy density or infinite spacetime curvature.
+A spacetime singularity represents some extreme situation where the general relativistic description of spacetime as a manifold breaks down due to infinite energy density or infinite spacetime curvature.
 This, in turn, implies that spacetime singularity should not be considered as part of the spacetime manifold.
-This leads to the idea that if a freely falling observer reaches a singularity, her/his world-line suddenly ends at a finite proper time and cannot be extended any further.
+This leads to the idea that if a freely falling observer reaches a singularity, her/his world-line cannot be continued beyond a certain proper time.
 Such a world-line (i.e., endless but never attains arbitrary large values in its proper time) can be described as an inextendible, incomplete timelike geodesic curve.
@@
 […]
-{itembox}
-[l]Proposition 2.5.1 Any achronal null geodesic (hence, any achronal null hypersurface) does not admit a pair of conjugate points on it.
+  Proposition 2.5.1   Any achronal null geodesic (hence, any achronal null hypersurface) does not admit a pair of conjugate points on it.
 Figure 3 illustrates what happens when a conjugate point appears.
-Figure 3: Null geodesics orthogonally emanating from $S$.
+Figure 3: Null geodesics emanating from a subset $S$ of $M$.
 One such null geodesic $\gamma$ is passing through $p$, where its expansion changes sign from positive to negative, and then has a conjugate point $r$ in $J^{+}(p)$.
@@
 See also Figure 4 below.
-An important example of achronal set is given by, {itembox}[l]Proposition 2.5.2 The boundary $\partial J^{+}(S)$ of the causal future of $S$ is a closed, $3$-dimensional, achronal submanifold.
-Let $\Sigma$ be any closed achronal set.
+An important example of achronal set is given by,
+  Proposition 2.5.2   The boundary $\partial J^{+}(S)$ of the causal future of $S$ is a closed, $3$-dimensional, achronal submanifold.
+Let $M$ be connected and let $\Sigma$ be any closed achronal set in $M$.
 The future domain of dependence of $\Sigma$, denoted $D^{+}(\Sigma)$, is defined by the set of all points $p\in M$ such that every past-directed inextendible causal curve from $p$ intersects $\Sigma$.
@@
 $L[\lambda]=\int_{p}^{q}{\rm d}t\,(-g_{\mu\nu}t^{\mu}t^{\nu})^{1/2}\,.$ (30)
-Then, one has {itembox}[l]Proposition 2.5.3 (Existence of maximum length curve) Let $N$ be globally hyperbolic (possibly as a subset of $M$).
+Then, one has
+  Proposition 2.5.3 (Maximum length curve)   Let $N$ be globally hyperbolic (possibly as a subset of $M$).
 Then, for any two points $p,q\in N$ with $q\in J^{+}(p)$, there exists a causal geodesic curve $\mu$ from $p$ to $q$ which attains the maximum length among all causal curves from $p$ to $q$.
@@
 As mentioned briefly just below the focusing theorems in the previous subsection, if a timelike geodesic connecting two points $p,q\in M$ contains a pair of conjugate points between the two, then there exists another causal curve whose length (proper time) is larger than the original timelike geodesic.
-The Lorentzian distance function $d:M\times M\rightarrow{\mathbb{R}}\cup\infty$ between two points $p,q\in M$ is defined by $d(p,q):=\sup\{L(\lambda):\lambda\in C(p,q)\}$.
+The Lorentzian distance function $d:M\times M\rightarrow{\mathbb{R}}\cup\{\infty\}$ between two points $p,q\in M$ is defined by $d(p,q):=\sup\{L(\lambda):\lambda\in C(p,q)\}$.
 If $p,q\in I^{+}(p),\>p\neq q$, then $d(p,q)>0$, and if $q\in E^{+}(p)$ or not causally related, then $d(p,q)=0$.
@@
 Similarly, if a null geodesic connecting two points $p,q\in M$ contains a pair of conjugate points on it, then it cannot be achronal since $p,q$ can be then connected by some timelike curve.
-In other words, it fails to be the fastest causal curve.
 Combining this observation with Prop.
 2.5.3, we have the following:
-{itembox}
-[l]Proposition 2.5.4 Any causal geodesic curve which connects two points $p$ and $q$, and which attains the maximum length between them cannot have a pair of conjugate points between $p$ and $q$.
+  Proposition 2.5.4   Any causal geodesic curve which connects two points $p$ and $q$, and which attains the maximum length between them cannot have a pair of conjugate points between $p$ and $q$.
 The singularity theorems predict the existence of an inextendible, incomplete causal (either timelike or null) geodesic curve, under the following three conditions concerning:
@@
 After demonstrating contradictory consequences, instead of claiming that the assumption of the geodesic completeness fails, one may rather be able to conclude that the assumption of the global hyperbolicity must fail.
-In this respect, it is worth noting that under the same conditions of Penrose’s theorem, except the assumption of the global hyperbolicity, one can construct geodesically complete spacetimes, among which most well-known example is the Bardeen’s regular black hole.
-    For this reason, after Penrose’s theorem Hawking:1970zqf , Hawking and Penrose proposed the improved version of the singularity theorem, in which the unwanted assumption of the global hyperbolicity is dropped, and instead, the chronology condition—which demands that no closed timelike curves exist—is imposed.
+In this respect, it is worth noting that under the same conditions of Penrose’s theorem, except the assumption of the global hyperbolicity, one can construct geodesically complete spacetimes, among which most well-known example is the Bardeen’s regular black hole 2767662 .
+    For this reason, after Penrose’s theorem Penrose:1964wq , Hawking and Penrose Hawking:1970zqf proposed the improved version of the singularity theorem, in which the unwanted assumption of the global hyperbolicity is dropped, and instead, the chronology condition—which demands that no closed timelike curves exist—is imposed.
 This is a reasonable condition as to consider a physically realistic universe.
@@
     Strong gravity.
-    This property is described by the requirement of the trapped set defined as follows: {itembox}[l]Definition 2.5.5 (Future trapped set) A closed achronal subset $S$ is said to be a future trapped set if $E^{+}(S)$ is compact.
+    This property is described by the requirement of the trapped set defined as follows:
+      Definition 2.5.5 (Future trapped set)   A closed achronal subset $S$ is said to be a future trapped set if $E^{+}(S)$ is compact.
 The past trapped set is defined analogously.
@@
 It is important to note, however, that in a closed universe, one cannot generally distinguish between ingoing and outgoing.
-    In the context of gravitational collapse and black hole physics, the most important example of a trapped set is the closed trapped surface, which is a compact $2$-dimensional surface $\cal T$} such that the expansions $\theta_{\pm}$ of the two families of future (or past) null geodesic congruence orthogonal to $\cal T$} both are negative.
+Note also that in the static Einstein universe, every Cauchy surface $\Sigma$ itself is in fact a trapped set since $E^{+}(\Sigma)=\Sigma$ is compact.
+    In the context of gravitational collapse and black hole physics, the most important example of a trapped set is the closed trapped surface, which is a compact, orientable, and spacelike $2$-dimensional surface $\cal T$} such that the expansions $\theta_{\pm}$ of the two families of future (or past) null geodesic congruence orthogonal to $\cal T$} both are negative.
 An example is any $2$-sphere of constant area radius inside the Schwarzschild (horizon) radius $r_{h}:=2G{\cal M}$.
-To be concrete, let us express the Schwarzschild metric666 In the standard coordinates $x^{\mu}=(t,r,\vartheta,\varphi)$, the Schwarzschild metric is given as ${\rm d}s^{2}=-f(r)\,{\rm d}t^{2}+f^{-1}(r)\,{\rm d}r^{2}+r^{2}({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (33) where $f(r):=1-2G{\cal M}/r$, ${\cal M}=const.$.
+To be concrete, let us express the Schwarzschild metric [^7] In the standard coordinates $x^{\mu}=(t,r,\vartheta,\varphi)$, the Schwarzschild metric is given as ${\rm d}s^{2}=-f(r)\,{\rm d}t^{2}+f^{-1}(r)\,{\rm d}r^{2}+r^{2}({\rm d}\vartheta^{2}+\sin^{2}\vartheta\,{\rm d}\varphi^{2})\,$, (33) where $f(r):=1-2G{\cal M}/r$, ${\cal M}=const.$.
 In one of the exterior region, U<0,V>0U<0,\>V>0, the two coordinates are related as $U=-e^{-(t-r_{*})/2r_{h}},\>V=e^{(t+r_{*})/2r_{h}}$, where $r_{*}=r+r_{h}\log(r-r_{h})$.
@@
 As discussed before, the null and timelike focusing theorems ensures that any complete causal geodesics inevitably contain a pair of conjugate points.
-Let us state Penrose’s version of the singularity theorem Penrose:1964wq . {itembox}[l]Theorem 2.5.6 (Penrose 1965) Spacetime $(M,g_{\mu\nu})$ cannot be null geodesically complete if:
-  • (i)
-    there is non-compact Cauchy surface $\Sigma$ in $M$,
-  • (ii)
-    there is a closed trapped surface $\cal T$} in $M$,
-  • (iii)
-    the NEC holds.
+Let us state Penrose’s version of the singularity theorem Penrose:1964wq .
+  Theorem 2.5.6 (Penrose 1965)   Spacetime $(M,g_{\mu\nu})$ cannot be null geodesically complete if: (i) there is non-compact Cauchy surface $\Sigma$ in $M$, (ii) there is a closed trapped surface $\cal T$} in $M$, (iii) the NEC holds.
 The proof goes roughly as follows.
@@
 $\Box$
-Next, consider Hawking and Penrose’s improved version of the theorem Hawking:1970zqf , in which the assumption of the global hyperbolicity (i.e., the existence of a Cauchy surface) is eliminated: {itembox}[l]Theorem 2.5.7 (Hawking-Penrose 1970) The following three conditions cannot all hold:
-  • (i)
-    the chronology condition holds,
-  • (ii)
-    there is a trapped set $\cal S$},
-  • (iii)
-    every inextendible causal geodesic contains a pair of conjugate points.
+Next, consider Hawking and Penrose’s improved version of the theorem Hawking:1970zqf , in which the assumption of the global hyperbolicity (i.e., the existence of a Cauchy surface) is eliminated:
+  Theorem 2.5.7 (Hawking-Penrose 1970)   The following three conditions cannot all hold: (i) the chronology condition holds, (ii) there is a trapped set $\cal S$}, (iii) every inextendible causal geodesic contains a pair of conjugate points.
   • This is an alternative expression HE73 of the Hawking-Penrose theorem Hawking:1970zqf .
@@
 […]
-{itembox}
-[l]Proposition 2.6.1 If the NEC holds, then a closed trapped surface $\cal T$}, if it exists, cannot intersect $J^{-}(\mathscr{I}^{+})$.
-  • Propositoin 9.2.1 HE73 .
+  Proposition 2.6.1   If the NEC holds, then a closed trapped surface $\cal T$}, if it exists, cannot intersect $J^{-}(\mathscr{I}^{+})$.
+  • Proposition 9.2.1 HE73 .
   • This implies that a closed trapped surface $\cal T$} cannot be seen from an ideal distant observer $\mathscr{I}^{+}$, and therefore indicates that $\cal T$} should be hidden inside the black hole region $B$.
@@
 The proof goes roughly as follows.
-To derive a contradiction, suppose ${\cal T}\in J^{-}(\mathscr{I}^{+})$.
+To derive a contradiction, suppose ${\cal T}\cap J^{-}(\mathscr{I}^{+})\neq\emptyset$.
 Then, a null geodesic generator $\mu$ of $\partial J^{+}({\cal T})$ has an intersection with $\mathscr{I}^{+}$ in $\tilde{M}$.
@@
 Applying a similar type of argument to the event horizon itself, one obtains the following result:
-{itembox}
-[l]Theorem 2.6.2 (Area theorem: The second law) Under the NEC, the cross-sectional area of the event horizon never decreases toward future.
-  • Propositoin 9.2.1 HE73 , Theorem 12.2.6 Wald84 .
+  Theorem 2.6.2 (Area theorem: The second law)   Under the NEC, the cross-sectional area of the event horizon never decreases toward future.
+  • Proposition 9.2.1 HE73 , Theorem 12.2.6 Wald84 .
   • One-quarter of the horizon cross-sectional area can be interpreted as the entropy of the black hole Bekenstein:1973ur (see Sec 4 and Formula 4.3.1), and accordingly the above property of the horizon area increase is phsyically interpreted as the second law of the black hole thermodynamics.
@@
 We also assume that the two Killing vector fields $t^{\mu}$ and $\varphi^{\mu}$ commute.
-An important consequence is that: {itembox}[l]Theorem 2.6.3 (Rigidity theorem) If an asymptotically flat, analytic, stationary spacetime solving the electrovacuum Einstein equations contains a black hole, then the event horizon of the black hole must be a Killing horizon.
+An important consequence is that:
+  Theorem 2.6.3 (Rigidity theorem)   If an asymptotically flat, analytic, stationary spacetime solving the electrovacuum Einstein equations contains a black hole, then the event horizon of the black hole must be a Killing horizon.
 Furthermore, if a stationary black hole is rotating, it must also be axisymmetric.
@@
 $\kappa^{2}=-\dfrac{1}{2}(\nabla^{\mu}\chi^{\nu})\nabla_{\mu}\chi_{\nu}\,.$ (38)
-Although the surface gravity $\kappa$, defined as above, is a local notion, the following result holds: {itembox}[l]Theorem 2.6.4 (The zero-th law) The surface gravity $\kappa$ of a stationary black hole must be a constant over its event horizon.
+Although the surface gravity $\kappa$, defined as above, is a local notion, the following result holds:
+  Theorem 2.6.4 (The zero-th law)   The surface gravity $\kappa$ of a stationary black hole must be a constant over its event horizon.
   • Bardeen-Carter-Hawking (1973) Bardeen:1973gs with DEC.
@@
 […]
-$\delta M=\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}t^{\nu}\,,\quad\delta J=-\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}\varphi^{\nu}\,$, (39)
+$\delta M$ $=$ $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}t^{\nu}\,$,
+$\delta J$ $=$ $-\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}\varphi^{\nu}\,$, (39)
 where $\sigma$ denotes a horizon cross-section with the area element ${\rm d}S$, and $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ the tangent to the horizon null geodesic generator with $\lambda$ being an affine parameter so that $k^{\mu}=\chi^{\mu}{\rm d}v/{\rm d}\lambda=\chi^{\mu}/(\kappa\lambda)=(t^{\mu}+\Omega_{H}\varphi^{\mu})/(\kappa\lambda)$.
-At this stage, we regard $T_{\mu\nu}$ as a small perturbation, which generates non-vanishing expansion and shear via the Raychaudhur equation.
+At this stage, we regard $T_{\mu\nu}$ as a small perturbation, which generates non-vanishing expansion and shear via the Raychaudhuri equation.
 In the linear order of perturbation, neglecting the quadratic terms of the expansion and shear, we can write the Raychaudhuri equation for the horizon null geodesic generators as
-$\dfrac{{\rm d}\theta}{{\rm d}\lambda}\simeq-8\pi GT_{\mu\nu}k^{\mu}k^{\nu}=-\dfrac{8\pi G}{\kappa\lambda}T_{\mu\nu}(t^{\mu}+\Omega_{H}\varphi^{\mu})k^{\nu}\,.$ (40)
-Then, by integrating the Raychaudhuri equation along the horizon generators [^9] By multiplying $\kappa\lambda$ on both side of (40), and taking integration over $\sigma$ and along the null generator, the right-hand side immediately yields $8\pi G(\delta M-\Omega_{H}\delta J)$, while the left-hand side gives $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,(\kappa\lambda)\,\dfrac{{\rm d}\theta}{{\rm d}\lambda}=\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\left[\dfrac{{\rm d}(\lambda\theta)}{{\rm d}\lambda}-\theta\right]=-\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,\theta=-\kappa\delta A\,$, (41) where we have used the fact that before and sufficiently after one has dropped matter fields into the horizon, $\theta=0$ as a stationary state.
-, as well as horizon cross-section $\sigma$, we have {itembox}[l]Theorem 2.6.5 (The first law)
-$\dfrac{\kappa}{8\pi G}\delta A=\delta M-\Omega_{H}\delta J$ (42)
+$\dfrac{{\rm d}\theta}{{\rm d}\lambda}$ $\simeq$ $-8\pi GT_{\mu\nu}k^{\mu}k^{\nu}$ (40)
+$=$ $-\dfrac{8\pi G}{\kappa\lambda}T_{\mu\nu}(t^{\mu}+\Omega_{H}\varphi^{\mu})k^{\nu}\,.$
+Then, by integrating the Raychaudhuri equation along the horizon generators [^10] By multiplying $\kappa\lambda$ on both side of (40), and taking integration over $\sigma$ and along the null generator, the right-hand side immediately yields $8\pi G(\delta M-\Omega_{H}\delta J)$, while the left-hand side gives $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,(\kappa\lambda)\,\dfrac{{\rm d}\theta}{{\rm d}\lambda}$ $=$ $\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\left[\dfrac{{\rm d}(\lambda\theta)}{{\rm d}\lambda}-\theta\right]$ (41) $=$ $-\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,\theta$ $=$ $-\kappa\delta A\,$, where we have used the fact that before and sufficiently after one has dropped matter fields into the horizon, $\theta=0$ as a stationary state.
+, as well as horizon cross-section $\sigma$, we have
+  Theorem 2.6.5 (The first law)   $\dfrac{\kappa}{8\pi G}\delta A=\delta M-\Omega_{H}\delta J$ (42)
   • Bardeen-Carter-Hawking (1973) Bardeen:1973gs .
@@
 To define the notion of MOTS in more general setting, consider a closed, connected $2$-surface ${\cal T}_{0}$.
-Let $k^{\mu}$ be the tangent vector field for an out-going101010 In general, the notion of “out-going” or “outward” direction is not uniquely given.
+Let $k^{\mu}$ be the tangent vector field for an out-going [^11] In general, the notion of “out-going” or “outward” direction is not uniquely given.
 For an asymptotically flat spacetime with a single asymptotic region, if ${\cal T}_{0}$ separates a Cauchy surface $\Sigma$ into two disconnected parts $\Sigma_{\rm in}$ and $\Sigma_{\rm out}$, then “out-going” direction is defined as the one, say $\Sigma_{\rm out}$, that contains the asymptotic region.
@@
 The following theorem determines the topology of MOTS.
-{itembox}
-[l]Theorem 2.6.6 (Topology of horizon cross-section) Under the DEC, the horizon cross-section ${\cal T}_{0}$ of a stationary black hole must be topologically $2$-sphere.
-  • Propositoin 9.3.2 HE73 .
+  Theorem 2.6.6 (Topology of horizon cross-section)   Under the DEC, the horizon cross-section ${\cal T}_{0}$ of a stationary black hole must be topologically $2$-sphere.
+  • Proposition 9.3.2 HE73 .
 The proof consists of two parts.
@@
 By using the Einstein equations and the properties of the MOTS, $\theta_{+}=0$ and $\ell^{\nu}\nabla_{\nu}\theta_{+}\geqslant 0$, one can find the following inequality:
-$\oint_{{\cal T}_{0}}\!\!{\rm d}S\,{\cal R}\geqslant 8\pi T_{\mu\nu}k^{\mu}(-\ell^{\nu})\geqslant 0\,$, (47)
+$\oint_{{\cal T}_{0}}\!\!{\rm d}S\,{\cal R}\geqslant 16\pi G\oint_{{\cal T}_{0}}\!\!{\rm d}S\,T_{\mu\nu}k^{\mu}(-\ell^{\nu})\geqslant 0\,$, (47)
 […]
 Next, let us consider the topology of the black hole exterior region. We first describe two theorems concerning topological constraints on globally hyperbolic, asymptotically flat spacetimes.
-{itembox}
-[l]Theorem 2.6.7 (Topology and singularity) Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime, on which the NEC holds.
+  Theorem 2.6.7 (Topology and singularity)   Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime, on which the NEC holds.
 Suppose that $M$ contains a non-simply connected Cauchy surface $\Sigma$.
@@
 Again under the NEC, we have the following theorem.
-{itembox}[l]Theorem 2.6.8 (Topological censorship) Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime.
+  Theorem 2.6.8 (Topological censorship)   Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime.
 If the NEC holds in $M$, then any causal curve $\gamma$ from a point on $\mathscr{I}^{-}$ to a point on $\mathscr{I}^{+}$ can be continuously deformed to a timelike curve $\gamma_{0}$ which is contained in a simply connected neighborhood of $\mathscr{I}^{+}\cup\mathscr{I}^{-}$.
@@
 Figure 9: The domain of outer communications of an asymptotically flat black hole spacetime.
-Applying Theorem 2.6.8 to $\rm DOC$}, one can find the following constraints on the horizon topology:
-{itembox}
-[l]Theorem 2.6.9 Suppose $(M,g_{\mu\nu})$ is a stationary, asymptotically flat spacetime containing a single asymptotically flat region, whose $\rm DOC$} is globally hyperbolic.
+Applying Theorem 2.6.8 to $\rm DOC$}, one can find the following constraints on the horizon topology.
+  Theorem 2.6.9   Suppose $(M,g_{\mu\nu})$ is a stationary, asymptotically flat spacetime containing a single asymptotically flat region, whose $\rm DOC$} is globally hyperbolic.
 Suppose the NEC holds.
@@
 Then, applying a similar argument of the proof of Theorem 2.6.8, we immediately obtain the following:
-{itembox}
-[l]Proposition 2.6.10 (No traversable wormhole) The traversable wormhole structure (TW) must violate the NEC.
+  Proposition 2.6.10 (No traversable wormhole)   The traversable wormhole structure (TW) must violate the NEC.
 The idea of the proof is as follows.
@@
 Let $t^{\mu}=(\partial/\partial t)^{\mu}$ be a unit timelike vector orthogonal to any homogeneous and isotropic hypersurface.
-Then, due to the symmetry, $\phi$ can only be dependent on $t$, and we have $\left(T^{(\phi)}_{\mu\nu}-({1}/{2})T^{(\phi)}{}^{\lambda}{}_{\lambda}g_{\mu\nu}\right)t^{\mu}t^{\nu}=\dot{\phi}^{2}-V(\phi)$.
+Furthermore, let us assume that $\phi$ depends only on $t$ in accord with the FLRW symmetry.
+Then we have $\left(T^{(\phi)}_{\mu\nu}-({1}/{2})T^{(\phi)}{}^{\lambda}{}_{\lambda}g_{\mu\nu}\right)t^{\mu}t^{\nu}=\dot{\phi}^{2}-V(\phi)$.
 Thus, if $\dot{\phi}^{2}<V$, which can occur for e.g., a massive free scalar field, or inflation/dark energy models, the SEC fails to hold.
@@
 There are a number of known examples in which the expectation value $\langle T_{\mu\nu}\rangle$ violates the locally defined energy conditions, even when their classical counterparts do not violate the locally defined energy conditions.
-In fact, Casimir effects violate all local energy conditions Roman:1986tp .
+In fact, Casimir effects violate all locally defined (pointwise) energy conditions Roman:1986tp .
 For example, consider two parallel reflecting planes orthogonal to $z$-axis in Minkowski spacetime.
 Then the vacuum expectation value of the energy-momentum tensor for a scalar field between the two planes takes the form of $\langle 0|T_{\mu\nu}|0\rangle=-(C^{2}/L^{4})\mathrm{diag}(1,-1,-1,3)$ with $C^{2}$ being a positive constant and $L$ the separation distance between the two planes Birrell:1982ix , for which all the four locally defined energy conditions are violated.
+See also, e.g., Fewster Fewster:2012yh and Section 10.2 of Lobo:2017cay (and references therein) for a more elaborated explanation that all pointwise energy conditions can be violated in quantum field theory.
 While the pointwise energy conditions may fail in some restricted local regions, they may still hold in the rest of the universe.
@@
 For example, in the case of Casimir effect above, although between the two planes, the NEC is violated along $z$-axis, since the two planes supply positive energy, if we integrate $\langle 0|T_{\mu\nu}k^{\mu}k^{\nu}|0\rangle$ along $z$-axis long enough passing through the two planes, we can have a positive energy value.
+In this respect, see also Ford:1978qya for a discussion concerning the second law of thermodynamics and averaged energy conditions.
 In what follows, we consider generalizations of the energy conditions to the averaged forms.
@@
 As a weaker condition than SEC, the averaged strong-energy condition (ASEC) is simply formulated as follows.
-{itembox}
-[l]Definition 2.7.1 (ASEC) For every inextendible timelike geodesic curve with proper time $\tau$ and the tangent $\xi^{\mu}$, the integral along the timelike geodesic,
-$\int{\rm d}\tau\left(T_{\mu\nu}-\dfrac{1}{2}Tg_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\,.$ (50)
+  Definition 2.7.1 (Averaged SEC)   For every inextendible timelike geodesic curve with proper time $\tau$ and the tangent $\xi^{\mu}$, the integral along the timelike geodesic, $\int{\rm d}\tau\left(T_{\mu\nu}-\dfrac{1}{2}Tg_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\,.$ (50)
   • Integrating the local timelike focusing ${\rm d}\theta/{\rm d}\tau\leqslant-R_{\mu\nu}\xi^{\mu}\xi^{\nu}$ in a finite interval $0,\tau]$, and using the Einstein equations, we have
@@
 As for a weaker condition than NEC, we have the averaged null energy condition (ANEC)
-{itembox}
-[l]Definition 2.7.2 (ANEC) For every inextendible null geodesic curve with affine parameter $\lambda$ and the tangent $k^{\mu}$, the integral along the null geodesic,
-$\int{\rm d}\lambda\,T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0\,.$ (52)
+  Definition 2.7.2 (Averaged NEC)   For every inextendible null geodesic curve with affine parameter $\lambda$ and the tangent $k^{\mu}$, the integral along the null geodesic, $\int{\rm d}\lambda\,T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0\,.$ (52)
   • This is the condition used in the original statement of the topological censorship Friedman:1993ty (Theorem 2.6.8).
@@
 This theorem generalizes to the case when $A$ has the boundary on a null hypersurface Casini:2017roe , which allows us to derive the ANEC from the monotonicity of the vacuum modular Hamiltonian Faulkner:2016mzt .
-We also consider the quantum null energy condition (QNEC), which bounds the null-null component of the stress tensor by the second null derivative of entanglement entropy Bousso:2015mna ; Wall:2011kb .
+We also consider QNEC, which bounds the null-null component of the stress tensor by the second null derivative of entanglement entropy Bousso:2015mna ; Wall:2011kb .
 We sketch the derivation of QNEC following Kudler-Flam:2023hkl and leave the complete proof to Ceyhan:2018zfg ; Hollands:2025glm .
@@
 For a density matrix $\rho$ on $\mathcal{H}$, the reduced density matrix $\rho_{A}$ on the Hilbert space $\mathcal{H}_{A}$ of the subsystem $A$ is defined by
-$\rho_{A}\equiv{\rm tr}_{\bar{A}}\left[\,\rho\,\right]\ ,$ (53)
+$\rho_{A}:={\rm tr}_{\bar{A}}\left[\,\rho\,\right]\ ,$ (53)
 […]
@@
 The entanglement entropy $S(\rho_{A})$ of the subsystem $A$ is defined as the von Neumann entropy of the reduced density matrix $\rho_{A}$:
-$S(\rho_{A})\equiv-{\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]\ .$ (56)
+$S(\rho_{A}):=-{\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]\ .$ (56)
 Entanglement entropy is known to satisfy the following inequality known as the strong subadditivity (SSA) Lieb1973cp :
@@
 For a pair of two density matrices $\rho_{A}$ and $\sigma_{A}$ on $\mathcal{H}_{A}$, the (quantum) relative entropy is defined by
-$S(\rho_{A}||\sigma_{A})\equiv{\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]-{\rm tr}_{A}\left[\rho_{A}\,\log\sigma_{A}\right]\ .$ (58)
+$S(\rho_{A}||\sigma_{A}):={\rm tr}_{A}\left[\rho_{A}\,\log\rho_{A}\right]-{\rm tr}_{A}\left[\rho_{A}\,\log\sigma_{A}\right]\ .$ (58)
 Roughly speaking, the relative entropy measures the “distance” between two quantum states $\rho_{A}$ and $\sigma_{A}$. Indeed, it is non-negative
@@
 […]
+Having applications to QFTs in mind, we introduce the full modular Hamiltonian $\hat{K}^{(\rho)}_{A}$ as follows:
+$\hat{K}^{(\sigma)}_{A}$ $:=K^{(\sigma)}_{A}\otimes{\bf 1}_{\bar{A}}-{\bf 1}_{A}\otimes K^{(\sigma)}_{\bar{A}}\$. (66)
+The monotonicity of relative entropy (61) can be expressed as the operator inequality for the full modular Hamiltonian Borchers:1995zg ; Blanco:2013lea :
+$B\subset A\quad\rightarrow\quad\hat{K}^{(\sigma)}_{B}\leq\hat{K}^{(\sigma)}_{A}\$. (67)
 3.2 Modular Hamiltonian in quantum field theory
-The modular Hamiltonian defined by (62) does not have a local expression written by an integral of a local operator in general, unlike a Hamiltonian.
+From now on, we will extend the discussion in the previous section to quantum field theory.
+Let the Hilbert space $\mathcal{H}$ be defined on a Cauchy surface $\Sigma$, and denote by $\mathcal{H}_{A}$ the Hilbert space associated with a subregion $A\subset\Sigma$.
+In a finite-dimensional system, the Hilbert space factorizes as $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{\bar{A}}$, which allows us to define the reduced density matrix $\rho_{A}$, modular Hamiltonian $K_{A}^{(\sigma)}$ and entanglement entropy.
+In quantum field theory, however, such a factorization no longer holds, so these quantities are not well-defined in general.
+Nevertheless, in an algebraic formulation of QFT, one can rigorously define the relative entropy $S(\rho_{A}||\sigma_{A})$ and the full modular Hamiltonian $\hat{K}_{A}^{(\sigma)}$, which generates the modular flow of local operator algebra (see witten2018notes-8d6 for a review).
+In what follows, we will take a less rigorous approach and assume that the Hilbert space factorizes under a suitable ultraviolet regularization, so that the reduced density matrix, modular Hamiltonian, and entanglement entropy can be treated for practical purposes.
+The full modular Hamiltonian defined by (66) does not have a local expression written by an integral of a local operator in general, unlike a Hamiltonian.
 However, when $\sigma$ is the vacuum density matrix $\Omega$, i.e.
-$|\sigma=\Omega=\ket{\Omega}\bra{\Omega}$, the modular Hamiltonian (62) can be “geometric”, which is a local integral of the stress tensor, generating the time evolution.
+$|\sigma=\Omega=\ket{\Omega}\bra{\Omega}$, the full modular Hamiltonian (66) can be “geometric,” which is a local integral of the stress tensor, generating the time evolution.
 A well-known example is the Bisognano-Wichmann theorem Bisognano:1976za .
-Let region $A$ be the right half-space $A=\{t=0,x_{1}\geq 0\}$.
+Let region $A$ be the right half-space $A=\{t=0,x^{1}\geq 0\}$.
 We take the coordinate in Minkowski spacetime as
-$t=r\,\sinh\tau\ ,\qquad x_{1}=r\,\cosh\tau\ .$ (66)
-That is, the coordinate $(r,\tau)$ implies the Minkowski coordinate of the observer who is uniformly accelerating in the Minkowski spacetime, and covers only $|t|\leq x_{1}$, known as the (right) Rindler wedge (see the left panel of figure 12).
-The Rindler wedge is the causal domain (diamond) $D[A]$ of the region $A$.
-The causal domain $D[A]$ is the set of points $p$ such that all causal curves through $p$ intersect $A$, and the information at the point $p$ is completely determined by that of $A$.
-The Hamiltonian $K$ for the observer with respect to the Rindler time $\tau$ is given by
-$K=\int_{A=\{t=0,\,x_{1}\geq 0\}}{\rm d}^{d-1}x\,x_{1}\,T_{tt}\ .$ (67)
-The Bisognano-Wichmann theorem states that the vacuum modular Hamiltonian is the Hamiltonian for the uniformly accelerating observer, i.e.
-$K_{A}^{(\Omega)}=2\pi\,K$.
-The physical interpretation of this theorem is as follows.
-Starting from $x_{1}\geq 0$ at the initial time, the uniformly accelerating observer cannot obtain the information about $x_{1}\leq 0$ at the same time, since even the signal with the speed of light never reaches the observer (for this observer, $x_{1}<0$ is regarded as the inside of the black hole).
-Thus, the system with this observer is described by the vacuum reduced density matrix $\sigma_{A}$, even if the total system is in a pure state (vacuum).
-Furthermore, $\Omega_{A}=e^{-2\pi K}$ implies that the Rindler observer perceives a thermal bath at inverse temperature $\beta=2\pi$.
-Figure 12: [Left] The right Rindler wedge is the causal domain $D[A]$ of the region $A$ as shown in gray color.
-The vacuum modular Hamiltonian coincides with the boost operator $K_{A}^{(\Omega)}=2\pi\,K$ due to the Bisognano-Wichmann theorem.
+$t=r\,\sinh\tau\ ,\qquad x^{1}=r\,\cosh\tau\ .$ (68)
+That is, the coordinate $(r,\tau)$ implies the Minkowski coordinate of the observer who is uniformly accelerating in the Minkowski spacetime, and covers only $|t|\leq x^{1}$, known as the (right) Rindler wedge (see the left panel of figure 12).
+The Rindler wedge is the causal domain (diamond) $D(A)$ of the region $A$.
+The causal domain $D(A)$ is the set of points $p$ such that all causal curves through $p$ intersect $A$, and the information at the point $p$ is completely determined by that of $A$.
+The Hamiltonian $K$ with respect to the Rindler time $\tau$ is given by
+$K=\int_{t=0}{\rm d}^{d-1}x\,x^{1}\,T_{tt}\ .$ (69)
+[…]
+$K_{R}=\int_{A=\{t=0,\,x^{1}\geq 0\}}{\rm d}^{d-1}x\,x^{1}\,T_{tt}\ .$ (70)
+On the other hand, assuming that quantum field theory has a tensor factorized Hilbert space under some UV regularization, the system of the observer may be characterized by the reduced density matrix $\Omega_{A}=e^{-K_{A}^{(\Omega)}}$, where $K_{A}^{(\Omega)}$ is the modular Hamiltonian.
+The Bisognano-Wichmann theorem then implies $K_{A}^{(\Omega)}=2\pi\,K_{R}$, so that $\Omega_{A}=e^{-2\pi K_{R}}$.
+This shows that the Rindler observer perceives a thermal bath at inverse temperature $\beta=2\pi$, even though the total system is in a pure state.
+Figure 12: [Left] The right Rindler wedge is the causal domain $D(A)$ of the region $A$ as shown in gray color.
+The vacuum modular Hamiltonian coincides with the boost operator $K_{A}^{(\Omega)}=2\pi\,K_{R}$ due to the Bisognano-Wichmann theorem.
 [Right] The right Rindler wedge is mapped to the causal domain of the spherical region by the conformal transformation (71).
-While the Bisognano-Wichmann theorem holds in any relativistic quantum field theory, there is another expression for the modular Hamiltonian in conformal field theory (CFT).
+While the Bisognano-Wichmann theorem holds in any relativistic quantum field theory (under certain condition), there is another expression for the modular Hamiltonian in conformal field theory (CFT).
 Let region $A$ be the solid ball of radius $R$ centered at the origin: $A=\{X^{0}=0,\,0\leq r\leq R\}$.
-Here, $r\equiv\sqrt{\sum_{i=1}^{d-1}(X^{i})^{2}}$ is the radius of the spherical coordinates in the constant-time slice.
-The Rindler wedge is mapped to the causal domain $D[A]$ of the solid ball $A$, by conformal transformation, called the CHM map Casini:2011kv (figure 12):
+Here, $r:=\sqrt{\sum_{i=1}^{d-1}(X^{i})^{2}}$ is the radius of the spherical coordinates in the constant-time slice.
+The Rindler wedge is mapped to the causal domain $D(A)$ of the solid ball $A$, by conformal transformation, called the CHM map Casini:2011kv (figure 12):
 $X^{\mu}=\frac{x^{\mu}+(x\cdot x)\,c^{\mu}}{1+2\,(x\cdot c)+(x\cdot x)\,(c\cdot c)}-2\,R^{2}\,c^{\mu}\ ,$ (71)
-where $c^{\mu}=(0,1/(2R),\vec{0}\,)$.
+where $x^{0}=t$ and $c^{\mu}=(0,1/(2R),\vec{0}\,)$.
 This means that, in CFT, the vacuum reduced density matrix of the right half-space is related to $\Omega_{A}$ by a unitary transformation.
 Then, the modular Hamiltonian of $\Omega_{A}$ is given by Casini:2011kv ; Hislop:1981uh
-$K^{(\Omega)}_{A}=2\pi\int_{X^{0}=0,\,0\leq r\leq R}{\rm d}^{d-1}X\,\frac{R^{2}-r^{2}}{2R}\,T_{00}\ .$ (69)
-We can consider the null deformed Rindler wedge as the causal domain $D[A]$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ as shown in figure 13.
-The modular Hamiltonian associated with such a region $A$ is also shown in Casini:2017roe to have a local expression similar to (67)
-$K^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{\gamma_{A}(\bm{x}_{\perp})}^{\infty}{\rm d}x^{+}(x^{+}-\gamma_{A}(\bm{x}_{\perp}))\,T_{++}\left(x^{+},x^{-}=0,\bm{x}_{\perp}\right)\ .$ (70)
-Figure 13: The null deformed Rindler wedge as the causal domain $D[A]$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ .
-Having the applications to QFTs in mind, it is useful to extend the monotonicity of relative entropy (61) to the operator inequality Borchers:1995zg ; Blanco:2013lea :
-$D[B]\subset D[A]\quad\rightarrow\quad\hat{K}^{(\rho)}_{B}\leq\hat{K}^{(\rho)}_{A}\$, (71)
-where $D[A]$ is the causal domain of a spacelike hypersurface $A$, and $\hat{K}^{(\rho)}_{A}$ is the full modular Hamiltonian
-$\hat{K}^{(\rho)}_{A}$ $\equiv K^{(\rho)}_{A}\otimes{\bf 1}_{\bar{A}}-{\bf 1}_{A}\otimes K^{(\rho)}_{\bar{A}}\$. (72)
+$K^{(\Omega)}_{A}=2\pi\int_{X^{0}=0,\,0\leq r\leq R}{\rm d}^{d-1}X\,\frac{R^{2}-r^{2}}{2R}\,T_{X^{0}X^{0}}\ .$ (72)
+It follows that the full modular Hamiltonian $\hat{K}^{(\Omega)}_{A}$ takes the same form as (72) with the integration over the entire time slice $X^{0}=0$.
+While the derivation of $\hat{K}^{(\Omega)}_{A}$ was not fully rigorous, the resulting expression can be shown to hold exactly in the case of a free massless scalar field Longo:2020amm .
+Let us define the light-cone coordinates $x^{\mu}=(x^{+},x^{-},\bm{x}_{\perp})\ ,~x^{\pm}:=t\pm x^{1}$ in Minkowski spacetime.
+We can consider the null deformed Rindler wedge as the causal domain $D(A)$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ as shown in figure 13.
+The full modular Hamiltonian associated with such a region $A$ in quantum field theory is also shown, at a physicist’s level of rigor in Casini:2017roe , to have a local expression similar to (70):
+$\hat{K}^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{-\infty}^{\infty}{\rm d}x^{+}(x^{+}-\gamma_{A}(\bm{x}_{\perp}))\,T_{++}\left(x^{+},x^{-}=0,\bm{x}_{\perp}\right)\ .$ (73)
+We will use this expression in deriving averaged null energy condition on flat space in section 3.4.
+Figure 13: The null deformed Rindler wedge as the causal domain $D(A)$ of a spacelike hypersurface $A$ with $x^{-}\leq 0$ bounded by $\partial A=\{x^{-}=0,\,x^{+}=\gamma_{A}(\bm{x}_{\perp})\}$ .
 3.3 Bekenstein bound
@@
 $S\lesssim R\,E\ ,$ (74)
-where $S$ and $E$ are the entropy of the matter and energy in the region $A$ with typical size $R$.
+where $S$ and $E$ are the entropy of the matter and energy in the region $A$ with typycal size $R$.
 This bound arose through the following thought experiment.
@@
 Consider the reduced density matrices of an excited state and the vacuum state: $\rho_{A},\Omega_{A}$.
-The positivity of the relative entropy (59) and the relation to modular Hamiltonian (63) show
+The positivity of the relative entropy (59) and the relation to the modular Hamiltonian (63) show
 $\Delta S_{A}\leq\Delta K^{(\Omega)}_{A}\$, (75)
 where the $K^{(\Omega)}_{A}$ in the right hand side is the vacuum modular Hamiltonian in CFT (72).
-Let $E$ be the energy of the excited state in the region $A$, then we can estimate $T_{tt}\sim E/R^{d-1}$, since $T_{tt}$ is the energy density.
+Let $E$ be the energy of the excited state in the region $A$, then we can estimate $T_{X^{0}X^{0}}\sim E/R^{d-1}$, since $T_{X^{0}X^{0}}$ is the energy density.
 Integrating over the region $A$, we obtain $\Delta K_{A}^{(\Omega)}\sim R\,E$.
@@
 […]
-where we use the light-cone coordinates $x^{\mu}=(x^{+},x^{-},\bm{x}_{\perp})\ ,~x^{\pm}\equiv t\pm x_{1}$ in Minkowski spacetime (see also Definition 2.7.2).
 The ANEC holds for any QFT in Minkowski spacetime, which was proved in special cases (free theory, two dimensions, holography) Klinkhammer:1991ki ; Wald:1991xn ; Folacci:1992xg ; Verch:1999nt ; Kelly:2014mra , and in interacting theories Faulkner:2016mzt ; Hartman:2016lgu .
 Here, we give a proof using the monotonicity of the relative entropy along the line of Faulkner:2016mzt .
-The modular Hamiltonian (70) for the null deformed Rindler wedge can also be written as
-$K^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{0}^{\infty}{\rm d}x^{+}x^{+}\,T_{++}\left(x^{+}-\gamma_{A}(\bm{x}_{\perp}),\,x^{-}=0,\,\bm{x}_{\perp}\right)\ .$ (76)
-We also consider another region $B$, whose boundary satisfies $\partial B=\{x^{-}=0,x^{+}=\gamma_{B}(\bm{x}_{\perp})\}$ and $\Delta\gamma(\bm{x}_{\perp})\equiv\gamma_{B}(\bm{x}_{\perp})-\gamma_{A}(\bm{x}_{\perp})>0$ (figure 14).
-Then, since the causal diamonds have a relation $D[B]\subset D[A]$, the monotonicity of the full modular Hamiltonian (71) yields the inequality
+We begin with the full modular Hamiltonian (73) for the null deformed Rindler wedge $D(A)$.
+We also consider another region $B$, whose boundary satisfies $\partial B=\{x^{-}=0,x^{+}=\gamma_{B}(\bm{x}_{\perp})\}$ and $\Delta\gamma(\bm{x}_{\perp}):=\gamma_{B}(\bm{x}_{\perp})-\gamma_{A}(\bm{x}_{\perp})>0$ (figure 14).
+In quantum field theory, (67) implies
+$D(B)\subset D(A)\quad\rightarrow\quad\hat{K}^{(\rho)}_{B}\leq\hat{K}^{(\rho)}_{A}\$, (77)
+where $D(A)$ and $D(B)$ are the causal domains of spacelike hypersurfaces $A$ and $B$, respectively. Thus, in the present setup, (73) and (77) yield the inequality
 $\begin{aligned} 0&\leq\hat{K}^{(\Omega)}_{A}-\hat{K}^{(\Omega)}_{B}\\ &=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{-\infty}^{\infty}{\rm d}x^{+}\,\Delta\gamma(\bm{x}_{\perp})\,T_{++}(x^{+},x^{-}=0,\bm{x}_{\perp})\ .\end{aligned}$ (78)
@@
 While ANEC is a kind of global energy conditions, there is a local bound on the null energy density.
-This is called the quantum null energy condition (QNEC) Bousso:2015mna , whose statement is as follows: {itembox}[l]Definition 3.5.1 (QNEC)
+This is called the quantum null energy condition (QNEC) Bousso:2015mna , whose statement is as follows:
 […]
-Here $|]\rho_{A}={\rm tr}_{\bar{A}}\left[\,\ket{\rho}\bra{\rho}\,\right$] and $S(\rho_{A})$ is the entanglement entropy of a spacelike hypersurface $A$ in $x^{-}\leq 0$ whose boundary is anchored on $x^{-}=0,x^{+}=\gamma_{A}(\bm{x}_{\perp})$.
-Note that the second-order null shape derivative of the entanglement entropy in the right-hand side is not necessarily positive.
-ANEC follows from QNEC by integrating both sides of (78) along, assuming that the surface term $\delta S/\delta\gamma$ vanishes.
-QNEC arose in the context of quantum focusing conjecture (QFC) Bousso:2015mna , and was proven in free theory Bousso:2015wca ; Malik:2019dpg , in holographic theories Koeller:2015qmn , and in interacting theories Balakrishnan:2017bjg ; Ceyhan:2018zfg ; Hollands:2025glm .
-[^12] For interacting CFT, QNEC is conjectured to be saturated Leichenauer:2018obf ; Balakrishnan:2019gxl .
-QNEC also has applications to quenches Mezei:2019sla ; Kibe:2024icu ; Kibe:2025cqc , and renormalization group flow Casini:2023kyj .
-Let us give a sketch of a proof of QNEC (our proof mainly follows that of Kudler-Flam:2023hkl ).
-We will use the expression of the relative entropy (63) and the vacuum modular Hamiltonian (70).
+Let us give a sketch of a physicist’s derivation of QNEC (we closely follows that of Kudler-Flam:2023hkl ).
+We will use the expression of the relative entropy (63) and the vacuum modular Hamiltonian corresponding to (73)
+$K^{(\Omega)}_{A}=2\pi\int{\rm d}^{d-2}\bm{x}_{\perp}\int_{\gamma_{A}(\bm{x}_{\perp})}^{\infty}{\rm d}x^{+}(x^{+}-\gamma_{A}(\bm{x}_{\perp}))\,T_{++}\left(x^{+},x^{-}=0,\bm{x}_{\perp}\right)\ .$ (80)
 By taking $\sigma$ in (63) to be the vacuum reduced density matrix, the second-order null shape derivative of the relative entropy becomes
@@
 $\frac{\delta^{2}S(\rho_{A}||\Omega_{A})}{\delta\gamma_{A}(\bm{x}_{\perp})^{2}}\geq 0\ .$ (82)
-This positivity is guaranteed by the “ant-formula”:
+This inequality is mathematically well-defined while the original QNEC (79) is not as the latter involves entanglement entropy explicitly. (82) is guaranteed by the “ant-formula”:
 […]
@@
 4.1 Black hole entropy and holography
-Although our main focus is on the energy conditions, a deeper understanding of their quantum extensions is incomplete without addressing holography.This subsection introduces the holographic principle, which—though seemingly peripheral—sheds crucial light on the emergence of energy, entropy, and spacetime itself in quantum gravity.
+Although our main focus is on the energy conditions, a deeper understanding of their quantum extensions is incomplete without addressing holography.
+This subsection introduces the holographic principle, which—though seemingly peripheral—sheds crucial light on the emergence of energy, entropy, and spacetime itself in quantum gravity.
 In particular, black holes provide a striking arena where these ideas converge and force us to reconsider basic physical notions.
@@
 Hawking’s discovery Hawking:1974rv ; Hawking:1975vcx that black holes emit blackbody radiation implies that they possess a well-defined temperature.
-From the laws of black hole mechanics, first formulated by Bekenstein Bekenstein:1973ur and Hawking, a consistent thermodynamic description emerged:
+Building on Bekenstein’s insight that horizon area encodes entropy Bekenstein:1973ur , the four laws of black hole mechanics were formulated by Bardeen, Carter, and Hawking Bardeen:1973gs , yielding a consistent thermodynamic description:
 $T_{H}$ $=\dfrac{\hbar\kappa}{2\pi k_{\rm B}}\quad\mbox{(Hawking temperature)}$ (92)
-$S_{\rm BH}$ $=\frac{A}{4G\hbar}\quad\mbox{(Bekenstein-Hawking entropy)}$ (91)
-where $T_{\rm H}$ is called the Hawking temperature given by the surface gravity $\kappa$, and $S_{\rm BH}$ called the Bekenstein-Hawking entropy formula, proportional to the area $A$ of a cross-section of the event horizon as seen in Sec.
+which is called the Hawking temperature given by the surface gravity $\kappa$, and
+  Formula 4.1.1 (Bekenstein-Hawking entropy)   $S_{\text{BH}}=\frac{k_{\rm B}A}{4G\hbar}\,.$ (93)     This is called the Bekenstein-Hawking entropy formula, proportional to the area $A$ of a cross-section of the event horizon as seen in Sec.
 2.6.
@@
 Since black holes are believed to be the densest and most entropic objects in nature, they offer a natural upper bound on the information content in a given region.
-This inspired the conjecture that the fundamental theory underlying quantum gravity is holographic: {itembox}[l]Conjecture 4.1.1 (Holographic principle) The complete description of a gravitational system in a bulk spacetime is equivalently formulated as a lower-dimensional, non-gravitational quantum field theory living on its boundary.
+This inspired the conjecture that the fundamental theory underlying quantum gravity is holographic:
+  Conjecture 4.1.2 (Holographic principle)   The complete description of a gravitational system in a bulk spacetime is equivalently formulated as a lower-dimensional, non-gravitational quantum field theory living on its boundary.
 A concrete and precise realization of this principle was proposed by Maldacena Maldacena:1997re in 1997, now famously known as the AdS/CFT correspondence.
@@
 The cylinder’s interior corresponds to the AdS bulk $M$ and its surface represents the conformal boundary $\partial M\simeq{\mathbb{R}}\times S^{d-1}$.
-The diamond-shaped subregion on $\partial M$ corresponds to $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{4},\eta_{\mu\nu})$.
+The diamond-shaped subregion on $\partial M$ corresponds to $d$-dimensional Minkowski spacetime $({\mathbb{R}}^{d},\eta_{\mu\nu})$.
 This subregion is enclosed by the red curves, which represent the future and past null infinity ($\mathscr{I}^{\pm}\subset\partial M$), and a single point $i^{0}\in\partial M$, which represents spatial infinity.
@@
 Two vertical double lines represent conformal infinity (i.e., the surface of the cylinder).
-The triangular subregion inside $M$ is covered by the horospherical chart (also known as the Poincare chart) given by (94), where $\infty>z>0$ inside $M$.
+The triangular subregion inside $M$ is covered by the horospherical chart (also known as the Poincaré chart) given by (96), where $\infty>z>0$ inside $M$.
 Each vertical hyperboloid represents a constant-$z$ hypersurface, which is a $d$-dimensional Minkowski spacetime with the metric $(\ell/z)^{2}\eta_{\mu\nu}$.
 The boundary is located at $z=0$.
-The AdS metric $g_{MN}$ in $(d+1)$-dimension can be expressed in the horospherical (Poincare) chart as
+The AdS metric $g_{MN}$ in $(d+1)$-dimension can be expressed in the horospherical (Poincaré) chart as
 ${\rm d}s^{2}_{(d+1)}=g_{MN}\,{\rm d}x^{M}{\rm d}x^{N}=\dfrac{\ell^{2}}{z^{2}}\left({\rm d}z^{2}+\eta_{\mu\nu}{\rm d}x^{\mu}{\rm d}x^{\nu}\right)\,$, (96)
@@
 As (part of) the $d$-dimensional boundary spacetime $(\partial M,g_{\mu\nu})$, one can also consider different geometries, such as $d$-dimensional de Sitter, AdS, or the static Einstein universe ${\mathbb{R}}\times S^{d-1}$ (see the metric (32) for $d=4$ case), other than the Minkowski spacetime.
-As for the $(d+1)$-dimensional bulk geometry $(M,g_{MN})$, one can, in principle, consider any asymptotically AdS spacetime, such as those including black holes inside but at large distances approaching the above AdS metric.
-In general, the AdS/CFT correspondence asserts the following: {itembox}[l]Conjecture 4.1.2 (AdS/CFT correspondence) A duality between a $(d+1)$-dimensional string theory defined in asymptotically Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) without gravity residing on its conformal boundary.
+As for the $(d+1)$-dimensional bulk geometry $(M,g_{MN})$ in the AdS/CFT correspondence, one can, in principle, consider any asymptotically AdS spacetime, such as those including black holes inside but at large distances approaching the above AdS metric.
+In general, the AdS/CFT correspondence asserts the following:
+  Conjecture 4.1.3 (AdS/CFT correspondence)   A duality between a $(d+1)$-dimensional string theory defined in asymptotically Anti-de Sitter (AdS) spacetime and a $d$-dimensional conformal field theory (CFT) without gravity residing on its conformal boundary.
 In this framework, string theory or quantum gravity phenomena in the bulk AdS spacetime can be fully encoded in the quantum dynamics of a conventional QFT on the boundary.
@@
 This is called the Fefferman-Graham (FG) expansion.
-For given $g^{(0)}_{\mu\nu}(x)$, all the subleading coefficients $g^{(n)}_{\mu\nu}(x)$ with $n<d$ are determined solely by the boundary metric $g^{(0)}_{\mu\nu}(x)$ and the curvature tensors with respect to $g^{(0)}_{\mu\nu}(x)$, while the coefficient $g^{(d)}_{\mu\nu}(x)$ is related to the expectation value of the boundary field stress-energy tensor $\langle T_{\mu\nu}\rangledeHaro$:2000vlm : {itembox}[l]Formula 4.1.3 (Holographic stress-energy tensor) The renormalized stress-energy tensor on the boundary quantum field is given in terms of the FG expansion coefficient as,
-$\langle T_{\mu\nu}\rangle=\dfrac{d\ell^{d-1}}{16\pi G_{d+1}}g^{(d)}_{\mu\nu}(x)+{X}_{\mu\nu}\,.$ (97)
-Here $X_{\mu\nu}$ represents gravitational anomaly, which vanishes when the boundary dimension $d$ is odd.
+For given $g^{(0)}_{\mu\nu}(x)$, all the subleading coefficients $g^{(n)}_{\mu\nu}(x)$ with $n<d$ are determined solely by the boundary metric $g^{(0)}_{\mu\nu}(x)$ and the curvature tensors with respect to $g^{(0)}_{\mu\nu}(x)$, while the coefficient $g^{(d)}_{\mu\nu}(x)$ is related to the expectation value of the boundary field stress-energy tensor $\langle T_{\mu\nu}\rangledeHaro$:2000vlm :
+  Formula 4.1.4 (Holographic stress-energy tensor)   The renormalized stress-energy tensor on the boundary quantum field is given in terms of the FG expansion coefficient as, $\langle T_{\mu\nu}\rangle=\dfrac{d\ell^{d-1}}{16\pi G_{d+1}}g^{(d)}_{\mu\nu}(x)+{X}_{\mu\nu}\,.$ (99)     Here $X_{\mu\nu}$ represents gravitational anomaly, which vanishes when the boundary dimension $d$ is odd.
 This formula stems from the fact that the AdS/CFT correspondence essentially equates the partition functions—and thus their effective actions—of the bulk and boundary theories.
@@
 Since the geometry near the horizon remains nearly classical during most of the black hole’s lifetime, especially for such large black holes, there appears to be no mechanism within the semiclassical framework that could halt or reverse this entropy growth.
-[…]
+Figure 16: An entangled pair of particles A and B in a spherically symmetric spacetime of black hole formation.
+The thin dotted hyperboloids depict curves of constant area radius $r$ and $\mathscr{H}$$^{+}$ represents the event horizon.
+Particle A falls into the black hole and moves toward smaller $r$, whereas B moves toward larger $r$.
+As a result, the distance between the two particles increases with time.
+Particle B eventually escapes to $\mathscr{I}^{+}$ as Hawking radiation.
+This process generates entanglement between the black hole interior and the outgoing Hawking radiation.
 However, as Page pointed out in Page:1993df ; Page:1993wv , if the evaporation process is ultimately unitary, the entanglement entropy should not increase indefinitely.
@@
 A pivotal insight in the effort to resolve the black hole information paradox is that entropy, particularly entanglement entropy, plays a fundamental role in the structure of quantum gravity.
-In semiclassical gravity, the entropy of a black hole is famously captured by the Bekenstein-Hawking formula, {itembox}[l]Formula 4.3.1 (Bekenstein-Hawking entroy)
-$S_{\text{BH}}=\frac{A}{4G\hbar}\,$ (98)
-where $A$ denotes the area of the event horizon as mentioned at the beginning of this section.
+In semiclassical gravity, the entropy of a black hole is famously captured by the Bekenstein-Hawking formula (93).
 This area-law scaling stands in sharp contrast to the volume-law behavior characteristic of entropy in local quantum field theory, hinting at a radical departure from conventional locality.
@@
 This conceptual leap gained concrete form in the context of the AdS/CFT correspondence, especially with the introduction of the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef .
-For a spatial subregion $A$ of the boundary CFT, the RT formula computes its entanglement entropy as {itembox}[l]Formula 4.3.2 (Ryu-Takayanagi formula)
-$S_{A}=\min_{\gamma}\left[\frac{\mathrm{Area}(\gamma_{A})}{4G}\right]\,$, (99)
-where $\gamma_{A}$ is the codimension-2 minimal surface in the bulk AdS geometry, anchored to $\partial A$ and homologous to $A$.
+For a spatial subregion $A$ of the boundary CFT, the RT formula computes its entanglement entropy as [^16] In the Ryu–Takayanagi formula, the entanglement entropy has a UV divergence of the form $S_{A}\propto\mathrm{Area}(\gamma_{A})/\epsilon^{2}$.
+Here $\epsilon$ is a short-distance cutoff in the boundary CFT: in AdS Poincaré coordinates, instead of extending the minimal surface all the way to the boundary at $z=0$, one terminates it at $z=\epsilon$, which regulates the divergence.
+  Formula 4.3.1 (Ryu-Takayanagi formula)   $S_{A}=\min_{\gamma_{A}}\left[\dfrac{k_{\rm B}\mathrm{Area}(\gamma_{A})}{4G\hbar}\right]\,.$ (100)     Here $\gamma_{A}$ is the codimension-2 minimal surface in the bulk AdS geometry, anchored to $\partial A$ and homologous to $A$.
 This prescription geometrizes entanglement: it relates a quantum information-theoretic quantity to a geometric property of spacetime, suggesting that entanglement is not merely a diagnostic tool but a foundational principle underlying gravitational physics.
@@
 In a holographic theory with a black hole in the bulk, this entropy must match the Bekenstein-Hawking entropy of the black hole:
-$S_{A}=S_{\text{BH}}=\frac{A_{\text{horizon}}}{4G}.$ (100)
+$S_{A}=S_{\text{BH}}=\dfrac{k_{\rm B}A_{\text{horizon}}}{4G\hbar}\,.$ (101)
 This matching of entropies in the fully mixed limit strongly supports the identification of the black hole horizon itself as the minimal surface $\gamma_{A}$ in the RT formula (100).
@@
 These bring us to the notion of quantum extremal surfaces (QES), developed by Faulkner, Lewkowycz, and Maldacena Faulkner:2013ana and further formalized by Engelhardt and Wall Engelhardt:2014gca .
-The QES prescription replaces the classical minimal-area surface with a surface that extremizes the generalized entropy: {itembox}[l]Formula 4.3.3 (Quantum extremal surface)
-$S_{\text{gen}}(\gamma)=\mbox{Ext}_{\gamma}\left[\frac{\mathrm{Area}(\gamma)}{4G}+S_{\text{bulk}}(\Sigma_{\gamma})\right]\,$, (101)
-where $\Sigma_{\gamma}$ is the bulk region bounded by $\gamma$ and the boundary region $A$, and $S_{\text{bulk}}(\Sigma_{\gamma})$ is the von Neumann entropy of quantum fields in that region.
-The entanglement entropy of region $A$ is then computed by extremizing this generalized entropy over all candidate surfaces $\gamma$ anchored to $\partial A$.
-Note that the idea of including quantum corrections from matter fields in this way is conceptually rooted in Bekenstein’s notion of generalized entropy Bekenstein:1974ax .
+The QES prescription replaces the classical minimal-area surface with a surface that extremizes the generalized entropy:
+$S_{\mathrm{gen}}(\gamma)=\dfrac{k_{\rm B}\mathrm{Area}(\gamma)}{4G\hbar}+S_{\mathrm{bulk}}(\Sigma_{\gamma})\,$, (102)
+where $\Sigma_{\gamma}$ is the bulk region on an achronal slice bounded by $\gamma$ and the boundary region $A$, and $S_{\text{bulk}}(\Sigma_{\gamma})$ is the von Neumann entropy of quantum fields in that region.
+The quantum extremal surface $\gamma_{A}$ is then defined as follows:
+  Formula 4.3.2 (Quantum extremal surface)   $S(A)=\mbox{min-ext}_{\gamma}S_{\mathrm{gen}}(\gamma).$ (103)     Here $\mbox{min-ext}_{\gamma}$ represents the following process: first search for extremal surface(s) $\gamma_{A}$ (homologous to $A$) that satisfy the extremality $\delta_{\gamma_{A}}S_{\mathrm{gen}}=0$, and then choose the one which attains the minimum among all such extremal surfaces if there are many such $\gamma_{A}$.
+The surface obtained by this min-ext prescription (among QES homologous to $A$) provides the correct entanglement entropy $S(A)$ Engelhardt:2014gca .
+Note that the inclusion of matter entanglement in this way is conceptually rooted in Bekenstein’s generalized entropy Bekenstein:1974ax .
 There, Bekenstein proposed that the total entropy should consist of both the area term and the matter entropy outside the horizon, anticipating the structure of $S_{\text{gen}}$ long before the advent of the AdS/CFT correspondence, even before the discovery of Hawking radiation.
@@
 Bulk operators in $\Sigma_{\gamma}$ can, in principle, be reconstructed from data in $A$.
-A major conceptual advance came with the development of the island formula, especially in the context of evaporating black holes coupled to an external non-gravitating bath.
+A major conceptual advance came with the development of the island formula, especially in the context of evaporating black holes coupled to an external non-gravitating bath Almheiri:2019hni .
+In practice, this means working in a regime where gravity is dynamical only in the black hole region, while the bath lives on a fixed background.
 This framework was pioneered by Almheiri, Engelhardt, Marolf, and Maxfield Almheiri:2019hni and Penington Penington:2019npb , and further developed by Almheiri:2019psf .
-In such a setup, the entanglement entropy $S_{R}$ of the Hawking radiation $R$ is given by {itembox}[l]Formula 4.3.4 (Island formula)
-$S_{R}=\mbox{Ext}_{I}\left[\frac{\mathrm{Area}(\partial I)}{4G}+S_{\text{bulk}}(R\cup I)\right]\,.$ (102)
-Here $I$ denotes a candidate island region inside the black hole, and $\partial I$ is its boundary (see Figure 18).
+In such a setup, the entanglement entropy $S_{R}$ of the Hawking radiation $R$ is given by
+  Formula 4.3.3 (Island formula)   $S_{R}=\mbox{min-ext}_{I}\left[\frac{k_{\rm B}\mathrm{Area}(\partial I)}{4G}+S_{\text{bulk}}(R\cup I)\right]\,.$ (104)     Here $I$ denotes a candidate island region inside the black hole, and $\partial I$ is its boundary (see Figure 18).
 At early times, the minimal configuration has no island, reproducing Hawking’s prediction of monotonically increasing radiation entropy.
 However, after the Page time, a new configuration appears where a nontrivial island emerges near the black hole horizon that becomes dominant, yielding a smaller generalized entropy.
-As a result of this island, the entropy of the radiation begins to decrease after the Page time, successfully reproducing the Page curve.
+As a result of this island, the entropy of the radiation begins to decrease after the Page time, successfully reproducing the Page curve [^17] In this argument, we assume the existence of a nice foliation of regular (partial) Cauchy surfaces which cover the entire process, ranging from black hole formation to the final stage (an immediate vicinity) of black hole evaporation.
+For asymptotically flat spacetimes, each of these time slices can be anchored at some boundary surface in the asymptotic region, sometime referred to as cutoff surface or anchor cruve Penington:2019npb .
+The calculation of QES and the determination of the Page time can be performed based on such time slices.
+.
 […]
@@
 While boundary entanglement patterns determine the bulk geometry, the consistency and causal structure of the bulk spacetime also impose powerful constraints on the boundary theory.
-A notable example is the derivation of boundary energy conditions, such as the averaged null energy condition (ANEC), from bulk causality and modular flow.
+A notable example is the derivation of boundary energy conditions, such as the averaged null energy condition (ANEC), from bulk causality.
 In the next section, we explore how this bidirectional logic further deepens our understanding of quantum gravity.
@@
 $\int_{-\infty}^{\infty}{\rm d}\lambda\,\langle T_{\mu\nu}k^{\mu}k^{\nu}(\lambda)\rangle\geqslant 0\,$, (105)
-where $k^{\mu}$ is a future-directed null vector and the expectation value is given by the holographic Formula 4.1.1 deHaro:2000vlm .
+where $k^{\mu}$ is a future-directed null vector and the expectation value is given by the holographic Formula 4.1.4 deHaro:2000vlm .
 The key insight in the Kelly-Wall argument is that any violation of the ANEC would imply a breakdown of boundary causality via its holographic encoding in bulk geometry.
@@
 More precisely, consider an asymptotically AdS spacetime $(M,g_{MN})$ and its conformal boundary $\partial M$.
-Suppose a pair of boundary two points $p,q\in\partial M$ are connected by a null geodesic $\gamma$ which is lying entirely in $\partial M$ and is achronal with respect to the boundary metric $g^{(0)}_{\mu\nu}$.
+Suppose a pair of boundary points $p,q\in\partial M$ are connected by a null geodesic $\gamma$ which is lying entirely in $\partial M$ and is achronal with respect to the boundary metric $g^{(0)}_{\mu\nu}$.
 If there is a timelike curve through the bulk $M$ which also connects the boundary two points $p$ and $q$, then the entire spacetime $M\cup\partial M$ is said to admit a bulk-shortcut.
-{itembox}[l]Theorem 4.4.1 (Gao-Wald theorem) Consider an asymptotically AdS spacetime $M$ which has no causal pathology, such as a naked singularity or causality violation, in the bulk.
+  Theorem 4.4.1 (Gao-Wald theorem)   Consider an asymptotically AdS spacetime $M$ which has no causal pathology, such as a naked singularity or causality violation, in the bulk.
 Then, under the (A)NEC and the null generic condition in the bulk, there exists no bulk-shortcut and the fastest possible causal curve connecting two boundary points $p,q$ is an achronal null geodesic $\gamma$ lying entirely on the boundary $\partial M$.
@@
 4.1, the conformal boundary $\partial M$ of $(d+1)$-dimensional AdS spacetime is globally the static Einstein universe in $d$-dimension, $\partial M\simeq{\mathbb{R}}\times S^{d-1}$.
-Let us consider a pair of two boundary points $p,q\in\partial M$ such that $p$ and $q\in J^{+}(\partial M)$ can be connected by an achronal null geodesic $\gamma$ with the tangent vector $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ lying entirely in $\partial M$, and $q$ corresponds to the antipodal point of $p$ in spatial section $S^{d-1}$, as depicted in Figure 19.
+Let us consider a pair of boundary points $p,q\in\partial M$.
+Fix $p\in\partial M$ and choose $q\in J^{+}(p,{\partial M})$ so that there exists an achronal boundary null geodesic $\gamma\subset\partial M$ with tangent $k^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}\lambda$ connecting $p$ to $q$; take $q$ to be the antipodal point of $p$ on $S^{d-1}$ (Fig.
+19).
 According to Gao-Wald theorem, under the bulk NEC, there is no bulk-shortcut connecting $p,q$ and $\gamma$ must be the fastest causal curve connecting the two point.
@@
 Note that when the boundary dimension $d$ is odd, $X_{\mu\nu}$ in (99) identically vanishes, while when $d$ is even, $X_{\mu\nu}$ must be taken into account.
-Then, applying Gao-Wald theorem, we find the following inequality: {itembox}[l]Definition 4.5.1 (Holographic weighted ANEC)
-$\int_{\lambda_{-}}^{\lambda_{+}}\!\!\ {\rm d}\lambda\>\eta^{d}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\geqslant C\,.$ (104)
+Then, applying Gao-Wald theorem, we find the following inequality:
+  Definition 4.5.1 (Holographic weighted ANEC)   $\int_{\lambda_{-}}^{\lambda_{+}}\!\!\ {\rm d}\lambda\>\eta^{d}\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}\geqslant C\,.$ (106)     where $\eta$ is the magnitude of the Jacobi field of the null geodesic congruence $\gamma$, representing the separation of points between the two adjacent null geodesics on $\partial M$.
   • When the boundary dimension $d$ is odd, the lower bound given by constant $C$ in the right-hand side vanishes, and this inequality is manifestly conformally invariantIizuka:2019ezn , and is called the conformally invariant ANEC (CANEC).
@@
 Perhaps most strikingly, this result reverses the typical logic of holography: rather than using boundary data to constrain the bulk, bulk causality is used to derive nontrivial dynamical constraints on the boundary theory.
-The argument elegantly ties together extremal surfaces, entanglement entropy, and modular flow to derive a physical energy condition from first principles.
+The argument elegantly ties together extremal surfaces, entanglement entropy to derive a physical energy condition from first principles.
 Acknowledgments